### v1

# Results of Fractional Rough Burgers equation in $H^{s}$ space and its application

Abstract
In this paper, we study the well-posedness of Fractional Rough Burgers equation driven by space-time noise in $H^{s}(\mathbb{T})$ space. For the higher dissipation $\gamma\in(\frac{4}{3},2]$, we establish local well-posedness. Global well-posedness is further obtained when $\gamma$ is restricted to the interval $(\frac{3}{2},2]$. For the lower dissipation $\gamma\in(\frac{5}{4},\frac{4}{3}]$, we use the regularity analysis derivation the para-controlled solution.
Keywords: Rough PDEs; Paracontrolled solution; Well-posedness; Shallow water.
2020 Mathematics Subject Classification: 60H15, 60H17, 42B37, 35Q35
1 Introduction
In this paper, we will consider Fractional Rough Burgers (FRB) equation with weak dissipation in $H^{s}(\mathbb{T})$
$\partial_{t}u-\Lambda^{\gamma}u=\frac{1}{2}\partial_{x}(u^{2})+\xi,\quad u(0,x)=u_{0}.$ (1.1)
Here $\mathbb{T}$ is some toru satisfied $\int_{\mathbb{T}}e^{ikx}=0$ for any $k\in\mathbb{Z}/\{0\}$(for example $\mathbb{R}/2\pi\mathbb{Z}$) and $\Lambda^{\gamma}$ is a Fourier multiplier satisfied $\mathcal{F}(\Lambda^{\gamma}f)=-|\xi|^{\gamma}\mathcal{F}f$. $\xi$ is a centered Gaussian space-time random distribution with covariance
$\mathbb{E}[\xi(\omega,t,x)\xi(\omega,s,y)]=\delta_{0}(t-s)\delta_{0}(x-y),\quad t,s\geq 0,\quad x,y\in\mathbb{T},$ (1.2)
where $\delta_{0}(x)$ is a Dirac measure satisfied $\delta_{0}(x)=0$ for $x\neq 0$ and $\int_{\mathbb{T}}\mathrm{d}\delta_{0}(x)=1$.
Equation (1.1) is a classical model in SPDE driving by additional noise. Similar research topics have attracted extensive study in recent years, building upon the foundational work of Martin Hairer in [Hai13, Hai14] and Gubinelli, Perkowski in [GIP15, GP17]. In this paper, we primarily investigate how the degree of dissipation affects the equation and solutions. For the high dissipation which $\frac{5}{3}<\gamma\leq 2$, the Fractional Rough Burgers equation with nonlinearity $F(u)\partial_{x}u$ replacing $\frac{1}{2}\partial_{x}(u^{2})$ has been studied in [GIP15]. Additionally, they build the global well-posedness in $\mathcal{C}^{s}$ for $\frac{1}{3}<s<\frac{\gamma}{2}-\frac{1}{2}$.
Note that the performance of lower dissipation $\gamma<\frac{5}{3}$ in $H^{s}$ space is unknown, where Sobolev spaces $H^{s}$ are of significant mathematical and physical relevance in the study Burgers equation and Degasperis-Procesi(DP) equation. For more details, one can be refer to [Yin03, Yin04, GL11, GL15]. In this paper, we will consider equation (1.1) with lower dissipation $\frac{4}{3}<\gamma\leq 2$ for $u_{0}\in H^{\frac{1}{2}+}$ and $\frac{5}{4}<\gamma\leq\frac{4}{3}$ for $u_{0}\in H^{\frac{1}{2}+}(\mathbb{T})\cap\mathcal{C}^{\frac{1}{2}+}(\mathbb{T})=W^{\frac{1}{2}+}(\mathbb{T})$.
It is also need to particularly noteworthy that the first author and co-authors has previously studied the randomized initial data problems for both the DP equation and Burgers equation
$\partial_{t}u-\Lambda^{\gamma}u=uu_{x},\quad u_{0}^{\omega}=\sum_{n\in\mathbb{Z}}h_{n}(\omega)\varphi(\Lambda-n)u_{0}.$ (1.3)
in [CZG24] for $\gamma\in(\frac{5}{4},\frac{3}{2})$, which is an another hot topic in the field of SPDEs. We omit detailed discussion here and refer interested readers to [Bou96, BT08b, BT08a, CG15, BOP15, BOP19, DNY22, DNY24]. Therefore, this paper can be viewed as a natural continuation of that earlier work under additional noise driving.
The primary objective of this paper is to establish fundamental results for the FRB equation in Sobolev spaces $H^{s}$ and extend analogous conclusions to the DP equations. We will address the potential difficulties associated with these problems, as well as the methods we will employ, by organizing the discussion into three main sections.
1.1 Probabilistic setting
To have a better understanding of the problem, we first study (1.1) by writing it as mild form
$u=P(t)u_{0}+\frac{1}{2}\int_{0}^{t}P(t-s)\partial_{x}(u^{2})\mathrm{d}s+X(t).$ (1.4)
Here $P(t)=e^{t\Lambda^{\gamma}}$ is the fractional heat flow, and $X(t)=\int_{0}^{t}P(t-s)\xi(s)\mathrm{d}s$ denote the solution of linear equation
$\partial_{t}X(t)-\Lambda^{\gamma}X(t)=\xi,\quad X_{0}=0.$ (1.5)
Note that $X(t)\in C_{T}\mathcal{C}^{\alpha}(\mathbb{T})$ a.s for any $\alpha<\frac{\gamma}{2}-\frac{1}{2}$, which is applied to the study of equations describing thermal dissipative structures such as KPZ equation [Hai13] and Burgers equation [GIP15, GP17] to establish well-posedness. This also applies to Sobolev spaces $H^{s}$ (See Lemma 2.7) and keep in mind that $X\in C_{T}W^{\alpha}(\mathbb{T})$ for $\alpha<\frac{\gamma}{2}-\frac{1}{2}$.
Note that the range of $\alpha$ is determined by dissipation $\gamma$, our paper gives two works in different range. One is higher regularity $\alpha\in(\frac{1}{6},\frac{1}{2})$, in such case, we can directly establish the well-posedness in $H^{\alpha}(\mathbb{T})$. One is lower dissipation $\alpha\in(\frac{1}{8},\frac{1}{6})$(See Remark 3.1), in such case we need to work in $W^{\alpha}(\mathbb{T})$. A common feature is that we have isolated a higher-regularity term (More precisely $H^{\frac{1}{2}+}(\mathbb{T})$) in the solution. This process is made possible in lower disspation $\gamma\in(\frac{5}{4},\frac{4}{3}]$ with the assistance of para-controlled method which is a powerful method to study rough or singular SPDE in [GIP15, GP17, GKO23, CC18].
A more in-depth consideration is that how lower the dissipation be? Let us consider FRB driven by $|D|^{\beta}\xi$
$u_{t}-\Lambda^{\gamma}u=\frac{1}{2}\partial_{x}(u^{2})+|D|^{\beta}\xi,\quad u(0,x)=u_{0}$ (1.6)
using the similar method in [DNY22, DNY24], we propose a conjecture regarding the relationship between $\gamma$ and $\beta$
$\gamma\geq\beta+1$ (1.7)
in final section. Similar methods have been used in [DNY24] to study probabilistic critical problems for the heat equation, wave equation, and Schrödinger equation.
1.2 Regularity analysis
After resolving the regularity issues for Gaussian processes, the next challenge lies in the study of well-posedness. Consider the Burgers equation with addition term $f$
$\partial_{t}v-\Lambda^{\gamma}v=\frac{1}{2}\partial_{x}(v^{2})+f,\quad v(0,x)=v_{0},$ (1.8)
where $f$ may depend on $u$. It takes the different structure of solution in two different range $\gamma\in(\frac{4}{3},2]$ and $\gamma\in(\frac{5}{4},\frac{4}{3}]$.
More precisely, let $f(v)=\partial_{x}(vX)+\frac{1}{2}\partial_{x}(X^{2})$, where $X$ satisfied (1.5). Then $f(v)$ gains at most $\alpha$-order Sobolev regularity when $v$ belongs to $H^{\frac{1}{2}+}$. It derive the relation between $\alpha$ and $\gamma$ which satisfied $\alpha+\gamma\geq\frac{3}{2}$, since $v$ can gain $\alpha+\gamma-1$ at most. For the lower $\alpha$ and $\gamma$, we use the Little-wood Paley decomposition in [BCD11] which rewrite $vX$ and $X^{2}$ as
$vX=T_{v}X+R(v,X)+T_{X}v,\quad X^{2}=2T_{X}X+R(X,X).$ (1.9)
It takes the higher regularity $T_{X}v$, $R(v,X)$ and $R(X,X)$ gains since $X$ has $\alpha$-order Höler regularity. The only difficult for the case of smaller $\alpha$ is $T_{v}X$ and $2T_{X}X$ which can be countered by para-controlled solution(See more details in Proposition 4.1), and it derives the relation $2\alpha-1+\gamma\geq\frac{3}{2}$.
1.3 Shallow water equation
Next, we introduce the Degasperis-Procesi(DP) equation, which is one type of shallow water equation, has following form
$m_{t}+um_{x}+3u_{x}m=0,\quad m(t,x)=(1-\partial_{x}^{2})u(t,x),\quad u(0,x)=u_{0}(x).$ (1.10)
Then we can define the fluid velocity $u(t,x):=g\ast m:=(1-\partial_{x}^{2})m$ and rewrite the DP equation as
$u_{t}+\frac{1}{2}\partial_{x}(u^{2})+\frac{3}{2}(1-\partial_{x}^{2})^{-1}\partial_{x}(u^{2})=0,\quad u(0,x)=u_{0}(x).$ (1.11)
We provide following citation [DP99, DHH02] to facilitate a deeper understanding of the physical background for this class of equations. In particular, a systematic discussion can be found in [AG10]. For the mathematical properties of this equation, we refer readers to the following literature. Of particular note, the local well-posedness to (1.10) with initial data $H^{s}$, $s\geq\frac{3}{2}$ was proved in [LY06], and also the precise blow-up scenario and a blow-up result were derived in [LY07].
In this paper, we also consider the nonlocal Rough DP equation with space-time noise $\xi$
$u_{t}+\Lambda^{\gamma}u+\frac{1}{2}\partial_{x}(u^{2})+\frac{3}{2}(1-\partial_{x}^{2})^{-1}\partial_{x}(u^{2})=\xi,\quad u(0,x)=u_{0}(x).$ (1.12)
as an application of our result for Rough Burgers equation.
1.4 Main result
We now present the main conclusions.
Theorem 1.1.
Let $u_{0}(x)\in H^{\frac{1}{2}+\delta}(\mathbb{T})$, for some small $\delta>0$, and $\gamma\in(\frac{4}{3},2]$, then there exists a postive $T^{*}$ and a solution $u(t,x)$ of equation (1.1) satisfied
$u(t,x)=X(t,x)+v(t,x),$
where $v(t,x)\in C_{T^{*}}H^{\frac{1}{2}+\delta}(\mathbb{T})$ a.s. and $X(t,x)$ is satisfied (1.5) with $X(t,x)\in C_{T^{*}}H^{\alpha}(\mathbb{T})$ for $\alpha\in(\frac{1}{6},\frac{\gamma}{2}-\frac{1}{2})$, a.s. Moreover we have $u\in C_{T^{*}}H^{\alpha}(\mathbb{T})$ a.s.
Remark 1.1.
’a.s.’ represent almost surely. We say event $A$ is almost surely if and only if $\mathbb{P}(A)=1$, where $\mathbb{P}$ represent the probability associated with $(\Omega,\mathcal{F},\mathbb{P})$. We will omit this description in most situation, except we need distinct such meaning.
Remark 1.2.
Recall the equation in [GIP15]
$\partial_{t}u-\Lambda^{\gamma}u=G(u)\partial_{x}u+\xi,\quad u(0,x)=u_{0},$ (1.13)
where $G\in C_{b}^{3}(\mathbb{R}^{d},L(\mathbb{R}^{d},\mathbb{R}^{d}))$. The main difference between our result and [GIP15] lies in the fact that our solution can back to the original equation (1.1). Since $G(u)\partial_{x}u$ is ill-defined when $\alpha\in(\frac{1}{3},\frac{\gamma}{2}-\frac{1}{2})$, they actually study the “generalized" solution, which is the limit in probability sense of $u^{\epsilon}$ satisfied following mollified equation
$\partial_{t}u^{\epsilon}-\Lambda^{\gamma}u^{\epsilon}=G(u^{\epsilon})\partial_{x}u^{\epsilon}+\xi^{\epsilon},\quad u(0,x)=u_{0}.$ (1.14)
Here $\xi^{\epsilon}=\epsilon^{-1}\psi(\epsilon\cdot)\ast\epsilon$ for some $\psi\in\mathcal{S}$ with $\int\psi\mathrm{d}t=1$. Similar approaches are frequently used for some singular SPDEs in [CC18, Hai13, GP17].
When $\gamma>\frac{3}{2}$, we can get the global result as follows.
Theorem 1.2.
For $\gamma>\frac{3}{2}$, Rough Burgers equation (1.1) has a global weak solution $u=v+X$ in $C([0,T];L^{2}(\mathbb{T}))\cap L^{2}([0,T];H^{\frac{\gamma}{2}}(\mathbb{T}))$ for any $T\geq 0$ . Moreover, for $\gamma>\frac{5}{3}$, such solution is the global strong solution of (1.1).
Remark 1.3.
Theorem 1.1 and 1.2 is also valid for Rough DP equation. In fact, the nonlinear term $\frac{3}{2}(1-\partial_{x}^{2})^{-1}\partial_{x}(u^{2})$ in the DP equation does not introduce additional difficulties for the local well-posedness problem. For global results, we need to utilize the special structure of the DP equation to derive energy estimates as Proposition 3.3 and Lemma 3.4.
For the lower dissipation $\gamma\in(\frac{5}{4},\frac{4}{3}]$, let $\mathcal{Q}$ be the linear evolution of $\partial_{x}X$ such as
$\partial_{t}\mathcal{Q}(t)-\Lambda^{\gamma}\mathcal{Q}(t)=\partial_{x}X(t),\quad\mathcal{Q}(0)=0.$ (1.15)
We let $(u,u^{\prime}u^{\sharp})$ be a triple satisfied
$u=X+u^{\prime}\prec\!\!\!\prec\mathcal{Q}+u^{\sharp},$ (1.16)
where $u^{\prime}\prec\!\!\!\prec\mathcal{Q}$ is the temporal mollified version of $T_{u^{\prime}}\mathcal{Q}$(See (2.1)). $u^{\sharp}$ is the higher regularity term.
Theorem 1.3.
Let $u_{0}\in\mathcal{C}^{\frac{1}{2}+\delta}\cap H^{\frac{1}{2}+\delta}=W^{\frac{1}{2}+\delta}$ for some $\delta>0$, $\gamma\in(\frac{5}{4},\frac{4}{3}]$. Define $\mathcal{L}=\partial_{t}-\Lambda^{\gamma}$, then there exists a positive $T^{*}$ and a unique triple $(u,u^{\prime},u^{\sharp})\in C_{T}W^{\alpha}\times C_{T}W^{\frac{1}{8}+\delta}\times C_{T}W^{\frac{1}{2}+\delta}$ satisfied
$u^{\prime}=$ $X+2u^{\prime}\prec\!\!\!\prec u^{\sharp},$ (1.17)
$\mathcal{L}u^{\sharp}=$ $\mathcal{L}u-\mathcal{L}X-\mathcal{L}u^{\prime}\prec\!\!\!\prec\mathcal{Q},$
$u=$ $X+u^{\prime}\prec\!\!\!\prec\mathcal{Q}+u^{\sharp},$
where $X(t)$ is satisfied (1.5). Moreover, $u$ is a solution of (1.1).
Remark 1.4.
$\prec$ and $\prec\!\!\!\prec$ is referred to as the para-product and modified para-product, which will be introduced in Lemma 2.1 and (2.1). Let $v:=u^{\prime}\prec\!\!\!\prec\mathcal{Q}$, then $v$ satisfied the equation
$\partial_{t}v-\Lambda^{\gamma}v=^{\prime}regular\;term^{\prime}+u^{\prime}\prec\partial_{x}X$
This is a term that will appear in the equation for $u^{\sharp}$. Our strategy is to use a specific $u^{\prime}$ with $\mathcal{L}(u^{\prime}\prec\!\!\!\prec\mathcal{Q})$ to cancel out this term to make sure that $u^{\sharp}$ can gains higher regularity (See Proposition 4.1), we will show it in Section 4.
Remark 1.5.
In fact, we can’t prove such result with only $H^{\alpha}$ space under our setting. Let $(u^{\prime},\mathcal{Q},u^{\sharp})\in(\mathcal{C}^{\alpha}\times H^{\beta}\times H^{\eta})$, then $\alpha+\beta=\eta$ is always valid. Since $\mathcal{Q}$ can gains $(\alpha-1+\gamma)$-order Sobolev regularity, then $u^{\sharp}$ gains $(2\alpha-1+\gamma)$-order Sobolev regularity at most by proof of Lemma B.3.To establish the $C^{\alpha}$ Hölder regularity of $u^{\prime}$, it is necessary for the remainder $u^{\sharp}$ to possess Sobolev regularity of order at least $H^{\alpha+\frac{1}{2}+}$, this is valid only when $\gamma>\frac{4}{3}$.
Remark 1.6.
We can extend the range of $\gamma\in(1,2]$, which means the $\alpha\in(0,\frac{1}{2})$. We omit these proofs because in this paper, since we primarily aim to demonstrate the effectiveness of the paracontrolled method in enhancing the regularity of solutions.
1.5 Organization of the paper
This paper is organized as follows. In Section 2, we introduce some notations and basic tools. In Section 3, we prove Theorem 1.1 and Theorem 1.2 for FRB equation and Remark 1.3 for Rough DP equation. In the Section 4, we prove Theorem 1.3 for para-controlled solution. In the Section 5, we give a conjecture of the range of $\gamma$ and give some future directions.
2 Basic tool
In this section, we give some useful tools
2.1 Littlewood-Paley theory
In this section, we introduce some basic stochastic and deterministic tools, more useful technical estimation tools will be introduced in the appendix.
The first key tool is Littlewood-Paley decomposition. We introduce a operator named “nonhomogeneous dyadic blocks” $\Delta_{j}$ for $j\geq-1$, and satisfied for any $u$ belong to $\mathcal{S}^{\prime}(\mathbb{T}^{d})$ (Its also valid for $\mathbb{R}^{d}$), we have decomposition
$u=\sum_{j\geq-1}\Delta_{j}u.$
More precisely, its associated with two functions $\rho_{l}(k)$ and $\chi(k)$ under Fourier mode such that
$\mathcal{F}(\Delta_{j}u)(k)=\rho(2^{-l}k)u_{k},\quad l\geq 0\quad and\quad\mathcal{F}(\Delta_{-1}u)(k)=\mathcal{\chi}(k)u_{k}.$
Here $\chi$ and $\rho$ are two radial functions, valued in the interval $[0,1]$, belonging respectively to $\mathcal{D}(B(0,\frac{4}{3}))$ and $\mathcal{D}(\{k\in\mathbb{Z}^{d}:\frac{3}{4}\leq k\leq\frac{8}{3}\})$, such that
$\chi(k)+\sum_{l\geq 0}\rho(2^{-l}k)=1,\quad k\in\mathbb{Z}^{d}.$
It can define $B_{p,r}^{s}(\mathbb{T}^{d})$ which contains Hölder space $\mathcal{C}^{\alpha}(\mathbb{T}^{d}):=B_{\infty,\infty}^{s}(\mathbb{T}^{d})$ for $\alpha\in(0,1)$ and Sobolev space $H^{s}(\mathbb{T}^{d}):=B_{2,2}(\mathbb{T}^{d})$ such that
$B_{p,r}^{s}(\mathbb{T}^{d}):=\{u:(\sum_{l\geq-1}2^{lsr}\|\Delta_{l}u\|_{L^{p}(\mathbb{T}^{d})}^{r})^{\frac{1}{r}}<\infty\},$
also define the norm. Under this setting, we introduce para-product decomposition which give a product rules in $\mathcal{S}^{\prime}(\mathbb{T}^{d})$. Let $S_{j}u=\sum_{k\leq j-1}\Delta_{k}u$, then we have the estimate
Lemma 2.1 ([BCD11]).
Define $f\prec g:=\sum_{j}S_{j-1}f\Delta_{j}g$, $f\succ g=g\prec f$, $f\circ g:=\sum_{|k-j|\leq 1}\Delta_{k}f\Delta_{j}g$, let $p,r\in[1,\infty]$ and $s\in\mathbb{R}$, then we have following estimates
  • 1. If $k\in\mathbb{N}$, then
    $\|f\prec g\|_{B^{s}_{p,r}}\leq C\|f\|_{L^{\infty}}\|D^{k}g\|_{B_{p,r}^{s-k}}.$
  • 2. If $k\in\mathbb{N}$, $t<0$ and $\frac{1}{r}=\min(1,\frac{1}{r_{1}}+\frac{1}{r_{2}})$, then
    $\|f\prec g\|_{B_{p,r}^{s+t}}\leq C\|f\|_{B_{\infty,r_{1}}^{t}}\|D^{k}g\|_{B_{p,r_{2}}^{s-k}}.$
  • 3. If $\frac{1}{p}=\frac{1}{p_{1}}+\frac{1}{p_{2}}\leq 1$, $\frac{1}{r}=\frac{1}{r_{1}}+\frac{1}{r_{2}}\leq 1$ and $s=s_{1}+s_{2}>0$, then
    $\|f\circ g\|_{B_{p,r}^{s}}\leq C\|f\|_{B_{p_{1},r_{1}}^{s_{1}}}\|g\|_{B_{p_{2},r_{2}}^{s_{2}}}.$
By above estimate in $H^{s}$ space, we can check that Moser-type estimate hold.
Lemma 2.2 (Moser-type estimate).
Let $s>0$, $\epsilon>0$, then following estimates hold:
  • (1) for $0<s<\frac{1}{2}$,
    $\|fg\|_{H^{2s-\frac{1}{2}}}\leq C\|f\|_{H^{s}}\|g\|_{H^{s}}.$
  • (2) for $s>\frac{1}{2}$,
    $\|fg\|_{H^{s}}\leq C\|f\|_{H^{s}}\|g\|_{H^{s}}.$
  • (3) for $s=\frac{1}{2}$,
    $\|fg\|_{H^{s-\epsilon}}\leq C\|f\|_{H^{s}}\|g\|_{H^{s}}.$
To explain the para-product, it must be notice that if $f$ and $g$ are two distribution, we can’t generally multiply them. A sufficient condition to do that is Lemma 2.1 holding which need $s_{1}+s_{2}\geq 0$ at least. Moreover, we call that $f\prec g$ and $f\succ g$ para-products since they are always well defined, and $f\circ g$ resonant term since this product limits the range in which their frequencies are close.
For our convenient, give following notation
$W_{p}^{\alpha}=B_{p,p}^{\alpha},\quad W^{\alpha}=W_{2}^{\alpha}\cap W_{\infty}^{\alpha},\quad$
Note that $W^{\alpha}$ is a Banach algebra when $\alpha>0$.
Lemma 2.3.
If $f\in W^{\alpha}$, then $f\in W_{p}^{\alpha}$ for any $p\geq 2$. Moreover, we have estimate
$\|f\|_{W_{p}^{\alpha}}\leq\|f\|_{H^{\alpha}}^{\frac{2}{p}}\|f\|_{\mathcal{C}^{\alpha}}^{\frac{p-2}{p}}.$
Proof.
By definition of $B_{p,p}^{\alpha}$ and interpolation in $L^{p}$, we have
$\|f\|_{B_{p,p}^{\alpha}}^{p}=$ $\sum_{j\geq-1}2^{j\alpha p}\|\Delta_{j}f\|^{p}_{L^{p}}$
$\leq$ $\sum_{j\geq-1}2^{j\alpha p}(\|\Delta_{j}f\|^{\frac{2}{p}}_{L^{2}}\|\Delta_{j}f\|^{\frac{p-2}{p}}_{L^{\infty}})^{p}$
$\leq$ $(\sup_{j\geq-1}2^{j\alpha}\|\Delta_{j}f\|_{L^{\infty}})^{p-2}\sum_{j\geq-1}2^{j\alpha 2}\|\Delta_{j}f\|^{2}_{L^{2}}$
$\leq$ $\|f\|_{H^{\alpha}}^{2}\|f\|_{B_{\infty,\infty}^{\alpha}}^{p-2}.$
Then our lemma is proved. ∎
Sometimes it is also necessary to use a modified version of the paraproduct. Such operators have been introduced in [GIP15, GP17], and are defined as follows: Let $\varphi\in C^{\infty}(\mathbb{R},\mathbb{R}_{+})$ be nonnegative with compact support contained in $\mathbb{R}_{+}$ and with total mass 1, define $Q_{i}$ for all $i\geq 1$ as follows
$Q_{i}:C\mathcal{C}^{\beta}\to C\mathcal{C}^{\beta},\quad Q_{i}f(t)=\int_{\mathbb{R}}2^{-\gamma i}\varphi(2^{\gamma i}(t-s))f(s\wedge 0)\mathrm{d}s.$
With the help of $Q_{i}$, we define a modified para-product
$f\prec\!\!\!\prec g=\sum_{i}(Q_{i}S_{i-1}f)\Delta_{i}g.$ (2.1)
Similar to estimate for the paraproduct in Lemma 2.1, we also have estimates for this operator.
Lemma 2.4.
For any $\beta\in\mathbb{R}$ and $p\geq 2$, we have
$\|f\prec\!\!\!\prec g(t)\|_{W^{\beta}_{p}}\leq C\|f\|_{C_{t}L^{\infty}}\|g\|_{W^{\beta}_{p}},$
for all $t>0$. And for $\alpha<0$ furthermore
$\|f\prec\!\!\!\prec g(t)\|_{W^{\alpha+\beta}_{p}}\leq C\|f\|_{C_{t}\mathcal{C}^{\alpha}}\|g\|_{W^{\beta}_{p}}.$
The remaining estimates will be presented in the Appendix B. Finally, we introduce a class of temporal regularity spaces, which have been referenced in [GIP15, GP17, Hai13, CC18]. Define the such norm with time $T$ as follows. $C_{T}X=C([0,T],X)$ for the space of continuous maps from $[0,T]$ to $X$, equipped with the norm $\sup_{t\in[0,T]}\|\cdot\|_{X}$. For $\alpha\in(0,1)$, we also define $C_{T}^{\alpha}X$ as a space of $\alpha$-Hölder continuous continuous functions from $[0,T]$ to $X$ equipped with the norm $\|f\|_{C_{T}^{\alpha}X}=\sup_{0\leq s<t\leq T}\frac{\|f_{s,t}\|_{X}}{|t-s|^{\alpha}}$. For our convenient, for $\alpha\geq 0$ we have following notation
$\mathcal{W}_{p}^{0}(T)=C_{T}L^{p},\quad\mathcal{W}_{p}^{\alpha}(T)=C_{T}B_{p,p}^{\alpha}\cap C^{\frac{\alpha}{\gamma}}_{T}L^{p},\quad\mathcal{W}^{\alpha}(T)=\mathcal{W}_{2}^{\alpha}(T)\cap\mathcal{W}_{\infty}^{\alpha}(T).$
Its easy to check that $\mathcal{W}^{\alpha}(T)$ is a Banach algebra when $\alpha\geq 0$. For $\alpha<0$, we also give a definition $\mathcal{W}^{\alpha}_{p}(T)=C_{T}B_{p,p}^{\alpha}$ Next, we introduce some useful estimate
Lemma 2.5.
Let $\alpha\in\mathbb{R}$, $\epsilon>0$, $\rho\in(0,\gamma)$ satisfied $\gamma\leq\rho+\epsilon$ and $\alpha+\rho<\gamma$. Let $f\in\mathcal{W}_{\infty}^{\epsilon}(T)$, $g\in C_{T}W^{\alpha+\rho}(T)$ and $\mathcal{L}g\in C_{T}W^{\alpha}$. Then
$\|f\prec\!\!\!\prec g\|_{\mathcal{W}_{\infty}^{\alpha+\rho}(T)}\lesssim T^{1-\frac{\rho}{\gamma}}\|f\|_{\mathcal{W}_{\infty}^{\epsilon}(T)}(\|g\|_{C_{T}\mathcal{C}^{\alpha+\gamma-\epsilon}}+\|\mathcal{L}g\|_{C_{T}\mathcal{C}^{\alpha}}),$
and
$\|f\prec\!\!\!\prec g\|_{\mathcal{W}_{2}^{\alpha+\rho}(T)}\lesssim T^{1-\frac{\rho}{\gamma}}\|f\|_{\mathcal{W}_{\infty}^{\epsilon}(T)}(\|g\|_{C_{T}H^{\alpha+\gamma-\epsilon}}+\|\mathcal{L}g\|_{C_{T}H^{\alpha}}).$
Proof.
We only prove the case of $\mathcal{W}_{2}^{\alpha+\rho}(T)$ and the $p=\infty$ is similar. By Lemma A.3, Lemma 2.4 and Lemma B.3, we have
$\|f\prec\!\!\!\prec g\|_{\mathcal{W}_{2}^{\alpha+\rho}(T)}\lesssim$ $\|f\prec\!\!\!\prec g(0)\|_{\mathcal{W}_{2}^{\alpha+\rho}(T)}+T^{1-\frac{\rho}{\gamma}}\|\mathcal{L}(f\prec\!\!\!\prec g)\|_{C_{T}H^{\alpha}}$
$\leq$ $T^{1-\frac{\rho}{\gamma}}\|\mathcal{L}(f\prec\!\!\!\prec g)-f\prec\!\!\!\prec\mathcal{L}g\|_{C_{T}H^{\alpha}}+\|f\prec\!\!\!\prec\mathcal{L}g\|_{C_{T}H^{\alpha}}$
$\leq$ $T^{1-\frac{\rho}{\gamma}}\|f\|_{C_{T}\mathcal{C}^{\epsilon}}\|g\|_{C_{T}H^{\alpha+\gamma-\epsilon}}+\|f\|_{C_{T}\mathcal{C}^{\epsilon}}\|\mathcal{L}g\|_{C_{T}H^{\alpha}}$
$\leq$ $T^{1-\frac{\rho}{\gamma}}\|f\|_{C_{T}\mathcal{C}^{\epsilon}}(\|g\|_{C_{T}H^{\alpha+\gamma-\epsilon}}+\|\mathcal{L}g\|_{C_{T}H^{\alpha}})$
$\leq$ $T^{1-\frac{\rho}{\gamma}}\|f\|_{C_{T}\mathcal{C}^{\epsilon}}(\|g\|_{C_{T}H^{\alpha+\rho}}+\|\mathcal{L}g\|_{C_{T}H^{\alpha}}).$
Since $\gamma-\epsilon\leq\rho$, then we prove the lemma. ∎
2.2 Stochastic tools
In this little section, we introduce some stochastic tools we need in this paper. The first tools is the Kolmogorov’s continuity criterion
Lemma 2.6 ([LG16]).
Let $X=(X(t))_{t\in[0,T]}$ be a random process and take values in a complete metric space $(E,d)$. If there exist $p,\epsilon,C>0$ such that, for every $s,t\in[0,T]$,
$\mathbb{E}[d(X(s),X(t))^{p}]\leq C|t-s|^{1+\epsilon}.$ (2.2)
Then, there is a modification $\tilde{X}$ of $X$ whose sample paths are Hölder continuous with exponent $\alpha$ for every $\alpha\in(0,\frac{\epsilon}{p})$: This means that, for every $\omega\in\Omega$ and every $\alpha\in(0,\frac{\epsilon}{p})$, there exists a finite constant $C_{\alpha}(\omega)$ such that, for every $s,t\in I$,
$d(\tilde{X}(s,\omega),\tilde{X}(t,\omega))\leq C_{\alpha}(\omega)|t-s|^{\alpha},$
where $\tilde{X}(t)$ is a modification of $X(t)$ with continuous sample paths.
Now we give a basic analysis for white noise $\xi$. For the linear evolution $X(t):=\int_{0}^{t}e^{(t-s)\Lambda^{\gamma}}\xi(s)\mathrm{d}s$, given a decomposition
$X(t)=\int_{-\infty}^{t}e^{(t-s)\Lambda^{\gamma}}\xi(s)\mathrm{d}s-e^{t\Lambda^{\gamma}}\int_{-\infty}^{0}e^{-s\Lambda^{\gamma}}\xi(s)\mathrm{d}s.$ (2.3)
Its not difficult to show that the second term is smooth, which means all the irregularity of $X(t)$ is described by the first term of the right hand side of (2.3). Hence we define $Y(t):=\int_{-\infty}^{t}e^{(t-s)\Lambda^{\gamma}}\xi(s)\mathrm{d}s$. For $X(t)$ and $Y(t)$, we have following lemma.
Lemma 2.7.
The spatial Fourier transform $Y_{k}=\mathcal{F}_{x}Y(k)$ is a Gaussian process with zero-mean and satisfy
$\mathbb{E}[Y_{k}(t)Y_{k^{\prime}}(t^{\prime})]=e^{-|t-t^{\prime}||k|^{\gamma}}\frac{|\mathbb{T}|}{2|k|^{\gamma}}\textbf{I}_{k+k^{\prime}=0}(k,k^{\prime}).$
where $|\mathbb{T}|=\int_{\mathbb{T}}1\mathrm{d}x$ and $\textbf{I}_{k+k^{\prime}=0}(k,k^{\prime})$ is the indicator function. Moreover, for any $T>0$ and $\beta<\frac{\gamma(1-2\kappa)}{2}-\frac{1}{2}$, we have $X(t),Y(t)\in C_{T}^{\kappa}W^{\beta}$, for $\kappa<\frac{1}{2}$.
Proof.
It’s easy to check the first result by noting that
$\mathbb{E}[\mathcal{F}\xi(t,k){\mathcal{F}\xi(s,k^{\prime})}]=\delta(t-s)\textbf{I}_{k+k^{\prime}=0}(k,k^{\prime}).$ (2.4)
Considering the second result, we can rewrite $Y(t)$ as
$Y(t)=\sum_{k\in\mathbb{Z}}Y_{k}(t)e^{ikx}.$
Calculate $\mathbb{E}|\Delta_{j}(Y(t)-Y(t^{\prime}))|^{2}$ and $\mathbb{E}\|\Delta_{j}(Y(t)-Y(t^{\prime}))\|_{L^{2}}^{2p}$ as follows
$\mathbb{E}|\Delta_{j}(Y(t)-Y(t^{\prime}))|^{2}=$ $\mathbb{E}\sum_{k,k^{\prime}\in\mathbb{Z}}\rho_{j}(k)\rho_{j}(k^{\prime})[Y_{k}(t)-Y_{k}(t^{\prime})][Y_{-k^{\prime}}(t)-Y_{-k^{\prime}}(t^{\prime})]e^{i(k-k^{\prime})x}$
$=$ $\sum_{k,k^{\prime}\in\mathbb{Z}}\rho_{j}(k)\rho_{j}(k^{\prime})|\mathbb{T}|\frac{2-2e^{-|t-t^{\prime}||k|^{\gamma}}}{2|k|^{\gamma}}$
$\lesssim$ $\sum_{k\in\mathrm{supp}(\rho_{j})}k^{-\gamma(1-\delta)}|t-t^{\prime}|^{\delta}$
$\lesssim$ $2^{j(1-\gamma(1-\delta))}|t-t^{\prime}|^{\delta},$
for some $\delta\in[0,1]$. By Gaussian hypercontractivity [Jan97], we have
$\mathbb{E}|\Delta_{j}(Y(t,x_{i})-Y(t^{\prime},x_{i}))|^{2p}\lesssim(\mathbb{E}|\Delta_{j}(Y(t,x_{i})-Y(t^{\prime},x_{i}))|^{2})^{p}.$
It implies that
$\mathbb{E}\|\Delta_{j}(Y(t)-Y(t^{\prime}))\|_{L^{2}}^{2p}=$ $\mathbb{E}(\int_{\mathbb{T}}|\Delta_{j}(Y(t)-Y(t^{\prime}))|^{2}\mathrm{d}x)^{p}$
$\lesssim$ $\int_{\mathbb{T}^{\otimes p}}\mathbb{E}\prod_{i=1}^{p}|\Delta_{j}(Y(t,x_{i})-Y(t^{\prime},x_{i}))|^{2}\mathrm{d}x_{i}$
$\lesssim$ $\int_{\mathbb{T}^{\otimes p}}\prod_{i=1}^{p}(\mathbb{E}|\Delta_{j}(Y(t,x_{i})-Y(t^{\prime},x_{i}))|^{2p})^{\frac{1}{p}}\mathrm{d}x_{i}$
$\lesssim$ $\int_{\mathbb{T}^{\otimes p}}\prod_{i=1}^{p}(\mathbb{E}|\Delta_{j}(Y(t,x_{i})-Y(t^{\prime},x_{i}))|^{2})\mathrm{d}x_{i}$
$\lesssim$ $2^{pj(1-\gamma(1-\delta))}|t-t^{\prime}|^{\delta p}.$
Then we can estimate $\mathbb{E}\|Y(t)-Y(t^{\prime})\|_{H^{\alpha}}^{2p}$ as
$\mathbb{E}\|Y(t)-Y(t^{\prime})\|_{B_{2,2}^{\beta}}^{2p}\lesssim$ $(\sum_{j\geq-1}2^{2j\beta}\mathbb{(}E\|\Delta_{j}(Y(t)-Y(t^{\prime}))\|_{L^{2}}^{2p})^{\frac{1}{p}})^{p}$
$\leq$ $(\sum_{j\geq-1}2^{j(2\beta+1-\gamma(1-\delta))})^{p}|t-t^{\prime}|^{\delta p}<\infty,$
for $\beta<\frac{\gamma(1-\delta)-1}{2}$. By Lemma 2.6, we have $Y(t)\in C^{\frac{\delta}{2}-\frac{1}{2p}}_{T}H^{\frac{\gamma(1-\delta)-1}{2}}$. Choose $p$ large enough, we can prove the case of $H^{s}$. For Hölder space, we can also have the results by similar argument
$\mathbb{E}\|Y(t)-Y(t^{\prime})\|_{B_{2p,2p}^{\beta}}^{2p}=$ $\sum_{j\geq-1}2^{2j\beta p}\mathbb{E}\|\Delta_{j}(Y(t)-Y(t^{\prime}))\|_{L^{2p}}^{2p}$
$\lesssim$ $\sum_{j\geq-1}2^{2j\beta p}\|\mathbb{E}|\Delta_{j}(Y(t)-Y(s))|^{2}\|_{L^{p}}^{p}$
$\lesssim$ $\sum_{j\geq-1}2^{jp(2\beta+1-\gamma(1-\delta))}|t-t^{\prime}|^{\delta p}.$
Let $p$ large enough and by embedding for Besov space, we get $Y(t)\in C^{\frac{\delta}{2}-\frac{1}{2p}}_{T}\mathcal{C}^{\frac{\gamma(1-\delta)-1}{2}}$. Choosing $p$ large enough again, we can prove the case of $\mathcal{C}^{s}$, which we finish our proof. ∎
Remark 2.1.
By Lemma 2.7, we have $X(t),Y(t)\in C_{T}^{\kappa}W^{\beta}$ for $\frac{\alpha}{\gamma}<\frac{1}{2}-\frac{1}{\gamma}-\kappa$. Let $\beta=0$, then $\kappa$ can choose to be $\frac{1}{2}-\frac{1}{2\gamma}-\epsilon$ for any small $\epsilon>0$. Then we have $X(t),Y(t)\in\mathcal{W}^{\alpha}(T)$ a.s for any $T\geq 0$ and $\alpha<\frac{\gamma}{2}-\frac{1}{2}$ if we fix $\epsilon$ small enough.
2.3 Iteration argument
Next we introduce a generalized contraction mapping theorem which compared with Lemma 5.5($L=0$) in [BCD11].
Lemma 2.8.
Let $E$ be a Banach space, and $\mathcal{B}$ a continuous bilinear map from $E\times E$ to $E$, and $r$ a positive real number such that
$r<\frac{1}{8\|\mathcal{B}\|}\quad with\quad\|\mathcal{B}\|_{E}=\sup_{\|u\|_{E},\|v\|_{E}\leq 1}\|\mathcal{B}(u,v)\|_{E}.$
Let $L$ is a linear map $E$ to $E$ and satisfied
$\sup_{v\in E}\frac{\|L(v)\|_{E}}{\|v\|_{E}}\leq\frac{1}{4}.$
Then for any $a$ in a ball $B(0,r)\subset E$, there exists a unique $\tilde{x}$ in $\mathcal{B}(0,2r)$ such that
[…]
Proof.
The proof of first result can refer to [BCD11]. For second, we involved an application of iterative scheme defined by
$\tilde{x}_{0}=a\quad\tilde{x}_{n+1}=a+L(\tilde{x}_{n})+\mathcal{B}(\tilde{x}_{n},\tilde{x}_{n}).$
By induction argument, we show that $\|\tilde{x}_{n+1}\|\leq 2r$ when $\|\tilde{x}_{n}\|\leq 2r$. In fact, by our assumption, we get
$\|\tilde{x}_{n+1}\|_{E}\leq r+\frac{r}{4}+4\|\mathcal{B}\|r^{2}\leq 2r.$
Thus, $(\tilde{x}_{n})_{n\in\mathbb{N}}$ remains in $B(0,2r)$. Moreover,
$\|\tilde{x}_{n+1}-\tilde{x}_{n}\|_{E}\leq$ $\|L(\tilde{x}_{n}-\tilde{x}_{n-1})\|_{E}+4r\|\mathcal{B}\|\|\tilde{x}_{n}-\tilde{x}_{n-1}\|_{E}$
$\leq$ $\frac{3}{4}\|\tilde{x}_{n}-\tilde{x}_{n-1}\|_{E}.$
[…]
We also provide version for studying specific types of coupled equations. Let $f((x,y))$ be a linear functional on $E\times E$ which means
$f(c(x,y))=cf((x,y)),$ (2.5)
$f(x_{1}+x_{2},y_{1}+y_{2})=f(x_{1},y_{1})+f(x_{2},y_{2}).$
Assume that $f(x,y)$ satisfied some control property on $X$, such as
$\|f(x,y)\|_{X}\leq\|x\|_{X}+c\|y\|_{X},$ (2.6)
we have the following lemma
Lemma 2.9.
Let $\mathcal{E}:=\dot{E}\times E$ be a Banach space, where $\dot{E}$ is subspace of $E$. Define norm space $\dot{E}\times E$ as
$\|(x,y)\|_{\mathcal{E}}=2\|x\|_{\dot{E}}+\|y\|_{E}.$
Moreover define $\mathcal{B}$ a continuous bilinear map from $E\times E$ to $\dot{E}$, and $r$ a positive real number such that
$\|\mathcal{B}(f(x,y),f(x,y))\|_{\dot{E}}\leq\frac{1}{8r}(\|x_{n}\|_{\dot{E}}+c\|y_{n}\|_{E})^{2},$
where $f$ satisfied (2.5) and (2.6) for $X=E$. Let $\mathcal{G}(x)$ and $g(y)$ are linear map $E$ to $\dot{E}$ and satisfied
$\sup_{v\in E}(\frac{\|\mathcal{G}(v)\|_{\dot{E}}}{\|v\|_{E}},\frac{\|g(v)\|_{\dot{E}}}{\|v\|_{E}})\leq\frac{1}{16},\quad\frac{\|h(v)\|_{\dot{E}}}{\|v\|_{\dot{E}}}\leq\frac{1}{16}.$
Then there exist $\tilde{c}>0$, for any $(a,b)\in B(0,r)\subset\mathcal{E}$ and $0<c<\tilde{c}$, there exists unique $(y,x)\in B(0,4r)\subset\mathcal{E}$ such that
$y=b+f(x,y),\quad x=a+g(y)+h(x)+\mathcal{B}(f(x,y),f(x,y))+\mathcal{G}(f(x,y)).$ (2.7)
Proof.
We involved application of iterative scheme defined by
$x_{0}=a,\quad y_{0}=b,\quad y_{n+1}=b+f(x_{n},y_{n}),\quad x_{n+1}=a+g(y_{n})+h(x_{n})+\mathcal{G}(f(x_{n},y_{n}))+\mathcal{B}(f(x_{n},y_{n}),f(x_{n},y_{n})).$
We first prove that $(x_{n+1},y_{n+1})\in B(0,4r)\subset E$ when $(x_{n},y_{n})\in B(0,4r)\subset E$. In fact,
$\|(x_{n+1},y_{n+1})\|_{\dot{E}\times E}\leq$ $\|b\|_{E}+\|f(x_{n},y_{n})\|_{E}+2\|a\|_{\dot{E}}+2\|g(y_{n})\|_{\dot{E}}+2\|h(x_{n})\|_{\dot{E}}$
$+2\|\mathcal{G}(f(x_{n},y_{n}))\|_{\dot{E}}+2\|\mathcal{B}(f(x_{n},y_{n}),f(x_{n},y_{n}))\|_{\dot{E}}$
$\leq$ $2\|a\|_{\dot{E}}+\|b\|_{E}+\frac{1}{8}\|y_{n}\|_{E}+\frac{1}{8}\|x_{n}\|_{\dot{E}}+\frac{1}{8}\|f(x_{n},y_{n})\|_{E}+\frac{1}{4r}(\|x_{n}\|_{\dot{E}}+c\|y_{n}\|_{E})^{2}$
$\leq$ $2\|a\|_{\dot{E}}+\|b\|_{E}+\frac{1}{8}\|y_{n}\|_{E}+\frac{1}{8}\|x_{n}\|_{\dot{E}}+\frac{1}{8}(\|x_{n}\|_{E}+c\|y_{n}\|_{E})+\frac{1}{4r}(\|x_{n}\|_{E}+c\|y_{n}\|_{E})^{2}$
$\leq$ $r+\frac{r}{2}+\frac{r}{2}+\frac{r}{4}+\frac{1}{8}(2r+4cr)+\frac{1}{8r}(4r^{2}+16cr^{2}+16c^{2}r^{2})<4r,$
for $0<c<\tilde{c}=\frac{1}{10}$. Denote $f_{n,m}=f(x_{n},y_{n})-f(x_{m},y_{m})\in E$, to prove $\{(x_{n},y_{n})\}_{n\in\mathbb{N}}$ is a Cauchy sequence, we estimate that
$\|(x_{n+1},y_{n+1})-(x_{n},y_{n})\|_{\dot{E}\times E}\leq$ $2\|x_{n+1}-x_{n}\|_{\dot{E}}+\|y_{n+1}-y_{n}\|_{E}$
$\leq$ $2\|\mathcal{G}(f_{n,n-1})\|_{\dot{E}}+2\|\mathcal{B}(f_{n,n-1},f(x_{n},y_{n}))+\mathcal{B}(f(x_{n},y_{n}),f_{n,n-1})\|_{\dot{E}}$
$+\|f_{n,n-1}\|_{E}$
$\leq$ $(\frac{7}{4}+c)\|x_{n}-x_{n-1}\|_{\dot{E}}+(c^{2}+\frac{13}{8}c+\frac{1}{8})\|y_{n}-y_{n-1}\|_{E}$
$<$ $\frac{37}{40}\|(x_{n},y_{n})-(x_{n-1},y_{n-1})\|_{\dot{E}\times E}.$
Then we get the limit of $(y_{n},x_{n})$ is a unique fixed point of mapping (2.7) in Banach space $\mathcal{E}$, which prove the lemma. ∎
3 Result in the range of $\gamma\in(\frac{4}{3},2]$
In this section, we consider the highly dissipative regime, which ensures that the linear component $X(t)$ possesses enhanced regularity properties. In fact, assume that $\gamma\in(\frac{4}{3},2]$, by Lemma 2.7, we have $X\in C_{T}H^{\frac{1}{6}+\delta}$ for some small $\delta$. It is subcritical since $X(t)\in H^{\frac{3}{2}-\gamma}.$
3.1 Bi-linear estimate
First we build a standard bilinearity estimate. Define $B(f,g)$ as following
$B(f,g)(t)=\int_{0}^{t}e^{(t-s)\Lambda^{\gamma}}\frac{1}{2}D(f(s)g(s))\mathrm{d}s,$ (3.1)
then we have following estimate.
Lemma 3.1.
Fix $E_{T}:=L^{\infty}_{T}H^{s}$, with $0\leq\rho<\gamma-1$, $s=\sigma+\rho$, then we have
$\|B(f,g)\|_{E_{T}}\leq CT^{\frac{\gamma-1-\rho}{\gamma}}(1+T^{\frac{\rho}{\gamma}})\|fg\|_{H^{\sigma}}.$ (3.2)
Proof.
We write that
$\|B(f,g)(t)\|_{H^{s}}\lesssim$ $\int_{0}^{t}\|e^{-(t-\tau)\partial_{x}^{\gamma}}(1-\partial_{x}^{2})^{\frac{\rho}{2}}\partial_{x}(1-\partial_{x}^{2})^{\frac{\sigma}{2}}(fg)\|_{L^{2}}.$
We divide $\xi$ in two parts: $|\xi|\lesssim 1$ and $|\xi|\gtrsim 1$. For low frequencies $|\xi|\lesssim 1$, we have
$|e^{-(t-\tau)|\xi|^{\gamma}}(1+\xi^{2})^{\frac{\rho}{2}}\xi|\lesssim|e^{-(t-\tau)|\xi|^{\gamma}}|\xi||,$
For high frequencies $|\xi|\gtrsim 1$, we have
$|e^{-(t-\tau)|\xi|^{\gamma}}(1+\xi^{2})^{\frac{\rho}{2}}\xi|\lesssim|e^{-(t-\tau)|\xi|^{\gamma}}|\xi|^{\rho+1}|,$
Then we obtain that
$\|B(f,g)(t)\|_{H^{s}}\lesssim\int_{0}^{t}\max(\frac{1}{(t-\tau)^{\frac{1}{\gamma}}},\frac{1}{(t-\tau)^{\frac{1+\rho}{\gamma}}})\|fg\|_{H^{\sigma}}\lesssim T^{\frac{\gamma-1-\rho}{\gamma}}(1+T^{\frac{\rho}{\gamma}})\|fg\|_{H^{\sigma}},$
which finish our proof. ∎
Remark 3.1.
The case of end-point $\rho=\gamma-1$ will be more complex. Using the time-frequencies analysis by frequencies localization, we need the estimate in $B_{2,\infty}^{s}$. In this paper, we don’t touch this endpoint. Since for any $\alpha<\frac{\gamma}{2}-\frac{1}{2}$, we can always have $X\in\mathcal{W}^{\alpha+}(T)$ a.s.
By Moser-type estimate, its not difficult to prove that
Lemma 3.2.
Let $\gamma\in(\frac{4}{3},2]$, $\alpha\in(\frac{1}{6},\frac{1}{2})$, fix $2\delta=\alpha+\gamma-\frac{3}{2}$, then for $T\leq 1$ we have following bounds
  • (1) For $f,g\in L^{\infty}_{T}H_{x}^{\frac{1}{2}+\delta}$, we have
    $\|B(f,g)\|_{L^{\infty}_{T}H^{\frac{1}{2}+\delta}}\leq C_{1}T^{\frac{\gamma-1}{\gamma}}\|f\|_{L^{\infty}_{T}H^{\frac{1}{2}+\delta}}\|g\|_{L^{\infty}_{T}H^{\frac{1}{2}+\delta}}.$
  • (2) For $f,g\in L^{\infty}_{T}W^{\alpha}$ for any $p\in[2,\infty]$, we have
    $\|B(f,g)\|_{L^{\infty}_{T}H^{\frac{1}{2}+\delta}}\leq C_{2}T^{\frac{\delta}{\gamma}}\|f\|_{C_{T}W^{\alpha}}\|g\|_{C_{T}W^{\alpha}}.$
  • (3) For $f\in L_{T}^{\infty}{H^{\frac{1}{2}+\delta}}$ and $g\in L_{T}^{\infty}W^{\alpha}$ for any $p\in[2,\infty]$, then we have
    $\|B(f,g)\|_{L^{\infty}_{T}H^{\frac{1}{2}+\delta}}\leq C_{3}T^{\frac{\delta}{\gamma}}\|f\|_{L^{\infty}_{T}H^{\frac{1}{2}+\delta}}\|g\|_{L^{\infty}_{T}H^{\alpha}}.$
3.2 Well-posedness
In this little section, we study the well-posedness of (1.1) for $\gamma>\frac{4}{3}$. Firstly, we consider the equation
$\partial_{t}v-\Lambda^{\gamma}v=\frac{1}{2}\partial_{x}(v^{2})+\partial_{x}(v\tilde{X})+\frac{1}{2}\partial_{x}(\tilde{X}^{2}),\quad u(0)=u_{0}.$ (3.3)
We solve the equation (3.3) by contraction mapping argument.
Proposition 3.1.
Let $\alpha\in(\frac{1}{6},\frac{1}{2})$, $\gamma\in(\frac{4}{3},2]$ and $u_{0}\in H^{\frac{1}{2}+\delta}$ for $2\delta=\alpha+\gamma-\frac{3}{2}$. If $\tilde{X}(t)\in{C_{T}W^{\alpha}}$, then there exists a unique solution $v(t,x)$ of equation (3.3) with a postive $T^{*}$ satisfied $v\in C_{T^{*}}H^{\frac{1}{2}+\delta}$.
Proof.
Let $r=2\max(\|u_{0}\|_{H^{\alpha}},\|\tilde{X}\|_{C_{1}W^{\alpha}})$. We rewrite equation (3.3) as
$u(t,x)=P(t)u_{0}+B(v,\tilde{X})+B(\tilde{X},v)+B(\tilde{X},\tilde{X})+B(v,v).$
where $P(t)f$ satisfied linear equation
$\partial_{t}P(t)f-\Lambda^{\gamma}P(t)f=0,\quad P(0)f=f.$ (3.4)
By Lemma 2.8, we claim that there exist $T^{*}$ such that
$a:=P(T^{*})u_{0}+B(\tilde{X},\tilde{X})\in B(0,r),\quad\|B\|_{C_{T^{*}}H^{\frac{1}{2}+\delta}}\leq\frac{1}{8r},\quad\sup_{v}\frac{2\|B(v,\tilde{X})\|_{C_{T^{*}}H^{\frac{1}{2}+\delta}}}{\|v\|_{C_{T^{*}}H^{\frac{1}{2}+\delta}}}\leq\frac{1}{4}.$ (3.5)
For the first claim of (3.5), fix $T>0$ small enough to make sure that $C_{2}T^{\frac{\delta}{\gamma}}\|\tilde{X}\|^{2}_{C_{1}W^{\alpha}}\leq\frac{r}{4}$, then by Lemma 3.2, we have
$\|a\|_{C_{T}H^{\frac{1}{2}+\delta}}\leq$ $\|P(\cdot)u_{0}\|_{C_{T}H^{\frac{1}{2}+\delta}}+\|B(\tilde{X},\tilde{X})\|_{C_{T}H^{\frac{1}{2}+\delta}}$
$\leq$ $\|u_{0}\|_{C_{T}H^{\frac{1}{2}+\delta}}+C_{2}T^{\frac{\delta}{\gamma}}\|\tilde{X}\|^{2}_{C_{1}W^{\alpha}}\leq r.$
For the second claim of (3.5), fix $T>0$ small enough to make sure that $C_{1}T^{\frac{\gamma-1}{\gamma}}\leq\frac{1}{8r}$, then by Lemma 3.2, we have
$\|B(v,v)\|_{L^{\infty}_{T}H^{\frac{1}{2}+\delta}}\leq$ $C_{1}T^{\frac{\gamma-1}{\gamma}}\|v\|_{C_{T}H^{\frac{1}{2}+\delta}}\|v\|_{C_{T}H^{\frac{1}{2}+\delta}}$
$\leq$ $\frac{1}{8r}\|v\|_{C_{T}H^{\frac{1}{2}+\delta}}^{2}$
For the third claim of (3.5), fix $T$ small enough to make sure that $C_{3}T^{\frac{\delta}{\gamma}}\|\tilde{X}\|_{C_{1}W^{\alpha}}\leq\frac{1}{8}$, then by Lemma 3.2, we have
$\frac{\|B(v,\tilde{X})\|_{C_{T}H^{s}}}{\|v\|_{C_{T}H^{s}}}\leq$ $C_{3}T^{\frac{\delta}{\gamma}}\|\tilde{X}\|_{C_{1}W^{\alpha}}\leq\frac{1}{8}.$
Fix $T^{*}$ satisfied the condition above all, we finish our proof. ∎
Proof of Theorem 1.1.
By Lemma 2.7, we have $X(t)\in C_{T}W^{\alpha}$ where $X(t)$ satisfied (1.5), by Proposition 3.1, we then we finish our proof by writing equation (1.1) as the (3.3) with fixing v=uXv=u-X and $\tilde{X}(t)=X(t)=\int_{0}^{t}P(t-s)\xi(s)\mathrm{d}s$. ∎
Remark 3.2.
We actually prove that the solution exists in the space $L_{t}^{\infty}H^{s}$ but depending on the equation itself, we can also easily demonstrate the continuity of $v$. Note that
$v(t)-v(s)=(S(t)-S(s))u_{0}+B_{s,t}(v+X,v+X)$
where $B_{s,t}(u,u):=\frac{1}{2}\int_{s}^{t}P(t-s)D(u^{2})\mathrm{d}s$. Use the similar argument for proving well-posedness, we can easily get the continuous.
To obtain the global result, we need more information of energy. Let us recall a classical energy estimate for general heat equation in [BCD11] in nonhomogeneous case and a energy control for Burgers equation.
Lemma 3.3.
Let $\gamma\in(1,2]$. Let $v$ be the solution in $\mathcal{C}([0,T];\mathcal{S}^{\prime}(\mathbb{T}^{d}))$ of the Cauchy problem
[…]
with $f\in L^{2}([0,T];\dot{H}^{s-1})$ and $v_{0}\in\dot{H}^{s}(\mathbb{T}^{d})$. Then,
$v\in(\cap_{p=2}^{\infty}L^{p}([0,T];\dot{H}^{s+\frac{\gamma}{p}}))\cap\mathcal{C}([0,T];\dot{H}^{s}).$
Moreover we have the following estimates:
$\|v(t)\|_{\dot{H}^{s}}^{2}+2\int_{0}^{t}\|\Lambda^{\frac{\gamma}{2}}v(t^{\prime})\|_{\dot{H}^{s}}^{2}\mathrm{d}t^{\prime}=$ $\|v_{0}\|_{\dot{H}^{s}}^{2}+2\int_{0}^{t}\langle f(t^{\prime}),v(t^{\prime})\rangle_{s}\mathrm{d}t^{\prime},$
$(\int_{\mathbb{T}^{d}}\langle\xi\rangle^{2s}(\sup_{0\leq t^{\prime}\leq t}|\hat{v}(t^{\prime},\xi)|)^{2}\mathrm{d}\xi)^{\frac{1}{2}}\leq$ $\|v_{0}\|_{\dot{H}^{s}}+\|f\|_{L_{T}^{2}(\dot{H}^{s-\frac{\gamma}{2}})},$
$\|v(t)\|_{L_{T}^{p}(\dot{H}^{s+\frac{\gamma}{p}})}\leq$ $\|v_{0}\|_{\dot{H}^{s}}+\|f\|_{L_{T}^{2}(\dot{H}^{s-\frac{\gamma}{2}})},$
with $\langle a,b\rangle_{s}=\int\langle\xi\rangle^{2s}\hat{a}(\xi)\overline{\hat{b}(\xi)}\mathrm{d}\xi$
Proof.
The case of $\gamma=2$ is proved in [BCD11], the extension to the case $\gamma\in(1,2]$ is straightforward. ∎
Proposition 3.2.
For $\gamma\in(\frac{3}{2},2]$, let $v$ be the smooth solution of (1.1) on $[0,T]$ for $u_{0}\in H^{\frac{1}{2}+\delta}$ and $X(t)$ satisfied (1.5). Then we have energy estimate as follows
$\|v\|_{L^{2}}^{2}+(1-\nu)\int_{0}^{t}\|v\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}\tau\lesssim\|u_{0}\|_{L^{2}}^{2}e^{C_{\nu}(1+\|X\|_{C_{t}W^{\alpha}}^{\frac{\gamma}{\gamma-1}})t}+C_{\nu}\int_{0}^{t}e^{C_{\nu}(1+\|X\|_{C_{t}W^{\alpha}}^{\frac{\gamma}{\gamma-1}})(t-t^{\prime})}\|X\|_{W^{\alpha}}^{4}\mathrm{d}t^{\prime}.$ (3.7)
Proof.
Note that $v$ satisfied (3.6), where $f=\frac{1}{2}\partial_{x}(v+X)^{2}$. By Lemma 3.3 and integral by parts, we have
$\|v\|_{L^{2}}^{2}+\int_{0}^{t}\|v\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}\leq\|u_{0}\|_{L^{2}}^{2}+2\int_{0}^{t}\|u\|_{L^{2}}^{2}\mathrm{d}t^{\prime}+\int_{0}^{t}\langle v,2\partial_{x}(vX)+\partial_{x}(X^{2})\rangle_{0}\mathrm{d}t^{\prime}.$
By Hölder inequality and Lemma 2.1 and Lemma 2.7, there exist $\alpha\in(\frac{1}{4},\frac{1}{2})$ such that $\frac{\gamma}{2}-\frac{1}{2}>\alpha>1-\frac{\gamma}{2}$ satisfied
$\langle v,2\partial_{x}(vX)+\partial_{x}(X^{2})\rangle_{0}\leq$ $2\|v\|_{H^{\frac{\gamma}{2}}}\|vX\|_{H^{1-\frac{\gamma}{2}}}+\|v\|_{H^{\frac{\gamma}{2}}}\|X^{2}\|_{H^{1-\frac{\gamma}{2}}}$
$\lesssim$ $2\|v\|_{H^{\frac{\gamma}{2}}}^{\frac{2}{\gamma}}\|v\|_{L^{2}}^{2-\frac{2}{\gamma}}\|X\|_{\mathcal{C}^{\alpha}}+\|v\|_{H^{\frac{\gamma}{2}}}\|X\|_{W^{\alpha}}^{2},$
where we use the interpolation $\|f\|_{H^{1-\frac{\gamma}{2}}}\leq C\|f\|_{L^{2}}^{\frac{2\gamma-2}{\gamma}}\|f\|_{H^{\frac{\gamma}{2}}}^{\frac{2-\gamma}{\gamma}}$. Using Young’s inequality, we have
$\|v\|_{L^{2}}^{2}+\int_{0}^{t}\|v\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}\leq\|u_{0}\|_{L^{2}}^{2}+\nu\int_{0}^{t}\|v\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}+C_{\nu}\int_{0}^{t}(1+\|X\|_{W^{\alpha}}^{\frac{\gamma}{\gamma-1}})\|v\|_{L^{2}}^{2}\mathrm{d}t^{\prime}+\frac{2}{\nu}\|X\|_{C_{T}W^{\alpha}}^{4}t.$ (3.8)
By Gronwall’s inequality, the proposition is proved. ∎
Moreover, next lemma note that we can improve such $v$ into higher regularity space.
Lemma 3.4.
Under the assumption in Lemma 3.3, for $v\in C_{T}L^{2}\cap L^{2}_{T}H^{\frac{\gamma}{2}}$, we have energy estimate as follows
$\|\Lambda^{s}v\|_{L^{2}}+2(1-\nu)\int_{0}^{t}\|\Lambda^{s}v\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}\leq\|u_{0}\|^{2}_{L^{2}}e^{C_{\nu}\int_{0}^{t}\|v\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}}+\int_{0}^{t}e^{C_{\nu}\int_{t^{\prime}}^{t}\|v\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime\prime}}(\|X\|_{W^{\alpha}}^{2}\|v\|_{H^{\frac{\gamma}{2}}}^{2}+\|X\|_{W^{\alpha}}^{4})\mathrm{d}t^{\prime}$ (3.9)
for some small $\nu<1$ and any $0<s<\gamma-\frac{3}{2}$. Moreover, $v\in C_{T}H^{s}\cap L_{T}^{2}H^{s+\frac{\gamma}{2}}$.
Proof.
Recall $\Lambda^{s}v$ satisfied following equation
$\partial_{t}(\Lambda^{s}v)-\Lambda^{\gamma}(\Lambda^{s}v)=\frac{1}{2}\Lambda^{s}\partial_{x}(v^{2})+\Lambda^{s}\partial_{x}(vx)+\frac{1}{2}\Lambda^{s}\partial_{x}(X^{2}).$ (3.10)
Multiply $\Lambda^{s}v$ both sides and integral, we have
$\frac{1}{2}\|\Lambda^{s}v\|_{L^{2}}^{2}+\int_{0}^{t}\|\Lambda^{s}v\|_{\dot{H}^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}=\frac{1}{2}\langle\Lambda^{s}\partial_{x}(v^{2}),\Lambda^{s}v\rangle+\langle\Lambda^{s}\partial_{x}(vX),\Lambda^{s}v\rangle+\frac{1}{2}\langle\Lambda^{s}\partial_{x}(X^{2}),\Lambda^{s}v\rangle.$ (3.11)
For the first term in the right side of (3.11), by Parsavel formula, we have
$\frac{1}{2}\langle\Lambda^{s}\partial_{x}(v^{2}),\Lambda^{s}v\rangle\leq$ $\|v\|_{\dot{H}^{s+\frac{\gamma}{2}}}\|D^{{s+1-\frac{\gamma}{2}}}v\|_{L^{2}}$
$\leq$ $\|v\|_{\dot{H}^{s+\frac{\gamma}{2}}}\|D^{{s+1-\frac{\gamma}{2}}}v-2vD^{s+1-\frac{\gamma}{2}}\|_{L^{2}}+\|v\|_{\dot{H}^{s+\frac{\gamma}{2}}}\|2vD^{s+1-\frac{\gamma}{2}}\|_{L^{2}}.$
By the estimate in [KPV93, GO14] and the interpolation, we have
$\|D^{{s+1-\frac{\gamma}{2}}}v-2vD^{s+1-\frac{\gamma}{2}}\|_{L^{2}}\lesssim$ $\|D^{\alpha}v\|_{L^{4}}^{2}$
$\lesssim$ $\|v\|_{\dot{H}^{s}}\|v\|_{\dot{H}^{\frac{\gamma}{2}}},$
for $\gamma\geq\frac{3}{2}$. By Hölder’s inequality and Sobolev embedding in [BCD11], for $0<s+1-\frac{\gamma}{2}<1$ and $\alpha=\frac{s}{2}+\frac{1}{2}-\frac{\gamma}{4}$, we have
$\|2vD^{s+1-\frac{\gamma}{2}}\|_{L^{2}}\lesssim$ $2\|v\|_{L^{\frac{2}{1-2s}}}\|D^{s+1-\frac{\gamma}{2}}v\|_{L^{\frac{2}{2s}}}$
$\lesssim$ $2\|v\|_{\dot{H}^{s}}\|D^{s+1-\frac{\gamma}{2}}v\|_{\dot{H}^{\frac{1}{2}-s}}$
$\lesssim$ $2\|v\|_{\dot{H}^{s}}\|v\|_{\dot{H}^{\frac{3}{2}-\frac{\gamma}{2}}}$
$\lesssim$ $2\|v\|_{\dot{H}^{s}}\|v\|_{\dot{H}^{\frac{\gamma}{2}}},$
for $s<\frac{1}{2}$ and $\gamma>\frac{3}{2}$. Then we have
$\frac{1}{2}\langle\Lambda^{s}\partial_{x}(v^{2}),\Lambda^{s}v\rangle\leq\frac{\nu}{3}\|v\|_{\dot{H}^{s+\frac{\gamma}{2}}}^{2}+C_{\nu}\|v\|_{\dot{H}^{s}}^{2}\|v\|_{\dot{H}^{\frac{\gamma}{2}}}^{2}.$
For the second term and third term, by Hölder’s inequality and Lemma 2.1, we have
$\langle\Lambda^{s}\partial_{x}(vX),\Lambda^{s}v\rangle\leq$ $\|vX\|_{\dot{H}^{s+1-\frac{\gamma}{2}}}\|v\|_{\dot{H}^{s+\frac{\gamma}{2}}}$
$\leq$ $\|X\|_{W^{\alpha}}\|v\|_{\dot{H}^{s+1-\frac{\gamma}{2}}}\|v\|_{\dot{H}^{s+\frac{\gamma}{2}}}$
$\leq$ $\|X\|_{W^{\alpha}}\|v\|_{\dot{H}^{\frac{\gamma}{2}}}\|v\|_{\dot{H}^{s+\frac{\gamma}{2}}}$
$\leq$ $C_{\nu}\|X\|_{W^{\alpha}}^{2}\|v\|^{2}_{\dot{H}^{\frac{\gamma}{2}}}+\frac{\nu}{3}\|v\|_{\dot{H}^{s+\frac{\gamma}{2}}}^{2},$
and
$\frac{1}{2}\langle\Lambda^{s}\partial_{x}(X^{2}),\Lambda^{s}v\rangle\leq$ $\|vX\|_{\dot{H}^{s+1-\frac{\gamma}{2}}}\|v\|_{\dot{H}^{s+\frac{\gamma}{2}}}$
$\leq$ $C_{\nu}\|X\|_{W^{\alpha}}^{2}\|v\|_{\dot{H}^{s+\frac{\gamma}{2}}}$
$\leq$ $C_{\nu}\|X\|_{W^{\alpha}}^{4}+\frac{\nu}{3}\|v\|_{\dot{H}^{s+\frac{\gamma}{2}}}^{2}.$
Summary above all, we have
$\frac{1}{2}\|\Lambda^{s}v\|_{L^{2}}^{2}+(1-\nu)\int_{0}^{t}\|\Lambda^{s}v\|_{\dot{H}^{\frac{\gamma}{2}}}\leq$ $C_{\nu}\int_{0}^{t}\|X\|_{W^{\alpha}}^{2}\|v\|_{\dot{H}^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}+C_{\nu}\int_{0}^{t}\|X\|_{W^{\alpha}}^{4}\mathrm{d}t^{\prime}+C_{\nu}\int_{0}^{t}\|v\|_{\dot{H}^{s}}^{2}\|v\|_{\dot{H}^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}$
since $\int_{0}^{t}\|v\|_{\dot{H}^{\frac{\gamma}{2}}}\mathrm{d}t^{\prime}$ is uniformly bounded by by constant only dependent on $\|u_{0}\|_{L^{2}}$ and $\|X\|_{C_{T}W^{\alpha}}$. By Gronwall’s inequality and $H^{s}\subset\dot{H}^{s}$ for $s\geq 0$, we prove (3.9) for all $v\in C_{T}L^{2}\cap L_{T}^{2}H^{\frac{\gamma}{2}}$. ∎
Finally, we give a proof of Theorem 1.2
Proof of Theorem 1.2.
Firstly, we claim that there exist $v$ satisfied equation (3.3) and energy estimate (3.7). In fact, choosing a function $\phi:\mathbb{R}\mapsto\mathbb{R}_{+}$, that is even, smooth, compactly supported, decreasing on $\mathbb{R}_{+}$, and such that $\phi(0)=1$, and then setting
$\xi^{\epsilon}(x)=\mathcal{F}^{-1}(\phi(\epsilon k)\mathcal{F}(\xi)(k))(x),\quad u_{0}^{\epsilon}(x)=\mathcal{F}^{-1}(\phi(\epsilon k)\mathcal{F}(u_{0})(k))(x).$
And $X^{\epsilon}$ solve the equation
$\partial_{t}X^{\epsilon}-\Lambda^{\gamma}X^{\epsilon}=\xi^{\epsilon},\quad X^{\epsilon}(0)=0.$
Let $v^{\epsilon}$ be the solution of equation
$\partial_{t}v^{\epsilon}-\Lambda^{\gamma}v^{\epsilon}=\frac{1}{2}\partial_{x}(v^{\epsilon})^{2}+\partial_{x}(v^{\epsilon}X^{\epsilon})+\frac{1}{2}\partial_{x}(X^{\epsilon})^{2},\quad v^{\epsilon}(0)=u_{0}^{\epsilon}.$ (3.12)
Then $v^{\epsilon}$ satisfied Proposition 3.2. Without loss of generality, let $\|X^{\epsilon}\|_{C_{T}W^{\alpha}}\leq\|X\|_{C_{T}W^{\alpha}}$ and $\|u^{\epsilon}_{0}\|_{H^{s}}\leq\|u_{0}\|_{H^{s}}$, we get that $v^{\epsilon}$ is uniformly bounded in $C_{T}L^{2}$. Then by Fatou’s lemma and Proposition 3.2, there exists $\psi:\mathbb{R}_{+}\mapsto\mathbb{R}_{+}$ and a subsequence $v_{\phi(\epsilon)}$ satisfied
$\lim_{\epsilon\to 0}\psi(\epsilon)\to 0,\quad\lim_{\epsilon\to 0}v_{\phi(\epsilon)}\to v\quad in\>S^{\prime},\quad\|v\|_{C_{T}L^{2}}^{2}+\int_{0}^{t}\|v\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}s\leq C(\|u_{0}\|_{L^{2}},\|X\|_{C_{T}W^{\alpha}}).$
To prove $v$ is the solution of (3.3), we need the strong convergence of $v_{\phi(\epsilon)}$ in $L^{2}$ to make sure that $\partial_{x}(v_{\phi(\epsilon)})^{2}$ can be weak-converge to $\partial_{x}(v^{2})$(its already valid for $\partial_{x}(X^{\epsilon})^{2}$ and $\partial_{x}(X^{\epsilon}v^{\epsilon})$). We use the Aubin-Lions-Simon Theorem in [Sim86]. It suffice to show following two conditions
$\partial_{t}v\in L_{T}^{p}(H^{-\sigma}),\quad and\quad v\in C_{T}H^{s},$ (3.13)
be satisfied for some $\sigma,s>0$, $p>1$. Note that $\partial_{t}v=\Lambda^{\gamma}v+\frac{1}{2}\partial_{x}(v^{2})+\partial_{x}(vX)+\frac{1}{2}\partial_{x}(X^{2})$. By Lemma 2.1 and Proposition 3.2, we have $\partial_{t}v\in L_{T}^{2}H^{-\sigma}$ for any $\sigma\geq 1$. The second condition of (3.13) is a consequence of (3.4). Then we prove $v$ is the solution of (3.3) which prove our claim.
To obtain the global solution, we employ induction argument. Fix $T^{*}\in[0,\infty)$, let $T^{(0)}\leq T^{*}$ small enough satisfied right hand side of (3.15) less than $\|u_{0}\|_{L^{2}}$($\tilde{T}$ can be only dependent on $\|X\|_{C_{T^{*}}W^{\alpha}}$). Consider equation which is begin at $v(\frac{T^{(0)}}{2})$. Fix $v^{(1)}=v(\frac{T^{(0)}}{2}+t)$, $X^{(1)}=X(\frac{T^{(0)}}{2}+t)$ then $v^{(1)}(t)$ satisfied the equation
$\partial_{t}v^{(1)}(t)-\Lambda^{\gamma}v^{(1)}(t)=-\partial_{x}(\frac{1}{2}(v^{(1)}+X^{(1)})^{2}),\quad v^{(1)}(0)=v(\frac{T^{(0)}}{2}).$ (3.14)
then we can get a solution $v^{(1)}$ of equation on $\frac{T^{(0)}}{2},\frac{3T^{(0)}}{2}]$ since $\|v(\frac{T}{2})\|_{L^{2}}\leq\|u_{0}\|_{L^{2}}<c$ and
$\|X^{(1)}\|_{C_{T}(W^{\frac{1}{4}-\kappa})}=\|X\|_{C_{[\frac{T^{(0)}}{2},\frac{T^{(0)}}{2}+T]}W^{\frac{1}{4}-\kappa}}\lesssim\|X\|^{2}_{C_{T^{*}}W^{\frac{1}{4}-\kappa}}.$
$v^{(1)}$ on $\frac{T^{(0)}}{2},\frac{3T^{(0)}}{2}]$ is also a solution of (3.3). Then define $v^{new}:=v^{(0)}$ on $0,\frac{T^{(0)}}{2}]$ and $v^{new}:=v^{(1)}$ on $\frac{T^{(0)}}{2},\frac{3T^{(0)}}{2}]$. Then $v^{new}$ is a solution in $C([0,\frac{3T^{(0)}}{2}];L^{2}(\mathbb{T}))\cap L^{2}([0,\frac{3T^{(0)}}{2}];H^{\frac{\gamma}{2}}(\mathbb{T}))$, and $\|v^{new}(\frac{3T^{(0)}}{2})\|_{L^{2}}\leq\|v(\frac{T^{(0)}}{2})\|_{L^{2}}<c$, repeat the above steps, we get $v\in C([0,T^{*}];L^{2}(\mathbb{T}))\cap L^{2}([0,T^{*}];H^{\frac{\gamma}{2}}(\mathbb{T}))$. Since $T^{*}$ is arbitrary, we prove that $v$ is the weak global solution of (3.3), let $X$ satisfied (1.5), we prove that $v$ is the weak global solution of (1.1).
Finally, we prove that for $\gamma>\frac{5}{3}$, $v$ can be the strong global solution of (1.1). In fact, its suffice to show the uniqueness of $v$. Let $u_{0}^{1}$, $u_{0}^{2}$ belongs to $H^{\frac{1}{2}+\delta}$ for $\delta$ satisfied Proposition 3.1, and $v(t;u_{0})$ be the solution of (3.3) with $v(0)=u_{0}$. Let $v^{1}=v(t;u_{0}^{1})$, $v^{2}=v(t;u_{0}^{2})$ and $\overline{v}:=v(t;u^{1}_{0})-v(t;u^{2}_{0})$, then $\overline{v}$ satisfied the equation
$\partial_{t}\overline{v}-\Lambda^{\gamma}\overline{v}=\frac{1}{2}\partial_{x}(\overline{v}(v^{1}+v^{2}))+\partial_{x}(\overline{v}X).$
Similar as the proof of Proposition 3.2, we obtain the estimates
$\|\overline{v}\|_{L^{2}}^{2}+\int_{0}^{t}\|\overline{v}\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}\leq 2\int_{0}^{t}\|\overline{v}\|_{L^{2}}^{2}\mathrm{d}t^{\prime}+\int_{0}^{t}\langle\overline{v},\partial_{x}(\overline{v}(v^{1}+v^{2}))+2\partial_{x}(\overline{v}X)\rangle_{0}\mathrm{d}t^{\prime}.$
By Hölder’s inequality, Parsavel formula and Lemma 2.1, for $\gamma>\frac{5}{3}$, we have
$\langle\overline{v},\partial_{x}(\overline{v}(v^{1}+v^{2}))\rangle_{0}\leq\|\overline{v}\|_{H^{\frac{\gamma}{2}}}\|\overline{v}(v^{1}+v^{2})\|_{H^{1-\frac{\gamma}{2}}}\leq\|\overline{v}\|_{H^{\frac{\gamma}{2}}}^{2}\|(v^{1}+v^{2})\|_{H^{s}}$
for some $\frac{\gamma}{2}-1>s>1-\frac{\gamma}{2}$ which can be choose by Lemma 3.4. Similarly,
$\langle\overline{v},2\partial_{x}(\overline{v}X))\rangle_{0}\leq\|\overline{v}\|_{H^{\frac{\gamma}{2}}}^{2}\|X\|_{W^{\alpha}}.$
By Gronwall lemma, we get $\|\overline{v}\|_{L^{2}}^{2}+\int_{0}^{t}\|\overline{v}\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}=0$, which finish our proof. ∎
3.3 Application to DP equation
This section demonstrates the application of the aforementioned theory to the study of a class of shallow water wave equations. The only difference between it and the Burgers equation lies on $\frac{3}{2}(1-\partial_{x}^{2})^{-1}\partial_{x}(u^{2})$. Note that the operator $(1-\partial_{x}^{2})^{-1}$ is an ‘stable’ operator in $H^{s}$ since
$\|(1-\partial_{x}^{2})^{-1}f\|_{H^{s}}\leq\|f\|_{H^{s-2}},\quad for\>s\in\mathbb{R}.$
Hence, we can easily get the local result for DP equation which finish the proof of Theorem 1.1 by incorporating Proposition 3.1.
For the global result, we need following energy estimate for DP equation
Proposition 3.3.
For $\gamma\in(\frac{3}{2},2]$, let $v$ be a smooth solution of (1.12) on $[0,T]$ for $u_{0}\in H^{\frac{1}{2}+\delta}$ and $X(t)$ satisfied (1.5), then we have energy estimate as follows
$\|v\|_{L^{2}}^{2}+(2-2\nu)\int_{0}^{t}\|v\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}\tau\lesssim\|u_{0}\|_{L^{2}}^{2}e^{(1+C_{\nu}\|X\|_{C_{T}H^{\alpha}}^{\frac{\gamma}{\gamma-1}}+\|X\|_{C_{T}W^{\alpha}})t}+(1+C_{\nu})\|X\|_{W^{\alpha}}^{4}.$ (3.15)
Proof.
For the case $t\in[\delta,T]$, let $n=(1-\partial_{x}^{2})v$ and $w=(4-\partial_{x}^{2})^{-1}v$, similar as the proof of Lemma 4.1 in [CZG24], we have
$\int_{\mathbb{R}}v\partial_{x}nw+3\partial_{x}vnw\mathrm{d}x=0.$
Then we can derive to following equation
$\partial_{t}\int_{\mathbb{R}}nw\mathrm{d}x=$ $2\int_{\mathbb{R}}\langle\partial_{x}\rangle^{2}\partial_{t}vw\mathrm{d}x$
$=$ $-2\int_{\mathbb{R}}\langle\partial_{x}\rangle^{2}\Lambda^{\gamma}vw\mathrm{d}x-2\int_{\mathbb{R}}\langle\partial_{x}\rangle^{2}\partial_{x}(2vX+X^{2})w\mathrm{d}x-3\int_{\mathbb{R}}\partial_{x}(2vX+X^{2})w\mathrm{d}x.$
By Parseval formula, Hölder inequality and interpolation inequality, we have
$\int_{\mathbb{T}}\langle\partial_{x}\rangle^{2}\partial_{x}(vX)w\mathrm{d}x=$ $\int_{\mathbb{T}}\frac{1+|\xi|^{2}}{4+|\xi|^{2}}\mathcal{F}(\partial_{x}(vX))\mathcal{F}(v)\mathrm{d}x$
$\lesssim$ $\|v\|_{H^{\frac{\gamma}{2}}}\|vX\|_{H^{1-\frac{\gamma}{2}}}$
$\lesssim$ $\|v\|_{H^{\frac{\gamma}{2}}}^{\frac{2}{\gamma}}\|v\|_{L^{2}}^{2-\frac{2}{\gamma}}\|X\|_{\mathcal{C}^{\alpha}}$
$\lesssim$ $\frac{\nu}{3}\|v\|_{H^{\frac{\gamma}{2}}}^{2}+C_{\nu}\|v\|_{L^{2}}^{2}\|X\|_{W^{\alpha}}^{\frac{\gamma}{\gamma-1}},$
where we use the fact $\alpha>{1-\frac{\gamma}{2}}$. Similarly, use Lemma 2.1, we have
$\int_{\mathbb{T}}\langle\partial_{x}\rangle^{2}\partial_{x}X^{2}w\mathrm{d}x\lesssim$ $\frac{\nu}{3}\|v\|_{H^{\frac{\gamma}{2}}}^{2}+C_{\nu}\|X\|_{W^{\alpha}}^{4},$
$\int_{\mathbb{T}}\partial_{x}(2vX+X^{2})w\mathrm{d}x\lesssim$ $\frac{\nu}{3}\|v\|_{H^{\frac{\gamma}{2}}}^{2}+C_{\nu}\|v\|_{L^{2}}^{2}\|X\|_{W^{\alpha}}^{\frac{\gamma}{\gamma-1}}+C_{\nu}\|X\|_{W^{\alpha}}^{4}.$
Summarize all, by Gronwall’s inequality, we finish our proof. ∎
Similar as the proof of Lemma 3.4, we can improve the estimate in higher regularity as following, which we omit the proof.
Lemma 3.5.
Under the assumption in Proposition 3.3, for $v\in C_{T}L^{2}\cap L^{2}_{T}H^{\frac{\gamma}{2}}$, we have energy estimate as follows
$\|\Lambda^{s}v\|_{L^{2}}+2(1-\nu)\int_{0}^{t}\|\Lambda^{s}v\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}\leq$ $\|u_{0}\|^{2}_{L^{2}}e^{C_{\nu}\int_{0}^{t}\|v\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}}$ (3.16)
$+\int_{0}^{t}e^{C_{\nu}\int_{t^{\prime}}^{t}\|v\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime\prime}}(\|X\|_{W^{\alpha}}^{2}\|v\|_{H^{\frac{\gamma}{2}}}^{2}+\|X\|_{W^{\alpha}}^{4})\mathrm{d}t^{\prime}$
Using the similar argument in the proof of Theorem 1.2, we finish the proof of Remark 1.3.
4 Result in the range of $\gamma\in(\frac{5}{4},\frac{4}{3}]$
4.1 Regularity analysis
In this section, we will investigate how to obtain a remainder term with Sobolev regularity $\frac{1}{2}+$ under weaker dissipation conditions $\gamma\in(\frac{5}{4},\frac{4}{3}]$. We will employ the paracontrolled method, which was introduced in [GIP15, GST19]. This method is particularly powerful for analyzing the regularity structure of solutions or studying singular SPDEs.
Our strategy as follows. Recall our regularity analysis and difference equation
$\partial_{t}v-\Lambda^{\gamma}v=\mathcal{L}(v)=$ $\mathcal{L}(u-X)$
$=$ $\frac{1}{2}(u^{2})+\xi-\xi$
$=$ $\frac{1}{2}\partial_{x}(v^{2})+\frac{1}{2}\partial_{x}(X^{2})+\partial_{x}(vX).$
Assume that $v$ gains the critical regularity $H^{\frac{1}{2}+}$. By the critical analysis, $\partial_{x}(v^{2})$, $\partial_{x}(X^{2})$ and $\partial_{x}(vX)$ must gain $H^{\frac{1}{2}-\gamma+}$-order regularity at least. But its impossible in the weaker dissipation conditions $\gamma\in(\frac{5}{4},\frac{4}{3}]$ since $\partial_{x}(X^{2})$ and $\partial_{x}(vX)$ can gain $\alpha-1$-order regularity at most. To overcome this difficult, our strategy is to decompose such terms in “ higher-regularity-term" and “ lower- regularity-term" by para-product decomposition (1.9). Correspondingly, we express $v$ as $u^{\sharp}$ and $w$, where $\mathcal{L}u^{\sharp}$ is equal to “ higher-regularity-term" and $\mathcal{L}w$ can counter the “ lower- regularity-term".
We firstly assume that $v$ also belong to $\mathcal{C}^{0+}$. We introduce a method called "para-controlled solution" which developed in [GIP15, GP17] for some part of $u$ notation by $u^{\sharp}$ gaining $H^{\frac{1}{2}+}$ regularity. To get a higher order term in $H^{\frac{1}{2}+}$. Fix distribution $u^{\sharp}$ satisfied
$\mathcal{L}u^{\sharp}=\frac{1}{2}(u^{\sharp})^{2}+R(X,u^{\sharp},u),\quad u^{\sharp}(0)=u_{0}\in W^{\frac{1}{2}+}.$ (4.1)
where $R(X,u^{\sharp},u)$ is to be determined. We assume $R(X,u^{\sharp},u)\in W^{\frac{1}{2}-\gamma+}$. Then by regularity theory, $u^{\sharp}$ gain $W^{\frac{1}{2}+}$. Fix $u=X+w+u^{\sharp}$, then we have
$\mathcal{L}(u^{\sharp})=$ $\mathcal{L}(u-X-w)$
$=$ $\frac{1}{2}(u^{2})+\xi-\xi-\mathcal{L}w$
$=$ $\frac{1}{2}\partial_{x}(w^{2})+\frac{1}{2}\partial_{x}(X^{2})+\partial_{x}(wX)+\partial_{x}(wu^{\sharp})+\partial_{x}(Xu^{\sharp})+(\frac{1}{2}\partial_{x}((u^{\sharp})^{2})-\mathcal{L}w).$
Since $u^{\sharp}$ gains enough regularity, which means $u^{\sharp}\cdot w$ preserve regularity of $w$. The restricted term is $\partial_{x}(w\cdot X)$ and $\partial_{x}(X^{2})$, since they at most gains $\alpha-1$-order regularity which lower than $\frac{1}{2}-\gamma+$ in weak dissipation conditions $\gamma\in(\frac{5}{4},\frac{4}{3}]$. Using para-product decomposition (1.9) and $X\in C_{T}\mathcal{C}^{\alpha}$, we note that $\partial_{x}(X\circ X)$, $\partial_{x}(X\prec w)$ and $\partial_{x}(X\circ w)$ belong to $H^{\frac{1}{2}-\gamma}$, hence the main difficulty is term $\partial_{x}(v\prec X)$ and $\partial_{x}(X\prec X)$.
To solve this, we assume that $v$ satisfied some suitable structure to counter this difficult term. Inspired by para-controlled ansatz, we assume that $w$ as form $u^{\prime}\prec\!\!\!\prec\mathcal{Q}$ and $u^{\mathcal{Q}}:=w+u^{\sharp}$. Assume $u^{\prime}\in\mathcal{W}^{\frac{1}{8}+}(T)$, $\mathcal{Q}\in\mathcal{W}^{\frac{3}{8}+}$ and $u^{\sharp}\in\mathcal{W}^{\frac{1}{2}+}(T)$, then by Lemma 2.5, $w\in\mathcal{W}^{\frac{3}{8}+}(T)$. As $\gamma>\frac{5}{4}$, We have regularity analysis as following
$[\partial_{x}(X\prec X)-X\prec\partial_{x}X]+\frac{1}{2}\partial_{x}(w^{2})+\partial_{x}(wu^{\sharp})+\frac{1}{2}\partial_{x}(X\circ X)\in C_{T}W^{-\frac{3}{4}+},\quad\partial_{x}(X\prec u^{\mathcal{Q}})\in C_{T}W^{-\frac{5}{8}+},$
$\partial_{x}(X\circ u^{\mathcal{Q}})\in C_{T}W^{-\frac{1}{2}+},\quad[u^{\prime}\prec\!\!\!\prec\partial_{x}X-u^{\prime}\prec\partial_{x}X]\in C_{T}W^{-\frac{3}{4}+},\quad u^{\prime}\prec\!\!\!\prec\mathcal{L}\mathcal{Q}-\mathcal{L}(u^{\prime}\prec\!\!\!\prec\mathcal{Q})\in C_{T}W^{\frac{1}{2}-\gamma+}.$
Therefore we can determine $R(X,u^{\sharp},u)$ as
$R(X,u^{\sharp},u):=$ $\frac{1}{2}\partial_{x}(w^{2})+\partial_{x}(wu^{\sharp})+\partial_{x}(X\circ X)+[\partial_{x}(X\prec X)-X\prec\partial_{x}X]+X\prec\partial_{x}X$
$+\partial_{x}(X\prec u^{\mathcal{Q}}+X\circ u^{\mathcal{Q}}+[\partial_{x}(u^{\mathcal{Q}}\prec X)-u^{\mathcal{Q}}\prec\partial_{x}X]$
$+u^{\mathcal{Q}}\prec\partial_{x}X+u^{\prime}\prec\!\!\!\prec\mathcal{L}\mathcal{Q}-\mathcal{L}(u^{\prime}\prec\!\!\!\prec\mathcal{Q})+[u^{\prime}\prec\partial_{x}X-u^{\prime}\prec\!\!\!\prec\partial_{x}X]$
Here we see that the purpose of using modified paraproduct operator $f\prec\!\!\!\prec g$ is to ensure that the commutator $\mathcal{L}(u^{\prime}\prec\!\!\!\prec\mathcal{Q})-u^{\prime}\prec\!\!\!\prec\mathcal{L}\mathcal{Q}$ aligns with our regularity framework. Fix $u^{\prime}=X+u^{\mathcal{Q}}$, we counter the rough term.
$u^{\prime}\prec\!\!\!\prec\mathcal{L}Q=$ $w\prec\partial_{x}X+X\prec\partial_{x}X$
where we fix $\mathcal{L}\mathcal{Q}=\partial_{x}X$. By equation (4.1), we construct a paracontrolled solution if we can verify $u^{\sharp}$ has enough regularity. In fact, by regularity theory, for $\gamma>\frac{5}{4}$, (4.1) derive the regularity of $u^{\sharp}$ such as
$u^{\sharp}\in\mathcal{W}^{\frac{1}{2}+\delta}(T),$
which we finish our ansatz setting.
4.2 Construction of para-controlled solution
We frame this problem as solving a system of equations of $(u^{\prime},u^{\sharp})$. Firstly, we define the equation of $(u^{\prime},u^{\sharp})$ driven by $(\tilde{X},u_{0})$.
$\left\{\begin{aligned} &u^{\prime}=\tilde{X}+u^{\mathcal{Q}}\\ &\mathcal{L}u^{\sharp}=\frac{1}{2}[\partial_{x}(\tilde{X}^{2})-\tilde{X}\prec\partial_{x}\tilde{X}]+\frac{1}{2}\partial_{x}(u^{\mathcal{Q}})^{2}+[\partial_{x}(u^{\mathcal{Q}}\tilde{X})-u^{\mathcal{Q}}\prec\partial_{x}\tilde{X}]+[u^{\prime}\prec\partial_{x}\tilde{X}-\mathcal{L}(u^{\prime}\prec\!\!\!\prec\mathcal{Q})],\end{aligned}\right$. (4.2)
with $u^{\sharp}(0)=u_{0}$.
Proposition 4.1.
Let $\alpha\in(\frac{1}{8},\frac{1}{6})$, $\gamma\in(\frac{4}{3},2]$ and $u_{0}\in W^{\frac{1}{2}+\delta}$ for $\delta\leq\min(\gamma-\frac{5}{4},\alpha-\frac{1}{8})$. If $\tilde{X}\in\mathcal{W}^{\alpha}(T)$, then there exists a unique solution $(u^{\prime},u^{\sharp})$ of equation (4.2) with positive $T^{*}$ satisfied $(u^{\prime},u^{\sharp})\in\mathcal{W}^{\frac{1}{8}+\delta}(T^{*})\times\mathcal{W}^{\frac{1}{2}+\delta}(T^{*})$.
Proof.
Firstly, fix $u^{\sharp}(0)=u_{0}$ and $f(u^{\sharp},u^{\prime})=u^{\mathcal{Q}}=u^{\prime}\prec\!\!\!\prec Q+u^{\sharp}$, rewrite (4.2) as following
$u^{\prime}=\tilde{X}+u^{\mathcal{Q}},\quad\quad u^{\sharp}=a+g(u^{\prime})+\mathcal{B}(f(u^{\sharp},u^{\prime}),f(u^{\sharp},u^{\prime}))+\mathcal{G}(f(u^{\sharp},u^{\prime})).$
And
$\mathcal{L}a=P(t)u_{0}+\frac{1}{2}\partial_{x}(\tilde{X}^{2})-\tilde{X}\prec\partial_{x}\tilde{X}$, $\quad a(0)=0,$
$\mathcal{L}(\mathcal{G}(f(u^{\sharp},u^{\prime})))=[\partial_{x}(u^{\mathcal{Q}}\tilde{X})-u^{\mathcal{Q}}\prec\partial_{x}\tilde{X}$], $\quad\mathcal{G}(f(u^{\sharp},u^{\prime}))(0)=0,$
$\mathcal{L}{\mathcal{B}}(f(u^{\sharp},u^{\prime}),f(u^{\sharp},u^{\prime}))=\frac{1}{2}\partial_{x}(u^{\mathcal{Q}})^{2},$ $\quad{\mathcal{B}}(f(u^{\sharp},u^{\prime})(0)=0,$
$\mathcal{L}g(u^{\prime})=[u^{\prime}\prec\!\!\!\prec\partial_{x}\tilde{X}-\mathcal{L}(u^{\prime}\prec\!\!\!\prec\mathcal{Q})]+[u^{\prime}\prec\partial_{x}\tilde{X}-u^{\prime}\prec\!\!\!\prec\partial_{x}\tilde{X}$], $\quad g(u^{\prime})(0)=0.$
Define $\mathcal{E}_{T}$ as following
$\mathcal{E}_{T}:=\{(u,v)\in\mathcal{W}^{\frac{1}{2}+\delta}(T)\times\mathcal{W}^{\frac{1}{8}+\delta}(T):2\|u\|_{\mathcal{W}^{\frac{1}{2}+\delta}(T)}+\|v\|_{\mathcal{W}^{\frac{1}{8}+\delta}(T)}<\infty\}.$
Our goal is to find a $T^{*}\leq 1$ such that the conditions of Lemma 2.9 hold in $\mathcal{E}_{T^{*}}$. Firstly, by Lemma 2.5, for $\epsilon=\frac{1}{8}+\delta$ and $\rho=\gamma-\delta$, we have
$\|f(u^{\sharp},u^{\prime})\|_{\mathcal{W}^{\frac{1}{8}+\delta}(T)}\leq$ $\|u^{\sharp}\|_{\mathcal{W}^{\frac{1}{2}+\delta}(T)}+CT^{\frac{\delta}{\gamma}}\|u^{\prime}\|_{\mathcal{W}^{\frac{1}{8}+\delta}(T)}(\|\mathcal{Q}\|_{C_{T}\mathcal{C}^{\frac{1}{8}+\delta}}+\|\partial_{x}\tilde{X}\|_{C_{T}\mathcal{C}^{\frac{1}{8}-\gamma+2\delta}}),$
$\leq$ $\|u^{\sharp}\|_{\mathcal{W}^{\frac{1}{2}+\delta}(T)}+CT^{\frac{\delta}{\gamma}}\|u^{\prime}\|_{\mathcal{W}^{\frac{1}{8}+\delta}(T)}\|\tilde{X}\|_{\mathcal{W}^{\alpha}(T^{*})}.$
and
$\|f(u^{\sharp},u^{\prime})\|_{C_{T}W^{\frac{3}{2}-\gamma+\delta}}\leq$ $\|u^{\sharp}\|_{\mathcal{W}^{\frac{1}{2}+\delta}(T)}+CT^{\frac{\delta}{\gamma}}\|u^{\prime}\|_{\mathcal{W}^{\frac{1}{8}+\delta}(T)}(\|\mathcal{Q}\|_{C_{T}W^{\frac{3}{2}-\gamma+\delta}}+\|\partial_{x}\tilde{X}\|_{\frac{3}{2}-2\gamma+2\delta})$
$\leq$ $\|u^{\sharp}\|_{\mathcal{W}^{\frac{1}{2}+\delta}(T)}+CT^{\frac{\delta}{\gamma}}\|u^{\prime}\|_{\mathcal{W}^{\frac{1}{8}+\delta}(T)}\|\tilde{X}\|_{\mathcal{W}^{\alpha}(T^{*})}.$
fix $T$ small enough, we can make sure that $CT^{1-\frac{\delta}{\gamma}}\|\tilde{X}\|_{\mathcal{W}^{\frac{1}{8}+\delta}(T^{*})}<\frac{1}{10}$.
Fix $r=4C\|u_{0}\|_{W^{\frac{1}{2}+\delta}}$, by Lemma A.3 and Lemma B.1, we can choose $T$ small enough such that
$\|(\tilde{X},a)\|_{\mathcal{E}_{T}}\leq$ $\|\tilde{X}\|_{\mathcal{W}^{\alpha}(T)}+2C\|u_{0}\|_{W^{\frac{1}{2}+\delta}}+2CT^{\frac{\delta}{\gamma}}\|\mathcal{L}a\|_{W^{\frac{1}{2}-\gamma+2\delta}}$
$\leq$ $\frac{r}{4}+\frac{r}{2}+\frac{r}{4}\leq r,$
by Lemma B.3 and Lemma B.1, we have similar as above argument, we can choose $T$ small enough such that
$\|g(u^{\prime})\|_{\mathcal{W}^{\frac{1}{2}+\delta}}\leq CT^{\frac{\delta}{\gamma}}\|\mathcal{L}g(u^{\prime})\|_{\mathcal{W}^{\frac{1}{2}-\gamma+2\delta}}\leq C$ $T^{\frac{\delta}{\gamma}}\|u^{\prime}\|_{C_{{T}}\mathcal{C}^{\frac{1}{8}+\delta}}(\|\mathcal{Q}\|_{C_{{T}}W^{\frac{3}{8}+\delta}}+\|\partial_{x}\tilde{X}\|_{C_{T}W^{\frac{3}{8}-\gamma+\delta}})$
$\leq$ $CT^{\frac{\delta}{\gamma}}\|u^{\prime}\|_{\mathcal{W}^{\frac{1}{8}+\delta}(T)}\|\tilde{X}\|_{C_{{T}}W^{\alpha}}\leq\frac{1}{16}\|u^{\prime}\|_{\mathcal{W}^{\frac{1}{8}+\delta}(T)}.$
and
$\|\mathcal{G}(f(u^{\sharp},u^{\prime}))\|_{\mathcal{W}^{\frac{1}{2}+\delta}}\leq$ $CT^{\frac{\delta}{\gamma}}\|\mathcal{L}\mathcal{G}(f(u^{\prime},u^{\sharp}))\|_{\mathcal{W}^{\frac{1}{2}-\gamma+2\delta}}$
$\leq$ $CT^{\frac{\delta}{\gamma}}\|f(u^{\prime},u^{\sharp})\|_{\mathcal{W}^{\frac{1}{8}+\delta}}\|\tilde{X}\|_{\mathcal{W}^{\alpha}(T)}$
$\leq$ $\frac{1}{16}\|f(u^{\prime},u^{\sharp})\|_{\mathcal{W}^{\frac{1}{8}+\delta}}.$
where we use the fact $\gamma>\frac{3}{2}$. Finally, We can choose $T$ small enough such that
$\|\mathcal{B}(f(u^{\sharp},u^{\prime}),f(u^{\sharp},u^{\prime}))\|_{\mathcal{W}^{\frac{1}{2}+\delta}}\leq$ $CT^{\frac{\delta}{\gamma}}\|\mathcal{L}\mathcal{B}(f(u^{\prime},u^{\sharp}))\|_{C_{T}W^{\frac{1}{2}-\gamma+2\delta}}$
$\leq$ $CT^{\frac{\delta}{\gamma}}\|\partial_{x}(f(u^{\prime},u^{\sharp})^{2})\|_{C_{T}W^{\frac{1}{2}+\delta}}$
$\leq$ $CT^{\frac{\delta}{\gamma}}\|f(u^{\prime},u^{\sharp})^{2}\|_{C_{T}W^{\frac{3}{2}-\gamma+\delta}}$
$\leq$ $CT^{\frac{\delta}{\gamma}}\|f(u^{\prime},u^{\sharp})\|^{2}_{C_{T}W^{\frac{3}{2}-\gamma+\delta}}$
$\leq$ $\frac{1}{8r}(\|u^{\sharp}\|_{\mathcal{W}^{\frac{1}{2}+\delta}}+c\|y_{n}\|_{\mathcal{W}^{\frac{1}{8}+\delta}})^{2}.$
Fix $T^{*}$ satisfied above all, by Lemma 2.9, there exists $(u^{\sharp},u^{\prime})\in\mathcal{E}_{T^{*}}$ where we prove our proposition. ∎
Proof of Theorem 1.3.
Choose $\tilde{X}(t)=X(t)$ satisfied (1.5). By the Remark 2.1, we have $\tilde{X}\in\mathcal{W}^{\alpha}({T^{*}})$ for some $\alpha\in(\frac{1}{8},\frac{1}{6})$. Then by Proposition 4.1, there exists $(u^{\prime},u^{\sharp})\in C_{T^{*}}W^{\frac{1}{8}+\delta}\times C_{T^{*}}W^{\frac{1}{2}+\delta}$ satisfied (4.2) for some $\delta>0$. Its not difficult to check that $u=X+u^{\prime}\prec\!\!\!\prec\mathcal{Q}+u^{\sharp}$ is the solution of (1.1) by the analysis in the Section 4.1. Since $u\in C_{T^{*}}W^{\alpha}$, we finish the proof of Theorem 1.3. ∎
5 Discussion on the range of $\gamma$
In the final section, we give a conjecture on the range of $\gamma$. Consider the rough of singular Burgers equation driven by $|D|^{\beta}\xi,$
$u_{t}-\Lambda^{\gamma}u=\frac{1}{2}\partial_{x}(u^{2})+|D|^{\beta}\xi.$ (5.1)
where $\beta\geq 0$. Fix $\zeta=|D|^{\beta}\xi$, For dyadic number $N\neq 0$, fix $\zeta^{N}:=\mathcal{F}^{-1}(1_{|k|\sim N}\mathcal{F}\xi(k))$, which means the frequency of $\xi$ is focus on $\frac{1}{2}N\leq|k|\leq N$. Then the Fourier mode of the linear evolution $X:=\int_{0}^{t}e^{(t-s)\Lambda^{\gamma}}\zeta^{N}\mathrm{d}s$ can be written as
$X=N^{\beta-\frac{\gamma}{2}}\sum_{|k|\sim N}G_{k}(t)e^{ikx},\quad G_{k}(t)=N^{\frac{\gamma}{2}}\int_{0}^{t}e^{-(t-s)|k|^{\gamma}}\xi_{k}(s)\mathrm{d}s,$
Its not difficult to calculate that $G_{k}(t)$ from a collection of independent Gaussian variable with $\mathbb{E}[G_{k}(t)]\sim 1$, then we can get $X$ belong to $C_{T}H^{\frac{\gamma}{2}-\beta-\frac{1}{2}}$. Now we calculate the "first nonlinear iteration", fix $u^{(1)}(t):=\int_{0}^{t}e^{(t-s)\Lambda^{\gamma}}\partial_{x}(X^{2})\mathrm{d}s$, and fix $u_{k}^{(1)}(t)$ is the Fourier mode
$u_{k}^{(1)}(t)=N^{2\beta-\gamma+1}\sum_{\begin{subarray}{c}l,m\in\mathbb{Z},|l|\sim|m|\sim|k|\sim|N|\\ l+m=k\end{subarray}}\int_{0}^{t}e^{-(t-s)|k|^{\gamma}}G_{l}(s)G_{m}(s)\mathrm{d}s.$
Since the heat integral always provide $N^{-\gamma}$and the sum only has size $N^{0}$, then the inner sum integral has size $N^{-\frac{\gamma}{2}}$ with high probability by square root cancellation in [DNY22], then we get the ansatz that
$\beta\leq\gamma-1.$ (5.2)
We call $(\beta,\gamma)$ is a "probability admissible pair" for Burgers equation, if (5.2) satisfied. We can simply get that $(0,\gamma)$ for $\gamma>\frac{4}{3}$ and $(1,2)$ are probability admissible pairs, which the first range was studied in our paper and the second in [GP17].
5.1 Future directions
In future work, we will continue our investigation of stochastic Burgers equation, but in a more singular case. For example, $\gamma=2$ and $\beta=1$ for (5.1). Note that $u$ gains the regularity as $X(t)=\int_{0}^{t}P(t-s)\xi(s)\mathrm{d}s$ has, by Lemma 2.7, we priori get $u\in-\frac{1}{2}-\delta$ for any $\delta>0$. It shows that $\partial_{x}(u^{2})$ is maybe ill-defined since the condition of Lemma 2.1 is not hold! To solve this, a common approach is to study its generalized solution. More preciesly, let $u^{\epsilon}$ satisfied
$\partial_{t}u^{\epsilon}-\Lambda^{\gamma}u^{\epsilon}=\frac{1}{2}D(u^{\epsilon})^{2}+|D|^{\beta}\xi_{\epsilon},\quad u^{\epsilon}(0)=u^{\epsilon}_{0}+Y_{\epsilon}(0),$ (5.3)
and prove the $u_{\epsilon}$ has a limit $u$ in probability sense. This idea was first proposed by Martin Hairer and applied to the KPZ equation in [Hai13]. In his subsequent work [Hai14], they established a comprehensive theory of regularity structures, which resolved the problem of defining solutions for a broad class of singular SPDEs, earning him the Fields Medal. There, they build a equation in modeled distribution space $\mathcal{D}^{\gamma,\eta}_{P}$ on $[0,T]$ such that
$u=(\mathcal{K}_{\bar{\Gamma}}+R_{\gamma}\mathcal{R})\textbf{R}^{+}F(u)+Gu_{0}.$ (5.4)
Our next target is to build a generalized equation in the classical space for example $W^{\alpha}$, and prove that the generalized solution of singular Burgers equation 5.1 can be back to the generalized equation.
Appendix A Estimate for heat flow
In this chapter, we give some estimate for heat flow. Define operator $P(t)$ as
$\mathcal{F}(P(t)f)=e^{-t|\xi|^{\gamma}}\mathcal{F}f.$ (A.1)
for $f\in\mathcal{S}^{\prime}(\mathbb{R}^{d})$. Its not difficult to see that $P(t)f$ is the solution of
$\partial_{t}u-\Delta u=0,\quad u(0)=f.$ (A.2)
For the operator, we have following estimate
Lemma A.1.
For $\gamma>0$, we have the estimate
$\|P(t)f\|_{L^{p}}\leq Ce^{-ct\lambda^{\gamma}}\|f\|_{L^{p}},$
if the support of $\hat{f}$ is belong to some $\lambda\mathcal{C}$, where $\mathcal{C}$ is a annual.
Proof.
The proof is similar as Lemma 2.4 in [BCD11], which only need to notice that $\partial_{x}^{\beta}e^{-t|\xi|^{\gamma}}\leq C(1+t)^{|\beta|}e^{-ct}$ for $\xi\in\lambda\mathcal{C}$ will not cause difficult when $\gamma>0$, since $\xi$ take away from 0. ∎
Following this useful Lemma, we can get
Lemma A.2.
Let $p,q\in[1,\infty]$, $\beta\in\mathbb{Z}_{+}$, then for $\gamma\geq 1$, we have
$\|\partial_{x}^{\beta}P(t)f\|_{L^{p}}\lesssim t^{-\frac{\beta}{\gamma}-\frac{1}{\gamma}(\frac{1}{r}-\frac{1}{p})}\|f\|_{L^{r}}.$
Proof.
We rewrite
$\partial_{x}^{\beta}P(t)f=G(t,x)\ast f,$
where $g(t,x)=(\mathcal{F}^{-1}(ik)^{\beta}e^{-t|k|^{\gamma}})$. By Young inequality, it suffice to show that for $\frac{1}{q}=1+\frac{1}{p}-\frac{1}{r}$, we have $\|G(t,x)\|_{L^{q}}\leq t^{-\frac{\beta}{\gamma}-(1-\frac{1}{q})\frac{1}{\gamma}}$. In fact, rewrite $G(t,x)$ as
$G(t,x)=t^{-\frac{\gamma}{\beta}}\mathcal{F}^{-1}(it^{\frac{1}{\gamma}}k)^{\beta}e^{-|t^{\frac{1}{\gamma}}k|^{\gamma}}.$
Since $g(k)=k^{\beta}e^{-|k|^{\gamma}}$ is belong to $\mathcal{S}(\mathbb{R})$, we get $\check{g}(x)\in\mathcal{S}(\mathbb{R})\subset L^{q}(\mathbb{R})$. By Lemma 1.24 in [BCD11], the bound
$\|G(t,x)\|_{L^{q}(\mathbb{T})}\lesssim t^{-\frac{\beta}{\gamma}-\frac{1}{\gamma}}\|\check{g}(t^{-\frac{1}{\gamma}}x)\|_{L^{q}(\mathbb{T})}\lesssim t^{-\frac{\beta}{\gamma}-(1-\frac{1}{q})\frac{1}{\gamma}}\|\check{g}(\cdot)\|_{L^{q}(\mathbb{R})}\leq t^{-\frac{\beta}{\gamma}-(1-\frac{1}{q})\frac{1}{\gamma}},$
which we prove our claim. ∎
For Besov space $B_{p,r}^{s}$, we have result
Lemma A.3.
Let $1\leq p,r\leq\infty$, $\alpha\in\mathbb{R}$ and $\gamma\geq 1$, then we have following estimate
$\|P(t)f\|_{B_{p,r}^{\alpha}}\lesssim\|f\|_{B_{p,r}^{\alpha}}.$ (A.3)
Let $\delta\geq 0$ and $t>0$, we have
$\|P(t)f\|_{B_{p,r}^{\alpha+\delta}}\leq t^{-\frac{\delta}{\gamma}}\|f\|_{B_{p,r}^{\alpha}}.$
Moreover, fix $f(0),\mathcal{L}f:=(\partial_{t}-\Lambda^{\gamma})f\in B_{p,r}^{\alpha}$ , then for $\delta\in[0,\gamma)$ and $t\leq 1$, we have estimate
$\|f\|_{B_{p,r}^{\alpha+\delta}}\lesssim\|f(0)\|_{B_{p,r}^{\alpha+\delta}}+t^{1-\frac{\delta}{\gamma}}\|\mathcal{L}f\|_{B_{p,r}^{\alpha}}.$
Proof.
For $j\geq 0$, by Lemma A.1, for $t>0$, we have
$\|P(t)\Delta_{j}f(t)\|_{L^{p}}\lesssim$ $Ce^{-ct2^{j\gamma}}\|\Delta_{j}f\|_{L^{p}}$
$\lesssim$ $t^{-\frac{\delta}{\gamma}}2^{-j\delta}\|\Delta_{j}f\|_{L^{p}}$
For j=1j=-1, by Lemma A.2, we have
$\|P(t)\Delta_{j}f(t)\|_{L^{p}}\lesssim$ $\|\Delta_{j}f(t)\|_{L^{p}}$
$\lesssim$ $(2^{-\delta}t^{\frac{\delta}{\gamma}})t^{-\frac{\delta}{\gamma}}2^{\delta}\|\Delta_{j}f(t)\|_{L^{p}}$
$\lesssim$ $t^{-\frac{\delta}{\gamma}}2^{-j\delta}\|\Delta_{j}f\|_{L^{p}}.$
Multiplying $2^{j\alpha}$ both sides and taking the $l^{r}$-norm with respect to $1\leq j$, we prove the second result.
For the last claim, by the last result
$\|f\|_{B_{p,r}^{\alpha+\delta}}\leq$ $\|P(t)f(0)\|_{B_{p,r}^{\alpha+\delta}}+\int_{0}^{t}\|P(t-s)\mathcal{L}f(s)\|_{B_{p,r}^{\alpha+\delta}}$
$\lesssim$ $\|f(0)\|_{B_{p,r}^{\alpha+\delta}}+\int_{0}^{t}\|(t-s)^{-\frac{\delta}{\gamma}}\mathcal{L}f(s)\|_{B_{p,r}^{\alpha}}$
$\lesssim$ $\|f(0)\|_{B_{p,r}^{\alpha+\delta}}+t^{1-\frac{\delta}{\gamma}}\|\mathcal{L}f(s)\|_{B_{p,r}^{\alpha}}$
where we use $\delta<\gamma$. ∎
The similar result is also valid for $C_{T}^{\alpha}L^{p}$. We have the following lemma
Lemma A.4.
Let $\gamma\geq 1$ and $\alpha\in\mathbb{R}_{+}$, for any $T\geq 0$ we have estimate
$\|P(\cdot)f\|_{\mathcal{W}^{\alpha}_{\infty}(T)}\lesssim\|f\|_{\mathcal{C}^{\alpha}},\quad\|P(\cdot)f\|_{\mathcal{W}^{\alpha}_{2}(T)}\lesssim\|f\|_{H^{\alpha}}.$
Moreover, let $\alpha\in(0,\gamma)$, for $\delta\in[0,\gamma)$ satisfied $0<\alpha+\delta<\gamma$ and $0\leq T\leq 1$, it has the estimate
$\|f\|_{\mathcal{W}_{\infty}^{\alpha+\delta}(T)}\lesssim\|f(0)\|_{W_{\infty}^{\alpha+\delta}}+T^{1-\frac{\delta}{\gamma}}\|\mathcal{L}f\|_{C_{T}W_{\infty}^{\alpha}},\quad\|f\|_{\mathcal{W}_{2}^{\alpha+\delta}(T)}\lesssim\|f(0)\|_{W_{2}^{\alpha+\delta}}+T^{1-\frac{\delta}{\gamma}}\|\mathcal{L}f\|_{C_{T}W_{2}^{\alpha}}.$
Proof.
We only proof the case of $\mathcal{W}_{2}^{\alpha}(T)$, the case of $\mathcal{W}_{\infty}^{\alpha+\delta}(T)$ can be refer to Lemma 2.9 in [GP17]. The estimate for $C_{T}H^{\alpha}$ and $C_{T}H^{\alpha+\delta}$ is a direct consequence of Lemma A.3. For $C_{T}^{\frac{\alpha}{\gamma}}L^{2}$, by Parsavel formula, we have
$\|P(\cdot)f\|_{C_{T}^{\frac{\alpha}{\gamma}}L^{2}}=$ $\sup_{0\leq s<t\leq T}\frac{\|(P(t)-P(s))f\|_{L^{2}}}{|t-s|^{\frac{\alpha}{\gamma}}}$
$\leq$ $\sup_{0\leq s<t\leq T}\frac{\|e^{-t|k|^{\gamma}}(1-e^{-(t-s)|k|^{\gamma}})\hat{f}(k)\|_{L^{2}}}{|t-s|^{\frac{\alpha}{\gamma}}}\leq\|f\|_{H^{\alpha}},$
for $\alpha\in\mathbb{R}_{+}$, which we prove the first estimate. Note that
$\|f(t)-f(s)\|_{L^{2}}\leq\int_{s}^{t}\|P(t-\tau)\mathcal{L}f(\tau)\|_{L^{2}}\mathrm{d}\tau+\int_{0}^{s}\|[P(t-\tau)-P(s-\tau)]\mathcal{L}f(\tau)\|\mathrm{d}\tau.$
The first term can be handled using a method similar to that in Lemma A.3, in fact
$\int_{s}^{t}\|P(t-\tau)\mathcal{L}f(\tau)\|_{L^{2}}\mathrm{d}\tau\lesssim$ $\|\mathcal{L}f\|_{C_{T}H^{\alpha}}\int_{s}^{t}e^{-(t-\tau)|k|^{\gamma}}|k|^{-\alpha}\mathrm{d}\tau$
$\lesssim$ $\|\mathcal{L}f\|_{C_{T}H^{\alpha}}(t-s)^{1+\frac{\alpha}{\gamma}}.$
For the second term, we estimate as
$\int_{0}^{s}\|[P(t-\tau)-P(s-\tau)]\mathcal{L}f(\tau)\|_{L^{2}}\mathrm{d}\tau\lesssim$ $\|\mathcal{L}f\|_{C_{T}H^{\alpha}}\int_{0}^{s}e^{-(s-\tau)|k|^{\gamma}}(e^{-(t-s)|k|^{\gamma}}-1)|k|^{-\alpha}\mathrm{d}s$
$\lesssim$ $\|\mathcal{L}f\|_{C_{T}H^{\alpha}}(t-s)^{\frac{\alpha+\delta}{\gamma}}s^{1-\frac{\delta}{\gamma}}.$
Summary all, from the $C_{T}^{\frac{\alpha+\delta}{\gamma}}L^{2}$, we have
$\|f\|_{\mathcal{W}_{2}^{\alpha+\delta}(T)}=$ $\sup_{0\leq s<t\leq T}\frac{\|f(t)-f(s)\|_{L^{2}}}{|t-s|^{\frac{\alpha+\delta}{\gamma}}}$
$\leq$ $T^{1-\frac{\delta}{\gamma}}\|\mathcal{L}f\|_{C_{T}H^{\alpha}}.$
Then we finish the proof. ∎
There are also some similar estimates for $\tilde{P}(t)$ which define as
$\mathcal{F}(\tilde{P}(t)(f))=\frac{e^{-t|\xi|^{\gamma}}}{1+|\xi|^{2}}\mathcal{F}f.$
To clear that, fix $\varphi(t,z)=\frac{e^{-|z|^{\gamma}}}{1+t^{-\frac{2}{\gamma}}|z|^{2}}.$ We have $\tilde{P}(t)f=\varphi(t,t^{\frac{1}{\gamma}}D)f=\tilde{G}(t,x)f$. Fixing $t\in[0,T]$, its easy to see that $\mathcal{F}^{-1}\varphi(t,z)\in L^{1}(\mathbb{R})$. In fact, by Young’s inequality
$\|\mathcal{F}^{-1}_{z}\varphi(t,z)(\xi)\|_{L^{1}}=$ $\|\mathcal{F}^{-1}{e^{-|z|^{\gamma}}}\ast\mathcal{F}(\frac{1}{1+t^{-\frac{2}{\gamma}}|z|^{2}})\|_{L^{1}}$
$\leq$ $\|\check{g}\|_{L^{1}}\|\mathcal{F}(\frac{1}{1+t^{-\frac{2}{\gamma}}|z|^{2}})\|_{L^{1}}$
the first integral is always finite. For the second term, by $\mathcal{F}^{-1}_{z}{\frac{1}{1+|z|^{2}}}=\pi e^{-|\xi|}$, we have $\mathcal{F}(\frac{1}{1+t^{-\frac{2}{\gamma}}|z|^{2}})=\pi t^{\frac{1}{\gamma}}e^{-t^{\frac{1}{\gamma}}|\xi|}$ which also in $L^{1}$, which obtain the following claim
$\|\tilde{P}(t)f\|_{B_{p,r}^{\alpha}}\lesssim\|f\|_{B_{p,r}^{\alpha}}.$
On the other hand, for $j\geq 1$, we have the estimate
$\|(\mathcal{F}^{-1}\varphi(t,t^{\frac{1}{\gamma}}z))\ast\Delta_{j}u\|_{L^{p}}\lesssim$ $\|\mathcal{F}^{-1}(\varphi(t,t^{\frac{1}{\gamma}}z)\psi(2^{-j}z))\|_{L^{1}}\|\Delta_{j}u\|_{L^{p}}$
$\lesssim$ $\|\mathcal{F}^{-1}(\varphi(t,t^{\frac{1}{\gamma}}2^{j}z)\psi)\|_{L^{1}}\|\Delta_{j}u\|_{L^{p}}$
$\lesssim$ $\|(1+|x|)^{d+1}\mathcal{F}^{-1}(\varphi(t,t^{\frac{1}{\gamma}}2^{j}z)\psi)\|_{L^{\infty}}\|\Delta_{j}u\|_{L^{p}}$
$\lesssim$ $\|(1+|D|)^{d+1}\varphi(t,t^{\frac{1}{\gamma}}2^{j}z)\psi\|_{L^{1}}\|\Delta_{j}u\|_{L^{p}}$
$\lesssim$ $(1+2^{j}t^{\frac{1}{\gamma}})^{d+1}\max_{\mu\in\mathbb{N}^{d}:|\mu|\leq d+1}\|\partial^{\mu}\varphi(t,t^{\frac{1}{\gamma}}2^{j}\cdot)\|_{L^{\infty}(\mathrm{supp}(\psi))}\|\Delta_{j}u\|_{L^{p}}.$
By Lebnitz’rule $\partial^{2}(uv)=\sum_{\mu_{1}+\mu_{2}=2}\frac{\alpha!}{\mu_{1}!\mu_{2}!}\partial^{\mu_{1}}u\partial^{\mu_{2}}v$, we have
$\partial^{2}\varphi(t,z)=\sum_{\mu_{1}+\mu_{2}=2}\frac{\alpha!}{\mu_{1}!\mu_{2}!}\partial^{\mu_{1}}(e^{-|z|^{\gamma}})\partial^{\mu_{2}}(\frac{1}{1+t^{-\frac{2}{\gamma}}|z|^{2}})=c_{1}u^{\prime\prime}v+c_{2}u^{\prime}v^{\prime}+c_{3}uv^{\prime\prime},$
where $u(z)=e^{-|z|^{\gamma}}$ and $v(z),v^{\prime}(z),v^{\prime\prime}(z)$ satisfied
$v:=\frac{1}{1+cz^{2}},\quad v^{\prime}:=\frac{-2cz}{(1+cz^{2})^{2}},\quad v^{\prime\prime}=\frac{6c^{2}z^{2}-2c}{(1+cz^{2})^{3}}.$
Since $cz^{2}>0$, it’s obviously that $v,v^{\prime},v^{\prime\prime}\leq 1$ which shows that
$\sup_{|\mu|\leq 2}\sup_{z\geq 1}(1+|z|)^{\delta+2}|\partial_{z}^{\mu}\varphi(t,z)|\lesssim 1.$
On the other hand, since all partial derivatives of $u=e^{-|z|^{\gamma}}$ is decay faster than rational function for $\gamma\geq 1$, we have
$\sup_{|\mu|\leq 2}\sup_{z\geq 1}(1+|z|)^{\delta+2}|\partial_{z}^{\mu}\varphi(t,z)|\lesssim$ $\sup_{z\geq 1}(1+|z|)^{\delta+2d}\sum_{|\mu|\leq 2}|\partial_{z}^{\mu}\varphi(t,z)|$
$\lesssim$ $c^{-1}\sup_{z\geq 1}(1+|z|)^{\delta}|(u^{\prime\prime}+u^{\prime}+u)\frac{cz^{2}}{1+cz^{2}}|$
$+c^{-1}\sup_{z\geq 1}(1+|z|)^{\delta-1}|(u^{\prime}+u)\frac{-2c^{2}z^{4}}{(1+cz^{2})^{2}}|$
$+c^{-1}\sup_{z\geq 1}(1+|z|)^{\delta-2}|u\frac{6c^{3}z^{6}-2c^{2}z^{4}}{(1+cx^{2})^{3}}|$
$\lesssim$ $c^{-1}$
Note that $c=t^{-\frac{2}{\gamma}}$ and $d=1$. Then there exists $j_{0}$ make sure that $2^{j_{0}}t^{\frac{1}{\gamma}}\geq 1$ such that
$\|\tilde{P}(t)\Delta_{j}f(t)\|_{L^{p}}\lesssim$ $(1+2^{j}t^{\frac{1}{\gamma}})^{2}(1+2^{j}t^{\frac{1}{\gamma}})^{-2-\delta}\min(t^{\frac{2}{\gamma}},1)$
$\lesssim$ $\min(t^{\frac{2}{\gamma}},1)t^{-\frac{\delta}{\gamma}}2^{-j\delta}\|\Delta_{j}f\|_{L^{p}}$
For $1\leq j<j_{0}$, we have the truth
$\|\tilde{P}(t)\Delta_{j}f(t)\|_{L^{p}}\lesssim$ $\|\Delta_{j}f(t)\|_{L^{p}}$
$\lesssim$ $(\max(t^{-\frac{2}{\gamma}},1)2^{j_{0}\delta}t^{\frac{\delta}{\gamma}})\min(t^{\frac{2}{\gamma}},1)t^{-\frac{\delta}{\gamma}}2^{-j\delta}\|\Delta_{j}f(t)\|_{L^{p}}$
$\lesssim$ $\min(t^{\frac{2}{\gamma}},1)t^{-\frac{\delta}{\gamma}}2^{-j\delta}\|\Delta_{j}f\|_{L^{p}}$
Summary above, we prove following lemma
Corollary A.1.
Let $\gamma\geq 1$, $\tilde{P}(t)$ is defined as $\mathcal{F}(\tilde{P}(t)(f))=\frac{e^{-t|\xi|^{\gamma}}}{1+|\xi|^{2}}\mathcal{F}f$. Fix some $T>0$, then for any $t\in(0,T]$, $\alpha\in\mathbb{R}$, $\delta\in(0,\gamma)$, and $u\in S^{\prime}(\mathbb{R})$, we have
$\|\tilde{P}(t)u\|_{B^{\alpha+\delta}_{p,r}}\lesssim\min(1,t^{\frac{2}{\gamma}})t^{-\frac{\delta}{\gamma}}\|u\|_{B^{\alpha}_{p,r}}.$ (A.4)
Moreover, let $\tilde{L}f=(\partial_{t}-\Lambda^{\gamma})(1-\partial_{x}^{2})^{-1}f\in B_{p,r}^{\alpha}$, then for $\delta\in[0,\gamma+2)$ and $t\leq 1$, we have estimate
$\|f\|_{B_{p,r}^{\alpha+\delta}}\leq\|f(0)\|_{B_{p,r}^{\alpha+\delta}}+Ct^{1-\frac{\delta-2}{\gamma}}\|\mathcal{L}f\|_{B_{p,r}^{\alpha}}.$ (A.5)
Appendix B Commutator estimates
First we introduce a useful commutator estimates as Lemma 2.99 in [BCD11].
Lemma B.1.
[[BCD11]] Let $f$ be a smooth function on $\mathbb{R}^{d}$ and be homogeneous of degree $m$ away from 0. Let $\rho\in(0,1)$, $s\in\mathbb{R}$ and $p,r\in[1,\infty]$. If $p_{1},p_{2}\in[1,\infty]$ such that $\frac{1}{p_{1}}+\frac{1}{p_{2}}=\frac{1}{p}$, then the bound
$\|a\prec f(D)u-f(D)(a\prec u)\|_{B_{p,r}^{s-m+\rho}}\lesssim\|\nabla a\|_{B_{p_{1},\infty}^{\rho-1}}\|u\|_{B_{p_{2},r}^{s}}.$
Remark B.1.
Fix $f(x)=|x|^{\gamma}$, then we have $D(u\prec v)-u\prec Dv\in W_{p}^{\alpha+\beta-\gamma}$ for $u\in\mathcal{C}^{\alpha}$, $\alpha\in(0,1)$ and $v\in W_{p}^{\beta}$.
Following Lemma is proved
Lemma B.2.
Assume that $\alpha\in(0,1)$, $\beta,\gamma\in\mathbb{R}$ and satisfied $\beta+\gamma<0$, $\alpha+\beta+\gamma>0$, define $C(f,g,h)=(f\prec g)\circ h-f(g\circ h)$, then there is estimate that
$\|C(f,g,h)\|_{H^{\alpha+\beta+\gamma}}\leq\|f\|_{\mathcal{C}^{\alpha}}\|g\|_{H^{\beta}}\|v\|_{\mathcal{C}^{\gamma}}.$
Proof.
We rewrite
$C(f,g,h)=\sum_{j,k\geq-1}\sum_{|i-j|\leq 1}[1_{i\gtrsim k}R_{i}(\Delta_{k}f,g)\Delta_{j}h-1_{i\leq k-N}\Delta_{k}f\Delta_{j}g\Delta_{j}h].$
Its suffice to estimate above two terms. For any fixed $k$, we have estimate
$\|\sum_{j\geq-1}\sum_{|i-j|\leq 1}1_{i\leq k-N}\Delta_{k}f\Delta_{j}g\Delta_{j}h\|_{L^{2}}\lesssim$ $2^{-k\alpha}\|f\|_{\mathcal{C}^{\alpha}}\sum_{i=-1}^{k-N}2^{-i(\beta+\gamma)}\|g\|_{H^{\beta}}\|h\|_{\mathcal{C}^{\gamma}}$
$\lesssim$ $2^{-k(\alpha+\beta+\gamma)}\|f\|_{\mathcal{C}^{\alpha}}\|g\|_{H^{\beta}}\|h\|_{\mathcal{C}^{\gamma}},$
where we use the fact $\beta+\gamma<0$. Note that the Fourier transform of $\sum_{j,k\geq-1}\sum_{|i-j|\leq 1}1_{i\gtrsim k}R_{i}(\Delta_{k}f,g)\Delta_{j}h$ is supported in a ball $2^{k}\mathcal{B}$, by Lemma 2.84 in [BCD11], we obtain the estimate since $\alpha+\beta+\gamma>0$. For the first series, we can check that the Fourier transform of $\sum_{k\geq-1}\sum_{|i-j|\leq 1}1_{i\gtrsim k}R_{i}(\Delta_{k}f,g)\Delta_{j}h$ is supported in a ball $2^{j}\mathcal{B}$ for fixed $j$. And we calculate that
$\|\sum_{k\geq-1}\sum_{|i-j|\leq 1}1_{i\gtrsim k}R_{i}(\Delta_{k}f,g)\Delta_{j}h\|_{L^{2}}=$ $\|\sum_{|i-j|\leq 1}R_{i}(\sum_{1\leq k\lesssim i}\Delta_{k}f,g)\Delta_{j}h\|_{L^{2}}$
$\lesssim$ $\sum_{|i-j|\leq 1}2^{-i(\beta+\gamma)}\|\sum_{k\lesssim i}\Delta_{k}f\|_{L^{\infty}}\|g\|_{H^{\beta}}2^{-j\gamma}\|h\|_{\mathcal{C}^{\gamma}}$
$\lesssim$ $2^{-j(\alpha+\beta+\gamma)}\|f\|_{\mathcal{C}^{\alpha}}\|g\|_{H^{\beta}}\|h\|_{\mathcal{C}^{\gamma}}.$
Similar as above estimate, we finish our proof. ∎
Lemma B.3.
Let $\alpha\in(0,\gamma)$ and $\beta\in\mathbb{R}$. Then
$\|f\prec\!\!\!\prec g-f\prec g\|_{C_{T}W^{\alpha+\beta}}\lesssim\|f\|_{\mathcal{W}_{\infty}^{\alpha}(T)}\|g\|_{C_{T}W^{\beta}}.$ (B.1)
Moreover, define operator $\mathcal{L}=\partial_{t}-\Lambda^{\gamma}$, then for $\gamma\geq 1$, it has
$\|\mathcal{L}(f\prec\!\!\!\prec g)-f\prec\!\!\!\prec(\mathcal{L}g)\|_{C_{T}W^{\alpha+\beta-\gamma}}\lesssim\|f\|_{C_{T}\mathcal{C}^{\alpha}}\|g(t)\|_{C_{T}W^{\beta}}.$ (B.2)
Proof.
We only proof the case of $H^{\alpha}$, the case of $\mathcal{C}^{\alpha}$ can be similar. Firstly, control
$\|(\int_{\mathbb{R}}2^{\gamma j}\varphi(2^{\gamma j}(t-s))S_{j-1}f(s)\mathrm{d}s-S_{j-1}f(t))\Delta_{j}g(t)\|_{L^{2}}.$
By $\int_{\mathbb{R}}\varphi(s)\mathrm{d}s=1$, we have
$\|(\int_{\mathbb{R}}2^{\gamma j}\varphi(2^{\gamma j}(t-s))S_{j-1}f(s)\mathrm{d}s$- $S_{j-1}f(t))\Delta_{j}g(t)\|_{L^{2}}$
$=$ $\|\int_{\mathbb{R}}2^{\gamma j}\varphi(2^{2j}s)S_{j-1}f(t-s)-S_{j-1}f(t)\mathrm{d}s\Delta_{j}g(t)\|_{L^{2}}$
$\lesssim$ $\|\Delta_{j}g(t)\|_{L^{2}}\int_{\mathbb{R}}2^{\gamma j}\varphi(2^{\gamma j}s)\|S_{j-1}f(t-s)-S_{j-1}f(t)\|_{L^{\infty}}\mathrm{d}s$
$\lesssim$ $\|\Delta_{j}g(t)\|_{L^{2}}\|f(t)\|_{C_{t}^{\frac{\alpha}{\gamma}}L^{\infty}}\int_{\mathbb{R}}2^{\gamma j}\varphi(2^{\gamma j}s)|s|^{\frac{\alpha}{\gamma}}\mathrm{d}s$
$\lesssim$ $2^{-j(\alpha+\beta)}(2^{j\beta}\|\Delta_{j}g(t)\|_{L^{2}})\|f(t)\|_{\mathcal{W}_{\infty}^{\alpha}(T)}.$
Multiplying $2^{-j(\alpha+\beta)}$ both sides to the above inequality and sum over $l^{2}$, we prove (B.1).
For the second result, it’s trivial for fractional Laplacian commutator by Bernstein inequality and Lemma B.1. We only estimate
$\sum_{j}\partial_{t}(\int_{0}^{t}2^{\gamma j}\varphi(2^{\gamma j}(t-s))S_{j-1}f(s)\mathrm{d}s)\Delta_{j}g(t).$
For fixd $j$ we recall that $\mathrm{supp}(\varphi)\subset\mathbb{R}_{+}$, and therefore
$\int_{0}^{t}2^{2\gamma j}\varphi^{\prime}(2^{\gamma j}(t-s))S_{j-1}f(s)\mathrm{d}s=2^{\gamma j}\int_{\mathbb{R}}2^{\gamma j}\varphi^{\prime}(2^{\gamma j}(t-s))S_{j-1}f(s)1_{s\geq 0}\mathrm{d}s.$
Since $\varphi^{\prime}(0)=\varphi(0)=0$, (B.2) then follows as the first part of the proof. ∎