### v3

# Results of Fractional Rough Burgers equation in $H^{s}$ space and its application

Abstract
In this paper, we study the well-posedness of Fractional Rough Burgers equation driven by space-time white noise in $H^{s}(\mathbb{T})$ space. For the higher dissipation $\gamma\in(\frac{4}{3},2]$, we establish local well-posedness. Global well-posedness is further obtained when $\gamma$ is restricted to the interval $(\frac{5}{3},2]$. For the lower dissipation $\gamma\in(\frac{5}{4},\frac{4}{3}]$, we use the regularity analysis derivation the para-controlled solution.
Keywords: Rough PDEs; Paracontrolled solution; Well-posedness; Shallow water.
2020 Mathematics Subject Classification: 60H15, 60H17, 42B37, 35Q35
1 Introduction
In this paper, we will consider Fractional Rough Burgers (FRB) equation with weak dissipation in $H^{s}(\mathbb{T})$
$\partial_{t}u-\Lambda^{\gamma}u=\partial_{x}(u^{2})+\xi,\quad u(0,x)=u_{0}.$ (1.1)
Here $\mathbb{T}$ is some torus satisfied $\int_{\mathbb{T}}e^{ikx}=0$ for any $k\in\mathbb{Z}/\{0\}$(for example $\mathbb{R}/2\pi\mathbb{Z}$) and $\Lambda^{\gamma}$ is a Fourier multiplier satisfied $\mathcal{F}(\Lambda^{\gamma}f)=-|\xi|^{\gamma}\mathcal{F}f$. $\xi$ is a centered Gaussian space-time random distribution with covariance
$\mathbb{E}[\xi(\omega,t,x)\xi(\omega,s,y)]=\delta_{0}(t-s)\delta_{0}(x-y),\quad t,s\geq 0,\quad x,y\in\mathbb{T},$ (1.2)
where $\delta_{0}(x)$ is a Dirac measure satisfied $\delta_{0}(x)=0$ for $x\neq 0$ and $\int_{\mathbb{T}}\mathrm{d}\delta_{0}(x)=1$.
Equation (1.1) is a classical stochastic partial differential equation driven by additive noise. This class of problems has attracted considerable attention in recent years, largely motivated by the pioneering work on regularity structures by Hairer [Hai13, Hai14] and on paracontrolled calculus by Gubinelli and Perkowski [GIP15, GP17].
The central objective of this paper is to understand how the dissipation parameter $\gamma$ affects the well-posedness and regularity of solutions for the specific quadratic nonlinearity $\partial_{x}(u^{2})$. This nonlinearity is of particular interest due to its close connection to the KPZ equation; see, for example, [Cor12]. In contrast, previous work such as [GIP15] studied the fractional stochastic Burgers equation with a general nonlinearity $F(u)\partial_{x}u$, establishing well-posedness in $\mathcal{C}^{s}$ for $s\in(\frac{1}{3},\frac{\gamma}{2}-\frac{1}{2})$ in the regime $\gamma\in(\frac{5}{3},2]$.
The behavior of equation (1.1) with lower dissipation $\gamma<\frac{5}{3}$ in the Sobolev spaces $H^{s}$ remains largely unexplored. These spaces are of significant mathematical and physical relevance in the study of related equations such as the Burgers and Degasperis–Procesi equations (see, e.g., [Yin03, Yin04, GL11, GL15]). In this paper, we consider equation (1.1) under two regimes of lower dissipation:
  • For $\frac{4}{3}<\gamma\leq 2$, we assume initial data $u_{0}\in H^{\frac{1}{2}+}(\mathbb{T})$.
  • For $\frac{5}{4}<\gamma\leq\frac{4}{3}$, we work with initial data in the intersection $u_{0}\in H^{\frac{1}{2}+}(\mathbb{T})\cap\mathcal{C}^{\frac{1}{2}+}(\mathbb{T})=W^{\frac{1}{2}+}(\mathbb{T})$.
It is also need to particularly noteworthy that the first author and co-authors have previously studied the randomized initial data problems for both the DP equation and Burgers equation
$\partial_{t}u-\Lambda^{\gamma}u=uu_{x},\quad u_{0}^{\omega}=\sum_{n\in\mathbb{Z}}h_{n}(\omega)\varphi(\Lambda-n)u_{0}.$ (1.3)
in [CZG24] for $\gamma\in(\frac{5}{4},\frac{3}{2})$, which is an another hot topic in the field of SPDEs. We omit detailed discussion here and refer interested readers to [Bou96, BT08b, BT08a, CG15, BOP15, BOP19, DNY22, DNY24]. Therefore, this paper can be viewed as a natural continuation of that earlier work under additional noise driving.
The primary objection of this paper is to establish fundamental results for the FRB equation in Sobolev spaces $H^{s}$ and extend analogous conclusions to the DP equations. We will address the potential difficulties associated with these problems, as well as the methods we will employ, by organizing the discussion into three main sections.
1.1 Probabilistic setting
The fact that $\xi\in\mathcal{C}^{-\frac{1}{2}-}(\mathbb{R}_{+}\times\mathbb{T})$ introduces significant analytical difficulties for direct approaches. One strategy is to write equation as mild form
$u=P(t)u_{0}+\int_{0}^{t}P(t-s)\partial_{x}(u^{2})\mathrm{d}s+X(t).$ (1.4)
Here $P(t)=e^{t\Lambda^{\gamma}}$ is the fractional heat flow, and $X(t)=\int_{0}^{t}P(t-s)\xi(s)\mathrm{d}s$ denote the solution of linear equation
$\partial_{t}X(t)-\Lambda^{\gamma}X(t)=\xi,\quad X_{0}=0.$ (1.5)
Noting that $X(t)\in C_{T}\mathcal{C}^{\alpha}(\mathbb{T})$ holds almost surely for any $\alpha<\frac{\gamma}{2}-\frac{1}{2}$, Such regularity has played a key role in establishing well-posedness for equations modeling thermally driven dissipative structures, notably the KPZ equation [Hai13] and the Burgers equation [GIP15, GP17]. In this paper, we extend analogous regularity results to the Sobolev spaces $H^{s}$ and apply them to prove well-posedness for the class of equations considered.
Since the admissible range of $\alpha$ depends on the dissipation parameter $\gamma$, this paper presents two regimes with distinct analytical treatments.
  • Higher regularity regime ($\alpha\in(\frac{1}{6},\frac{1}{2})$): in this case, well-posedness can be established directly in the Sobolev space $H^{\alpha}(\mathbb{T})$.
  • Lower dissipation regime ($\alpha\in(\frac{1}{8},\frac{1}{6})$, see Remark 3.1): here the analysis requires working in $W^{\alpha}(\mathbb{T})$.
A common feature in both regimes is the isolation of a higher-regularity component in the solution—more precisely, a term belonging to $H^{\frac{1}{2}+}(\mathbb{T})$. This decomposition is made possible for lower dissipation levels $\gamma\in(\frac{5}{4},\frac{4}{3}]$ by means of the paracontrolled calculus, a powerful tool introduced for studying rough and singular stochastic partial differential equations [GIP15, GP17, GKO23, CC18].
A more in-depth consideration is that how lower the dissipation be? Let us consider FRB driven by $|D|^{\beta}\xi$
$u_{t}-\Lambda^{\gamma}u=\partial_{x}(u^{2})+|D|^{\beta}\xi,\quad u(0,x)=u_{0}.$ (1.6)
Using the similar method in [DNY22, DNY24], we propose a conjecture regarding the relationship between $\gamma$ and $\beta$
$\gamma\geq\frac{2}{3}\beta+1$ (1.7)
in final section. Similar methods have been used in [DNY24] to study probabilistic critical problems for the heat equation, wave equation, and Schrödinger equation.
1.2 Regularity analysis
After resolving the regularity issues for Gaussian processes, the next challenge lies in the study of well-posedness. Consider the Burgers equation with addition term $f$
$\partial_{t}v-\Lambda^{\gamma}v=\partial_{x}(v^{2})+f,\quad v(0,x)=v_{0},$ (1.8)
where $f$ may depend on $u$. It takes the different structure of solution in two different ranges $\gamma\in(\frac{4}{3},2]$ and $\gamma\in(\frac{5}{4},\frac{4}{3}]$.
More precisely, let $f(v)=2\partial_{x}(vX)+\partial_{x}(X^{2})$, where $X$ satisfied (1.5). Then $f(v)$ gains at most $\alpha$-order Sobolev regularity when $v$ belongs to $H^{\frac{1}{2}+}$. It derive the relation between $\alpha$ and $\gamma$ which satisfied $\alpha+\gamma\geq\frac{3}{2}$, since $v$ can gain $\alpha+\gamma-1$ at most.
For the lower $\alpha$ and $\gamma$, our objective is to similarly isolate a higher-regularity remainder term ($H^{\frac{1}{2}+}$) in the solution’s structure. As the method used in the highly dissipative regime is no longer applicable here, a more refined analysis is required. Using the Littlewood–Paley decomposition (see [BCD11]), we rewrite $vX$ and $X^{2}$ as
$vX=T_{v}X+R(v,X)+T_{X}v,\quad X^{2}=2T_{X}X+R(X,X).$ (1.9)
The terms $T_{X}v$, $R(v,X)$ and $R(X,X)$ inherit higher regularity due to the fact that $X$ possesses $\alpha$-order Höler regularity. The main difficulty in the small-$\alpha$ case lies in handling the remaining terms $T_{v}X$ and $2T_{X}X$. We address this difficulty using the paracontrolled method (see Proposition 4.1 for details), which ultimately yields the condition $2\alpha-1+\gamma\geq\frac{3}{2}$.
1.3 Shallow water equation
Next, we introduce the Degasperis-Procesi(DP) equation, which is one type of shallow water equation, has the following form
$m_{t}+um_{x}+3u_{x}m=0,\quad m(t,x)=(1-\partial_{x}^{2})u(t,x),\quad u(0,x)=u_{0}(x).$ (1.10)
Then we can define the fluid velocity $u(t,x):=g\ast m:=(1-\partial_{x}^{2})^{-1}m$ and rewrite the DP equation as
$u_{t}+\frac{1}{2}\partial_{x}(u^{2})+\frac{3}{2}(1-\partial_{x}^{2})^{-1}\partial_{x}(u^{2})=0,\quad u(0,x)=u_{0}(x).$ (1.11)
The following citations [DP99, DHH02] provide physical background for this class of equations; a more systematic discussion can be found in [AG10]. Regarding the mathematical theory, we note in particular that local well-posedness of (1.10) for initial data in $H^{s}$ with $s>\frac{3}{2}$ was established in [LY06], together with a precise blow-up criterion and blow-up results in [LY07].
In this paper, we also consider the nonlocal Rough DP equation with space-time noise $\xi$
$u_{t}-\Lambda^{\gamma}u+\frac{1}{2}\partial_{x}(u^{2})+\frac{3}{2}(1-\partial_{x}^{2})^{-1}\partial_{x}(u^{2})=\xi,\quad u(0,x)=u_{0}(x),$ (1.12)
as an application of our result for Rough Burgers equation. Such method is also suitable for other type of shallow water equation with fractional dissipation.
1.4 Main result
We now present the main conclusions.
Theorem 1.1.
Let $u_{0}(x)\in H^{\frac{1}{2}+\delta}(\mathbb{T})$ for $0<\delta<\frac{3\gamma}{8}-\frac{1}{2}$, $\gamma\in(\frac{4}{3},2]$, then there exists a postive $T^{*}$ and a solution $u(t,x)$ of equation (1.1) satisfying
$u(t,x)=X(t,x)+v(t,x),$
where $v(t,x)\in C_{T^{*}}H^{\frac{1}{2}+\delta}(\mathbb{T})$ a.s. and $X(t,x)$ satisfies (1.5) with $X(t,x)\in C_{T^{*}}H^{\alpha}(\mathbb{T})$ a.s for $\alpha\in(\frac{1}{6},\frac{\gamma}{2}-\frac{1}{2})$. Moreover we have $u\in C_{T^{*}}H^{\alpha}(\mathbb{T})$ a.s..
Remark 1.1.
‘a.s.’ represents almost surely. We say event $A$ is almost surely if and only if $\mathbb{P}(A)=1$, where $\mathbb{P}$ represents the probability associated with $(\Omega,\mathcal{F},\mathbb{P})$. We will omit this description in most situation, except we need distinct such meaning.
Remark 1.2.
Recall the equation in [GIP15]
$\partial_{t}u-\Lambda^{\gamma}u=G(u)\partial_{x}u+\xi,\quad u(0,x)=u_{0},$ (1.13)
where $G\in C_{b}^{3}(\mathbb{R}^{d},L(\mathbb{R}^{d},\mathbb{R}^{d}))$. The main difference between our result and [GIP15] lies in the fact that our solution can back to the original equation (1.1). Since $G(u)\partial_{x}u$ is ill-defined when $\alpha\in(\frac{1}{3},\frac{\gamma}{2}-\frac{1}{2})$, they actually study the “generalized" solution, which is the limit in probability sense of $u^{\epsilon}$ satisfying the following mollified equation
$\partial_{t}u^{\epsilon}-\Lambda^{\gamma}u^{\epsilon}=G(u^{\epsilon})\partial_{x}u^{\epsilon}+\xi^{\epsilon},\quad u(0,x)=u_{0}.$ (1.14)
Here $\xi^{\epsilon}=\epsilon^{-1}\psi(\epsilon\cdot)\ast\xi$ for some $\psi\in\mathcal{S}$ with $\int\psi\mathrm{d}t=1$. Similar approaches are frequently used for some singular SPDEs in [CC18, Hai13, GP17].
When $\gamma>\frac{3}{2}$, we can get the global result as follows.
Theorem 1.2.
Let $\gamma\in(\frac{5}{3},2]$, $u_{0}\in H^{\sigma}$ for $\sigma>0$ and $\int_{\mathbb{T}}u(x)\mathrm{d}x=0$, then Rough Burgers equation (1.1) has a global mild solution $u=v+X$ where $X$ satisfies (1.5) and $v$ belongs to $C([0,T];L^{2}(\mathbb{T}))\cap L^{2}([0,T];H^{\frac{\gamma}{2}}(\mathbb{T}))$ for any $T\geq 0$.
Remark 1.3.
Theorem 1.1 and 1.2 are also valid for Rough DP equation. In fact, the nonlinear term $\frac{3}{2}(1-\partial_{x}^{2})^{-1}\partial_{x}(u^{2})$ in the DP equation does not introduce additional difficulties for the local well-posedness problem. For global results, we need to utilize the special structure of the DP equation to derive energy estimates as Proposition 3.3 and Lemma 3.4.
For the lower dissipation $\gamma\in(\frac{5}{4},\frac{4}{3}]$, let $\mathcal{Q}$ be the linear evolution of $\partial_{x}X$ such that
$\partial_{t}\mathcal{Q}(t)-\Lambda^{\gamma}\mathcal{Q}(t)=\partial_{x}X(t),\quad\mathcal{Q}(0)=0.$ (1.15)
Let $(u,u^{\prime},u^{\sharp})$ be a triple satisfying
$u=X+u^{\prime}\prec\!\!\!\prec\mathcal{Q}+u^{\sharp},$ (1.16)
where $u^{\prime}\prec\!\!\!\prec\mathcal{Q}$ is the temporal mollified version of $T_{u^{\prime}}\mathcal{Q}$(See (2.2)). $u^{\sharp}$ is the higher regularity term.
Theorem 1.3.
Let $u_{0}\in\mathcal{C}^{\frac{1}{2}+\delta}\cap H^{\frac{1}{2}+\delta}=W^{\frac{1}{2}+\delta}$ for $0<\delta<\frac{\gamma}{2}-\frac{5}{8}$, $\gamma\in(\frac{5}{4},\frac{4}{3}]$. Define $\mathcal{L}=\partial_{t}-\Lambda^{\gamma}$, then there exists a positive $T^{*}$ and a unique triple $(u,u^{\prime},u^{\sharp})\in C_{T}W^{\alpha}\times C_{T}W^{\frac{1}{8}+\delta}\times C_{T}W^{\frac{1}{2}+\delta}$ satisfying
$u^{\prime}=$ $2X+2u^{\prime}\prec\!\!\!\prec\mathcal{Q}+2u^{\sharp},$ (1.17)
$\mathcal{L}u^{\sharp}=$ $\mathcal{L}u-\mathcal{L}X-\mathcal{L}u^{\prime}\prec\!\!\!\prec\mathcal{Q},\quad u^{\sharp}(0)=u_{0}$
$u=$ $X+u^{\prime}\prec\!\!\!\prec\mathcal{Q}+u^{\sharp},$
where $X(t)$ satisfies (1.5) with $X(t,x)\in C_{T^{*}}W^{\alpha}(\mathbb{T})$ a.s for $\alpha\in(\frac{1}{8},\frac{\gamma}{2}-\frac{1}{2})$. Moreover, $u$ is a solution of (1.1).
Remark 1.4.
$\prec$ and $\prec\!\!\!\prec$ are referred to as the para-product and modified para-product, which will be introduced in Lemma 2.1 and (2.2). Let $v:=u^{\prime}\prec\!\!\!\prec\mathcal{Q}$, then $v$ satisfies the equation
$\partial_{t}v-\Lambda^{\gamma}v=^{\prime}regular\;term^{\prime}+u^{\prime}\prec\partial_{x}X.$
This term will appear in the equation for $u^{\sharp}$. Our strategy is to use a specific $u^{\prime}$ with $\mathcal{L}(u^{\prime}\prec\!\!\!\prec\mathcal{Q})$ to cancel out this term to make sure that $u^{\sharp}$ can gain higher regularity (See Proposition 4.1). We will show it in Section 4.
Remark 1.5.
In fact, we can’t prove such result with only $H^{\alpha}$ space under our setting. Let $(u^{\prime},\mathcal{Q},u^{\sharp})\in(\mathcal{C}^{\alpha}\times H^{\beta}\times H^{\eta})$, then $\alpha+\beta=\eta$ is always valid. Since $\mathcal{Q}$ can gain $(\alpha-1+\gamma)$-order Sobolev regularity, $u^{\sharp}$ gains $(2\alpha-1+\gamma)$-order Sobolev regularity at most by proof of Lemma B.3. To establish the $C^{\alpha}$ Hölder regularity of $u^{\prime}$, it is necessary for the remainder $u^{\sharp}$ to possess Sobolev regularity of order at least $H^{\alpha+\frac{1}{2}+}$, this is valid only when $\gamma>\frac{4}{3}$.
Remark 1.6.
We can extend the range to $\gamma\in(1,2]$, which means $\alpha\in(0,\frac{1}{2})$. We omit these proofs because in this paper, we primarily aim to demonstrate the effectiveness of the paracontrolled method in enhancing the regularity of solutions.
1.5 Organization of the paper
This paper is organized as follows. In Section 2, we introduce some notations and basic tools. In Section 3, we prove Theorem 1.1 and Theorem 1.2 for the fractional stochastic Burgers equation, and we also prove the result in Remark 1.3 for the rough Degasperis–Procesi equation. In Section 4, we prove Theorem 1.3 for the paracontrolled solution. In Section 5, we provide a conjecture on the range of $\gamma$ in Sobolev space and discuss some future research directions.
2 Basic tool
In this section, we introduce some basic stochastic and deterministic tools, more useful technical estimation tools will be introduced in the appendix.
2.1 Littlewood-Paley theory
The first key tool is Littlewood-Paley decomposition. We introduce an operator named “nonhomogeneous dyadic blocks” $\Delta_{j}$ for $j\geq-1$, and satisfying that for any $u$ belong to $\mathcal{S}^{\prime}(\mathbb{T}^{d})$ (It’s also valid for $\mathbb{R}^{d}$), we have decomposition
$u=\sum_{j\geq-1}\Delta_{j}u.$
More precisely, it’s associated with two functions $\rho_{l}(k)$ and $\chi(k)$ under Fourier mode such that
$\mathcal{F}(\Delta_{l}u)(k)=\rho(2^{-l}k)u_{k},\quad l\geq 0\quad and\quad\mathcal{F}(\Delta_{-1}u)(k)=\mathcal{\chi}(k)u_{k}.$
Here $\chi$ and $\rho$ are two radial functions, valued in the interval $[0,1]$, belonging respectively to $\mathcal{D}(B(0,\frac{4}{3}))$ and $\mathcal{D}(\{k\in\mathbb{Z}^{d}:\frac{3}{4}\leq k\leq\frac{8}{3}\})$, such that
$\chi(k)+\sum_{l\geq 0}\rho(2^{-l}k)=1,\quad k\in\mathbb{Z}^{d}.$
It can define nonhomogeneous Besov space $B_{p,r}^{s}(\mathbb{T}^{d})$ which contains Hölder space $\mathcal{C}^{s}(\mathbb{T}^{d}):=B_{\infty,\infty}^{s}(\mathbb{T}^{d})$ for $\alpha\in(0,1)$ and nonhomogeneous Sobolev space $H^{s}(\mathbb{T}^{d}):=B^{s}_{2,2}(\mathbb{T}^{d})$ such that
$B_{p,r}^{s}(\mathbb{T}^{d}):=\left\{u\in S^{\prime}:\left(\sum_{l\geq-1}2^{lsr}\|\Delta_{l}u\|_{L^{p}(\mathbb{T}^{d})}^{r}\right)^{\frac{1}{r}}<\infty\right\},$
also define the norm. Also define the Fourier multiplier
$\mathcal{F}(\dot{\Delta}_{j}u)(k)=\rho(2^{-l}k)u_{k},\quad l\in\mathbb{Z},$
where $\rho$ is a smooth cut-off function supported in an annulus. Then we introduce the homogeneous Besov space
$\dot{B}_{p,r}^{s}(\mathbb{T}^{d}):=\left\{u\in\mathcal{S}_{h}^{\prime}:\left(\sum_{l\in\mathbb{Z}}2^{lsr}\|\dot{\Delta}_{l}u\|_{L^{p}(\mathbb{T}^{d})}^{r}\right)^{1/r}<\infty\right\},$
where $\mathcal{S}_{h}^{\prime}$ denotes the space of tempered distributions $u$ such that for any $\theta\in\mathcal{D}(\mathbb{T}^{d})$ with value $1$ near 0,
$\lim_{\lambda\to\infty}\|\theta(\lambda D)u\|_{L^{\infty}}=0.$ (2.1)
It can be verified that $\dot{B}_{2,2}^{s}(\mathbb{T}^{d})$ coincides with the homogeneous Sobolev space
$\dot{H}^{s}(\mathbb{T}^{d})=\left\{u\in\mathcal{S}_{h}^{\prime}:\left(\sum_{k\neq 0}|k|^{2s}|\mathcal{F}(u)(k)|^{2}\right)^{1/2}<\infty\right\}.$
It is straightforward to verify that condition (2.1) implies $\int_{\mathbb{T}}u(x)\,dx=\widehat{u}(0)=0$. Consequently, a function $f$ belongs to $\dot{H}^{s}$ if and only if $f\in H^{s}$ and $\int_{\mathbb{T}}f(x)\,dx=0$. Under this setting, we introduce para-product decomposition which give a product rules in $\mathcal{S}^{\prime}(\mathbb{T}^{d})$. Let $S_{j}u=\sum_{k\leq j-1}\Delta_{k}u$, then we have the estimate
Lemma 2.1 ([BCD11]).
Define $f\prec g:=\sum_{j}S_{j-1}f\Delta_{j}g$, $f\succ g:=g\prec f$, $f\circ g:=\sum_{|k-j|\leq 1}\Delta_{k}f\Delta_{j}g$, let $p,r\in[1,\infty]$ and $s\in\mathbb{R}$, then we have the following estimates
  • 1. If $k\in\mathbb{N}$, then
    $\|f\prec g\|_{B^{s}_{p,r}}\leq C\|f\|_{L^{\infty}}\|D^{k}g\|_{B_{p,r}^{s-k}}.$
  • 2. If $k\in\mathbb{N}$, $t<0$ and $\frac{1}{r}=\min(1,\frac{1}{r_{1}}+\frac{1}{r_{2}})$, then
    $\|f\prec g\|_{B_{p,r}^{s+t}}\leq C\|f\|_{B_{\infty,r_{1}}^{t}}\|D^{k}g\|_{B_{p,r_{2}}^{s-k}}.$
  • 3. If $\frac{1}{p}=\frac{1}{p_{1}}+\frac{1}{p_{2}}\leq 1$, $\frac{1}{r}=\frac{1}{r_{1}}+\frac{1}{r_{2}}\leq 1$ and $s=s_{1}+s_{2}>0$, then
    $\|f\circ g\|_{B_{p,r}^{s}}\leq C\|f\|_{B_{p_{1},r_{1}}^{s_{1}}}\|g\|_{B_{p_{2},r_{2}}^{s_{2}}}.$
The above estimate also holds if we replace $\dot{\Delta}_{j}$ with $\Delta_{j}$ and $\dot{S}_{j}$ with $S_{j}$.
By above estimate in $H^{s}$ space, we can check that Moser-type estimate hold.
Lemma 2.2 (Moser-type estimate).
Let $s>0$, $\epsilon>0$, then the following estimates hold:
  • (1) for $0<s<\frac{1}{2}$,
    $\|fg\|_{H^{2s-\frac{1}{2}}}\leq C\|f\|_{H^{s}}\|g\|_{H^{s}}.$
  • (2) for $s>\frac{1}{2}$,
    $\|fg\|_{H^{s}}\leq C\|f\|_{H^{s}}\|g\|_{H^{s}}.$
  • (3) for $s=\frac{1}{2}$,
    $\|fg\|_{H^{s-\epsilon}}\leq C\|f\|_{H^{s}}\|g\|_{H^{s}}.$
To explain the paraproduct, we first note that two distributions $f$ and $g$ cannot, in general, be multiplied directly. A sufficient condition for such a product to be well defined is that the requirement of Lemma 2.1 be satisfied, which at least necessitates $s_{1}+s_{2}\geq 0$ . Moreover, $f\prec g$ and $f\succ g$ are called para-products since they are always well defined regardless of the regularity of $f$ and $g$. The term $f\circ g$ is referred to as the resonant term, because it involves interactions of frequencies that are close to each other, which typically restricts the range of regularities for which it can be controlled.
For our convenience, we give the following notation
$W_{p}^{\alpha}=B_{p,p}^{\alpha},\quad W^{\alpha}=W_{2}^{\alpha}\cap W_{\infty}^{\alpha},\quad$
and note that $W^{\alpha}$ is a Banach algebra when $\alpha>0$.
Lemma 2.3.
If $f\in W^{\alpha}$, then $f\in W_{p}^{\alpha}$ for any $p\geq 2$. Moreover, we have estimate
$\|f\|_{W_{p}^{\alpha}}\leq\|f\|_{H^{\alpha}}^{\frac{2}{p}}\|f\|_{\mathcal{C}^{\alpha}}^{\frac{p-2}{p}}.$
Proof.
By definition of $B_{p,p}^{\alpha}$ and interpolation in $L^{p}$, we have
$\|f\|_{B_{p,p}^{\alpha}}^{p}=$ $\sum_{j\geq-1}2^{j\alpha p}\|\Delta_{j}f\|^{p}_{L^{p}}$
$\leq$ $\sum_{j\geq-1}2^{j\alpha p}(\|\Delta_{j}f\|^{\frac{2}{p}}_{L^{2}}\|\Delta_{j}f\|^{\frac{p-2}{p}}_{L^{\infty}})^{p}$
$\leq$ $(\sup_{j\geq-1}2^{j\alpha}\|\Delta_{j}f\|_{L^{\infty}})^{p-2}\sum_{j\geq-1}2^{j\alpha 2}\|\Delta_{j}f\|^{2}_{L^{2}}$
$\leq$ $\|f\|_{H^{\alpha}}^{2}\|f\|_{B_{\infty,\infty}^{\alpha}}^{p-2}.$
Then our lemma is proved. ∎
Sometimes it is also necessary to use a modified version of the paraproduct. Such operators have been introduced in [GIP15, GP17], and are defined as follows: Let $\varphi\in C_{c}^{\infty}(\mathbb{R},\mathbb{R}_{+})$ be nonnegative with compact support contained in $\mathbb{R}_{+}$ and with total mass 1, define $Q_{i}$ for all $i\geq 1$ as follows
$Q_{i}:C\mathcal{C}^{\beta}\to C\mathcal{C}^{\beta},\quad Q_{i}f(t)=\int_{\mathbb{R}}2^{\gamma i}\varphi(2^{\gamma i}(t-s))f(s\vee 0)\mathrm{d}s.$
We will often apply $Q_{i}$ and other operators on $C\mathcal{C}^{\beta}$ to functions $f\in C_{T}\mathcal{C}^{\beta}$ which we then simply extend from $[0,T]$ to R+ by considering $f(\cdot\wedge T)$. With the help of $Q_{i}$, we define a modified para-product
$f\prec\!\!\!\prec g=\sum_{i}(Q_{i}S_{i-1}f)\Delta_{i}g.$ (2.2)
Similar to estimate for the paraproduct in Lemma 2.1, we also have estimates for this operator.
Lemma 2.4.
For any $\beta\in\mathbb{R}$ and $p\geq 2$, we have
$\|f\prec\!\!\!\prec g(t)\|_{W^{\beta}_{p}}\leq C\|f\|_{C_{t}L^{\infty}}\|g\|_{W^{\beta}_{p}},$
for all $t>0$. And for $1<\alpha<0$ furthermore
$\|f\prec\!\!\!\prec g(t)\|_{W^{\alpha+\beta}_{p}}\leq C\|f\|_{C_{t}\mathcal{C}^{\alpha}}\|g\|_{W^{\beta}_{p}}.$
Proof.
Noting that $\mathcal{F}(\mathcal{Q}_{j}S_{j-1}f(s)\Delta_{j}g(t))\in 2^{j}\tilde{\mathcal{C}}$, where $\tilde{\mathcal{C}}$ is some annuals, we are left with proving an appropriate estimate for $\|\mathcal{Q}_{j}S_{j-1}f(s)\Delta_{j}g(t)\|_{L^{p}}$. By Hölder inequality, and definition of $\varphi$, we have
$\|\int_{0}^{t}2^{\gamma j}\varphi(2^{\gamma j}(t-s))S_{j-1}f(s)\mathrm{d}s\Delta_{j}g(t)\|_{L^{p}}\lesssim$ $\int_{0}^{t}2^{\gamma j}\varphi(2^{\gamma j}(t-s))\|S_{j-1}f(s)\|_{L^{\infty}}\mathrm{d}s\|\Delta_{j}g(t)\|_{L^{p}}$
$\lesssim$ $\int_{\mathbb{R}}2^{\gamma j}\varphi(2^{\gamma j}(t-s))\mathrm{d}s\|f\|_{C_{t}L^{\infty}}\|\Delta_{j}g(t)\|_{L^{p}}$
$\lesssim$ $\|f\|_{C_{t}L^{\infty}}\|\Delta_{j}g(t)\|_{L^{p}}.$
Since $g(t)\in W_{p}^{\beta}$, by Lemma 2.23 in [BCD11], we get the first estimate. The second estimate is similar, requiring only the observation that
$\sup_{j\geq-1}2^{j\alpha}\|S_{j-1}f(s)\|_{L^{\infty}}\leq C\|f(s)\|_{\mathcal{C}^{\alpha}},$ (2.3)
for $1<\alpha<0$ by Proposition 2.79 in [BCD11]. ∎
The remaining estimates will be presented in the Appendix B. Finally, we introduce a class of temporal regularity spaces, which have been referenced in [GIP15, GP17, Hai13, CC18]. Define the such norm with time $T$ as follows. $C_{T}X=C([0,T],X)$ for the space of continuous maps from $[0,T]$ to $X$, equipped with the norm $\sup_{t\in[0,T]}\|\cdot\|_{X}$. For $\alpha\in(0,1)$, we also define $C_{T}^{\alpha}X$ as a space of $\alpha$-Hölder continuous functions from $[0,T]$ to $X$ equipped with the norm $\|f\|_{C_{T}^{\alpha}X}=\sup_{0\leq s<t\leq T}\frac{\|f(t)-f(s)\|_{X}}{|t-s|^{\alpha}}$. For our convenience, for $\alpha\geq 0$ we have the following notation
$\mathcal{W}_{p}^{0}(T)=C_{T}L^{p},\quad\mathcal{W}_{p}^{\alpha}(T)=C_{T}B_{p,p}^{\alpha}\cap C^{\frac{\alpha}{\gamma}}_{T}L^{p},\quad\mathcal{W}^{\alpha}(T)=\mathcal{W}_{2}^{\alpha}(T)\cap\mathcal{W}_{\infty}^{\alpha}(T).$
It’s easy to check that $\mathcal{W}^{\alpha}(T)$ is a Banach algebra when $\alpha\geq 0$. For $\alpha<0$, we also give a definition $\mathcal{W}^{\alpha}_{p}(T)=C_{T}B_{p,p}^{\alpha}$ Next, we introduce some useful estimate
Lemma 2.5.
Let $\alpha\in\mathbb{R}$, $\epsilon,T>0$, $\rho\in(0,\gamma)$ satisfy $\gamma\leq\rho+\epsilon$ and $\alpha+\rho<\gamma$. Let $f\in\mathcal{W}_{\infty}^{\epsilon}(T)$, $g(0)=0$, $g\in C_{T}W^{\alpha+\rho}(T)$ and $\mathcal{L}g\in C_{T}W^{\alpha}$. Then
$\|f\prec\!\!\!\prec g\|_{\mathcal{W}_{\infty}^{\alpha+\rho}(T)}\lesssim T^{1-\frac{\rho}{\gamma}}\|f\|_{\mathcal{W}_{\infty}^{\epsilon}(T)}(\|g\|_{C_{T}\mathcal{C}^{\alpha+\gamma-\epsilon}}+\|\mathcal{L}g\|_{C_{T}\mathcal{C}^{\alpha}}),$
and
$\|f\prec\!\!\!\prec g\|_{\mathcal{W}_{2}^{\alpha+\rho}(T)}\lesssim T^{1-\frac{\rho}{\gamma}}\|f\|_{\mathcal{W}_{\infty}^{\epsilon}(T)}(\|g\|_{C_{T}H^{\alpha+\gamma-\epsilon}}+\|\mathcal{L}g\|_{C_{T}H^{\alpha}}).$
Proof.
We only prove the case of $\mathcal{W}_{2}^{\alpha+\rho}(T)$ and the $p=\infty$ is similar. By Lemma A.2, Lemma 2.4 and Lemma B.3, we have
$\|f\prec\!\!\!\prec g\|_{\mathcal{W}_{2}^{\alpha+\rho}(T)}\lesssim$ $\|f\prec\!\!\!\prec g(0)\|_{\mathcal{W}_{2}^{\alpha+\rho}(T)}+T^{1-\frac{\rho}{\gamma}}\|\mathcal{L}(f\prec\!\!\!\prec g)\|_{C_{T}H^{\alpha}}$
$\leq$ $T^{1-\frac{\rho}{\gamma}}\|\mathcal{L}(f\prec\!\!\!\prec g)-f\prec\!\!\!\prec\mathcal{L}g\|_{C_{T}H^{\alpha}}+\|f\prec\!\!\!\prec\mathcal{L}g\|_{C_{T}H^{\alpha}}$
$\leq$ $T^{1-\frac{\rho}{\gamma}}\|f\|_{C_{T}\mathcal{C}^{\epsilon}}\|g\|_{C_{T}H^{\alpha+\gamma-\epsilon}}+\|f\|_{C_{T}\mathcal{C}^{\epsilon}}\|\mathcal{L}g\|_{C_{T}H^{\alpha}}$
$\leq$ $T^{1-\frac{\rho}{\gamma}}\|f\|_{C_{T}\mathcal{C}^{\epsilon}}(\|g\|_{C_{T}H^{\alpha+\gamma-\epsilon}}+\|\mathcal{L}g\|_{C_{T}H^{\alpha}})$
$\leq$ $T^{1-\frac{\rho}{\gamma}}\|f\|_{C_{T}\mathcal{C}^{\epsilon}}(\|g\|_{C_{T}H^{\alpha+\rho}}+\|\mathcal{L}g\|_{C_{T}H^{\alpha}}).$
Since $\gamma-\epsilon\leq\rho$, then we prove the lemma. ∎
2.2 Stochastic tools
In this little section, we introduce some stochastic tools we need in this paper. The first tool is the Kolmogorov’s continuity criterion.
Lemma 2.6 ([LG16]).
Let $X=(X(t))_{t\in[0,T]}$ be a random process and take values in a complete metric space $(E,d)$. If there exist $p,\epsilon,C>0$ such that, for every $s,t\in[0,T]$,
$\mathbb{E}[d(X(s),X(t))^{p}]\leq C|t-s|^{1+\epsilon}.$ (2.4)
Then, there is a modification $\tilde{X}$ of $X$ whose sample paths are Hölder continuous with exponent $\alpha$ for every $\alpha\in(0,\frac{\epsilon}{p})$. This means that, for every $\omega\in\Omega$ and every $\alpha\in(0,\frac{\epsilon}{p})$, there exists a finite constant $C_{\alpha}(\omega)$ such that, for every $s,t\in I$,
$d(\tilde{X}(s,\omega),\tilde{X}(t,\omega))\leq C_{\alpha}(\omega)|t-s|^{\alpha},$
where $\tilde{X}(t)$ is a modification of $X(t)$ with continuous sample paths.
Now we give a basic analysis for white noise $\xi$. For the linear evolution $X(t):=\int_{0}^{t}e^{(t-s)\Lambda^{\gamma}}\xi(s)\mathrm{d}s$, we have the following lemma.
Lemma 2.7.
The spatial Fourier transform $X_{k}=\mathcal{F}_{x}X(k)$ is a Gaussian process with zero-mean and satisfy
$\mathbb{E}[X_{k}(t)X_{k^{\prime}}(t^{\prime})]=(e^{-|t-t^{\prime}||k|^{\gamma}}-e^{-(t+t^{\prime})|k|^{\gamma}})\frac{|\mathbb{T}|}{2|k|^{\gamma}}\textbf{I}_{k+k^{\prime}=0}(k,k^{\prime}),\quad\mathbb{E}[X_{0}(t)X_{k}(t^{\prime})]=(t\wedge t^{\prime})|\mathbb{T}|\textbf{I}_{k=0}(k).$
where $|\mathbb{T}|=\int_{\mathbb{T}}1\mathrm{d}x$ and $\textbf{I}_{k+k^{\prime}=0}(k,k^{\prime})$ is the indicator function. Moreover, for any $T>0$ and $\beta<\frac{\gamma(1-2\kappa)}{2}-\frac{1}{2}$, we have $X(t)\in C_{T}^{\kappa}W^{\beta}$ for $\kappa<\frac{1}{2}$.
Proof.
It’s easy to check the first result by noting that
$\mathbb{E}[\mathcal{F}\xi(t,k){\mathcal{F}\xi(s,k^{\prime})}]=\delta(t-s)\textbf{I}_{k+k^{\prime}=0}(k,k^{\prime}).$ (2.5)
Considering the second result, we can rewrite $X(t)$ as
$X(t)=\sum_{k\in\mathbb{Z}}X_{k}(t)e^{ikx}.$
For $j\geq 0(k\neq 0)$, we calculate $\mathbb{E}|\Delta_{j}(X(t)-X(t^{\prime}))|^{2}$ as follows
$\mathbb{E}|\Delta_{j}(X(t)-X(t^{\prime}))|^{2}=$ $\mathbb{E}\sum_{k,k^{\prime}\in\mathbb{Z}}\rho_{j}(k)\rho_{j}(k^{\prime})[X_{k}(t)-X_{k}(t^{\prime})][X_{-k^{\prime}}(t)-X_{-k^{\prime}}(t^{\prime})]e^{i(k-k^{\prime})x}$
$\lesssim$ $2^{j(1-\gamma(1-\delta))}|t-t^{\prime}|^{\delta},$
for some $\delta\in[0,1]$. The case j=1j=-1 can be treated using essentially the same arguments, except that then we need to distinguish the cases $k=0$ and $k\neq 0$. By Gaussian hypercontractivity [Jan97], we have
$\mathbb{E}|\Delta_{j}(X(t,x_{i})-X(t^{\prime},x_{i}))|^{2p}\lesssim(\mathbb{E}|\Delta_{j}(X(t,x_{i})-X(t^{\prime},x_{i}))|^{2})^{p}.$
It implies that
$\mathbb{E}\|\Delta_{j}(X(t)-X(t^{\prime}))\|_{L^{2}}^{2p}=$ $\mathbb{E}(\int_{\mathbb{T}}|\Delta_{j}(X(t)-X(t^{\prime}))|^{2}\mathrm{d}x)^{p}$
$\lesssim$ $\int_{\mathbb{T}^{\otimes p}}\mathbb{E}\prod_{i=1}^{p}|\Delta_{j}(X(t,x_{i})-X(t^{\prime},x_{i}))|^{2}\mathrm{d}x_{i}$
$\lesssim$ $\int_{\mathbb{T}^{\otimes p}}\prod_{i=1}^{p}(\mathbb{E}|\Delta_{j}(X(t,x_{i})-X(t^{\prime},x_{i}))|^{2p})^{\frac{1}{p}}\mathrm{d}x_{i}$
$\lesssim$ $\int_{\mathbb{T}^{\otimes p}}\prod_{i=1}^{p}(\mathbb{E}|\Delta_{j}(X(t,x_{i})-X(t^{\prime},x_{i}))|^{2})\mathrm{d}x_{i}$
$\lesssim$ $2^{pj(1-\gamma(1-\delta))}|t-t^{\prime}|^{\delta p}.$
Then we can estimate $\mathbb{E}\|X(t)-X(t^{\prime})\|_{H^{\beta}}^{2p}$ as
$\mathbb{E}\|X(t)-X(t^{\prime})\|_{B_{2,2}^{\beta}}^{2p}\lesssim$ $(\sum_{j\geq-1}2^{2j\beta}\mathbb{(}E\|\Delta_{j}(X(t)-X(t^{\prime}))\|_{L^{2}}^{2p})^{\frac{1}{p}})^{p}$
$\leq$ $(\sum_{j\geq-1}2^{j(2\beta+1-\gamma(1-\delta))})^{p}|t-t^{\prime}|^{\delta p}\leq C|t-s|^{\delta p},$
for $\beta<\frac{\gamma(1-\delta)-1}{2}$. By Lemma 2.6, we have $X(t)\in C^{\frac{\delta}{2}-\frac{1}{2p}}_{T}H^{\frac{\gamma(1-\delta)-1}{2}}$. Choose $p$ large enough, we can prove the case of $H^{s}$. For Hölder space, we can also have the results by similar argument
$\mathbb{E}\|X(t)-X(t^{\prime})\|_{B_{2p,2p}^{\beta}}^{2p}=$ $\sum_{j\geq-1}2^{2j\beta p}\mathbb{E}\|\Delta_{j}(X(t)-X(t^{\prime}))\|_{L^{2p}}^{2p}$
$\lesssim$ $\sum_{j\geq-1}2^{2j\beta p}\|\mathbb{E}|\Delta_{j}(X(t)-X(t^{\prime}))|^{2}\|_{L^{p}}^{p}$
$\lesssim$ $\sum_{j\geq-1}2^{jp(2\beta+1-\gamma(1-\delta))}|t-t^{\prime}|^{\delta p}\leq C|t-s|^{\delta p}.$
Let $p$ be large enough and by embedding for Besov space, we get $X(t)\in C^{\frac{\delta}{2}-\frac{1}{2p}}_{T}\mathcal{C}^{\frac{\gamma(1-\delta)-1}{2}}$. Choosing $p$ large enough again, we can prove the case of $\mathcal{C}^{s}$, which we finish our proof. ∎
Remark 2.1.
By Lemma 2.7, we have $X(t)\in C_{T}^{\kappa}W^{\beta}$ for $\frac{\beta}{\gamma}<\frac{1}{2}-\frac{1}{\gamma}-\kappa$. Let $\beta=0$, then $\kappa$ can choose to be $\frac{1}{2}-\frac{1}{2\gamma}-\epsilon$ for any small $\epsilon>0$. Then we have $X(t)\in\mathcal{W}^{\alpha}(T)$ a.s for any $T\geq 0$ and $\alpha<\frac{\gamma}{2}-\frac{1}{2}$ if we fix $\epsilon$ small enough.
2.3 Iteration argument
Next we introduce a generalized contraction mapping theorem which compared with Lemma 5.5($L=0$) in [BCD11].
Lemma 2.8.
Let $E$ be a Banach space, and $\mathcal{B}$ a continuous bilinear map from $E\times E$ to $E$, and $r$ a positive real number such that
$r<\frac{1}{8\|\mathcal{B}\|}\quad with\quad\|\mathcal{B}\|_{E}=\sup_{\|u\|_{E},\|v\|_{E}\leq 1}\|\mathcal{B}(u,v)\|_{E}.$
Let $L$ is a linear map $E$ to $E$ and satisfied
$\sup_{v\in E}\frac{\|L(v)\|_{E}}{\|v\|_{E}}\leq\frac{1}{4}.$
Then for any $a$ in a ball $B(0,r)\subset E$, there exists a unique $\tilde{x}$ in $\mathcal{B}(0,2r)$ such that
[…]
Proof.
The proof of first result can refer to [BCD11]. For second, we involved an application of iterative scheme defined by
$\tilde{x}_{0}=a\quad\tilde{x}_{n+1}=a+L(\tilde{x}_{n})+\mathcal{B}(\tilde{x}_{n},\tilde{x}_{n}).$
By induction argument, we show that $\|\tilde{x}_{n+1}\|\leq 2r$ when $\|\tilde{x}_{n}\|\leq 2r$. In fact, by our assumption, we get
$\|\tilde{x}_{n+1}\|_{E}\leq r+\frac{r}{4}+4\|\mathcal{B}\|r^{2}\leq 2r.$
Thus, $(\tilde{x}_{n})_{n\in\mathbb{N}}$ remains in $B(0,2r)$. Moreover,
$\|\tilde{x}_{n+1}-\tilde{x}_{n}\|_{E}\leq$ $\|L(\tilde{x}_{n}-\tilde{x}_{n-1})\|_{E}+4r\|\mathcal{B}\|\|\tilde{x}_{n}-\tilde{x}_{n-1}\|_{E}$
$\leq$ $\frac{3}{4}\|\tilde{x}_{n}-\tilde{x}_{n-1}\|_{E}.$
[…]
We also provide version for studying specific types of coupled equations. Let $f((x,y))$ be a linear functional on $E\times E$ which means
$f(c(x,y))=cf((x,y)),$ (2.6)
$f(x_{1}+x_{2},y_{1}+y_{2})=f(x_{1},y_{1})+f(x_{2},y_{2}).$
Assume that $f(x,y)$ satisfies some control property from $\dot{E}\times E$ to $X$, such as
$\|f(x,y)\|_{X}\leq\|x\|_{\dot{E}}+c\|y\|_{E},$ (2.7)
we have the following lemma
Lemma 2.9.
Let $\dot{E},E,X$ be Banach spaces satisfying $\dot{E}\subset X\subset E$, $\mathcal{E}:=\dot{E}\times E$ be a Banach space with norm
$\|(x,y)\|_{\mathcal{E}}=2\|x\|_{\dot{E}}+\|y\|_{E}.$
Moreover let $\mathcal{B}$ be a continuous bilinear map from $X\times X$ to $\dot{E}$, and $r$ a positive real number such that
$\|\mathcal{B}(u,v)\|_{\dot{E}}\leq\frac{1}{16r}\|u\|_{X}\|v\|_{X},$
Let $\mathcal{G}(x)$ and $g(y)$ be linear map from $E$ to $\dot{E}$ and satisfying
$\sup_{v\in E}(\frac{\|\mathcal{G}(v)\|_{\dot{E}}}{\|v\|_{E}},\frac{\|g(v)\|_{\dot{E}}}{\|v\|_{E}})\leq\frac{1}{16},\quad\frac{\|h(v)\|_{\dot{E}}}{\|v\|_{\dot{E}}}\leq\frac{1}{16}.$
Let $f$ satisfy (2.6) and (2.7) for $X$, then there exist $\tilde{c}>0$, such that for any $(a,b)\in B(0,\frac{r}{4})\subset\mathcal{E}$ and $0<c<\tilde{c}$, there exists unique $(x,y)\in B(0,4r)\subset\mathcal{E}$ such that
$y=b+f(x,y),\quad x=a+g(y)+h(x)+\mathcal{B}(f(x,y),f(x,y))+\mathcal{G}(f(x,y)).$ (2.8)
Proof.
We involved application of iterative scheme defined by
$x_{0}=a,\quad y_{0}=b,\quad y_{n+1}=b+f(x_{n},y_{n}),\quad x_{n+1}=a+g(y_{n})+h(x_{n})+\mathcal{G}(f(x_{n},y_{n}))+\mathcal{B}(f(x_{n},y_{n}),f(x_{n},y_{n})).$
We first prove that $(x_{n+1},y_{n+1})\in B(0,4r)\subset\mathcal{E}$ when $(x_{n},y_{n})\in B(0,4r)\subset\mathcal{E}$. Observing that $\|x_{n}\|\leq 2r$, $\|y_{n}\|\leq 4r$ and
$\|f(x_{n},y_{n})\|_{E}\leq\|f(x_{n},y_{n})\|_{X}\leq\|x_{n}\|_{\dot{E}}+c\|y_{n}\|_{E}\leq 2r+4cr,$
by a directly computation, we have
$\|(x_{n+1},y_{n+1})\|_{\mathcal{E}}\leq$ $\|b\|_{E}+\|f(x_{n},y_{n})\|_{E}+2\|a\|_{\dot{E}}+2\|g(y_{n})\|_{\dot{E}}+2\|h(x_{n})\|_{\dot{E}}$
$+2\|\mathcal{G}(f(x_{n},y_{n}))\|_{\dot{E}}+2\|\mathcal{B}(f(x_{n},y_{n}),f(x_{n},y_{n}))\|_{\dot{E}}$
$\leq$ $2\|a\|_{\dot{E}}+\|b\|_{E}+\frac{1}{8}\|y_{n}\|_{E}+\frac{1}{8}\|x_{n}\|_{\dot{E}}+\frac{9}{8}\|f(x_{n},y_{n})\|_{E}+\frac{1}{8r}(\|f(x_{n},y_{n})\|_{X}^{2})$
$\leq$ $2\|a\|_{\dot{E}}+\|b\|_{E}+\frac{1}{8}\|y_{n}\|_{E}+\frac{1}{8}\|x_{n}\|_{\dot{E}}+\frac{9}{8}(\|x_{n}\|_{E}+c\|y_{n}\|_{E})+\frac{1}{8r}(\|x_{n}\|_{E}+c\|y_{n}\|_{E})^{2}$
$\leq$ $\frac{r}{4}+\frac{r}{2}+\frac{r}{4}+\frac{9}{8}(2r+4cr)+\frac{1}{8r}(4r^{2}+16cr^{2}+16c^{2}r^{2})$
$\leq$ $\frac{15}{4}r+(\frac{13}{2}c+2c^{2})r<4r,$
for $0<c<\tilde{c}=\frac{1}{30}$. Noting that $(x_{0},y_{0})\in B(0,\frac{r}{4})\subset B(0,4r)$, we obtain that $(x_{n},y_{n})\in B(0,4r)$ for all $n\in\mathbb{N}$. Setting $f_{n,m}=f(x_{n},y_{n})-f(x_{m},y_{m})\in E$, for all $n,m\in\mathbb{N}$, we have
$\|f_{n,m}\|_{X}\leq\|x_{n}-x_{m}\|_{\dot{E}}+c\|y_{n}-y_{m}\|_{E},\quad\|f(x_{n},y_{n})\|_{X}\leq\|x_{n}\|_{\dot{E}}+c\|y_{n}\|_{E}\leq 2r+4cr$
To prove that $\{(x_{n},y_{n})\}_{n\in\mathbb{N}}$ is a Cauchy sequence, we estimate that
$\|(x_{n+1},y_{n+1})-(x_{n},y_{n})\|_{\mathcal{E}}\leq$ $2\|x_{n+1}-x_{n}\|_{\dot{E}}+\|y_{n+1}-y_{n}\|_{E}$
$\leq$ $2\|g(y_{n}-y_{n-1})\|_{\dot{E}}+2\|h(x_{n}-x_{n-1})\|_{\dot{E}}+2\|\mathcal{G}(f_{n,n-1})\|_{\dot{E}}$
$+2\|\mathcal{B}(f_{n,n-1},f(x_{n},y_{n}))+\mathcal{B}(f(x_{n},y_{n}),f_{n,n-1})\|_{\dot{E}}+\|f_{n,n-1}\|_{E}$
$\leq$ $\frac{1}{8}\|y_{n}-y_{n-1}\|_{E}+\frac{1}{8}\|x_{n}-x_{n-1}\|_{\dot{E}}+\frac{1}{8r}(\|f_{n,n-1}\|_{X}\|f(x_{n},y_{n})\|_{X})$
$+\frac{1}{8r}(\|f_{n,n-1}\|_{X}\|f(x_{n-1},y_{n-1})\|_{X})+\frac{9}{8}\|f_{n,n-1}\|_{E}$
$\leq$ $(\frac{1}{8}+\frac{2cr+4c^{2}r}{4r}+\frac{9c}{8})\|y_{n}-y_{n-1}\|_{E}+(\frac{1}{8}+\frac{2r+4cr}{4r}+\frac{9}{8})\|x_{n}-x_{n-1}\|_{\dot{E}}$
$<$ $(\frac{7}{8}+\frac{13c}{8})\|(x_{n},y_{n})-(x_{n-1},y_{n-1})\|_{\mathcal{E}}<\frac{15}{16}\|(x_{n},y_{n})-(x_{n-1},y_{n-1})\|_{\mathcal{E}}.$
Then we get the limit of $(x_{n},y_{n})$ is a unique fixed point of mapping (2.8) in Banach space $\mathcal{E}$, which proves the lemma. ∎
3 Result in the range of $\gamma\in(\frac{4}{3},2]$
This section treats the highly dissipative regime, in which the linear component $X(t)$ enjoys improved regularity. Specifically, for $\gamma\in(\frac{4}{3},2]$, Lemma 2.7 yields $X\in C_{T}H^{\frac{1}{6}+\delta}$ for a sufficiently small $\delta$. This regime is subcritical because the critical regularity for $X$ is $H^{\frac{3}{2}-\gamma}.$
3.1 Bi-linear estimate
First we build a standard bilinearity estimate. Defining $B(f,g)$ as follows
$B(f,g)(t)=\int_{0}^{t}e^{(t-s)\Lambda^{\gamma}}\partial_{x}(f(s)g(s))\mathrm{d}s,$ (3.1)
then we have the following estimate.
Lemma 3.1.
Fixing $E_{T}:=L^{\infty}_{T}H^{s}$, with $0\leq\rho<\gamma-1$, $s=\sigma+\rho$, then we have
$\|B(f,g)\|_{E_{T}}\leq CT^{\frac{\gamma-1-\rho}{\gamma}}(1+T^{\frac{\rho}{\gamma}})\|fg\|_{C_{T}H^{\sigma}}.$ (3.2)
Proof.
By (3.1) and Parseval formula, we have
$\|B(f,g)(t)\|_{H^{s}}\lesssim$ $\int_{0}^{t}\|e^{-(t-\tau)\Lambda^{\gamma}}(1-\partial_{x}^{2})^{\frac{\rho}{2}}\partial_{x}(1-\partial_{x}^{2})^{\frac{\sigma}{2}}(fg)\|_{L^{2}}\mathrm{d}\tau$
$\lesssim$ $\int_{0}^{t}\|e^{-(t-\tau)|k|^{\gamma}}(1+|k|^{2})^{\frac{\rho}{2}}\partial_{x}(1+|k|^{2})^{\frac{\sigma}{2}}(fg)\|_{L^{2}}\mathrm{d}\tau$
Divide $k$ in two parts: $|k|\lesssim 1$ and $|k|\gtrsim 1$. For low frequencies $|k|\lesssim 1$, we have
$|e^{-(t-\tau)|k|^{\gamma}}(1+k^{2})^{\frac{\rho}{2}}k|\lesssim|e^{-(t-\tau)|k|^{\gamma}}|k||.$
For high frequencies $|k|\gtrsim 1$, we have
$|e^{-(t-\tau)|k|^{\gamma}}(1+k^{2})^{\frac{\rho}{2}}k|\lesssim|e^{-(t-\tau)|k|^{\gamma}}|k|^{\rho+1}|.$
Then for $t\in[0,T]$ we obtain that
$\|B(f,g)(t)\|_{H^{s}}\lesssim\int_{0}^{t}\max(\frac{1}{(t-\tau)^{\frac{1}{\gamma}}},\frac{1}{(t-\tau)^{\frac{1+\rho}{\gamma}}})\|fg\|_{H^{\sigma}}\mathrm{d}\tau\lesssim T^{\frac{\gamma-1-\rho}{\gamma}}(1+T^{\frac{\rho}{\gamma}})\|fg\|_{C_{T}H^{\sigma}},$
which finish our proof. ∎
Remark 3.1.
The endpoint case $\rho=\gamma-1$ requires a more refined analysis, typically involving frequencies localization in Besov spaces $B_{2,\infty}^{s}$. Such endpoint issues are not addressed in the present work. Indeed, for any $\alpha<\frac{\gamma}{2}-\frac{1}{2}$, one has $X\in\mathcal{W}^{\alpha+}(T)$ almost surely, which suffices for our purposes.
By Moser-type estimate, it’s not difficult to prove that
Lemma 3.2.
Let $\gamma\in(\frac{4}{3},2]$, $\alpha>\frac{1}{6}$, $\delta$ satisfy $0<2\delta\leq\alpha+\gamma-\frac{3}{2}$, then for $T\leq 1$ we have the following bounds
  • (1) For $f,g\in L^{\infty}_{T}H_{x}^{\frac{1}{2}+\delta}$, we have
    $\|B(f,g)\|_{L^{\infty}_{T}H^{\frac{1}{2}+\delta}}\leq C_{1}T^{\frac{\gamma-1}{\gamma}}\|f\|_{L^{\infty}_{T}H^{\frac{1}{2}+\delta}}\|g\|_{L^{\infty}_{T}H^{\frac{1}{2}+\delta}}.$
  • (2) For $f,g\in L^{\infty}_{T}W^{\alpha}$, we have
    $\|B(f,g)\|_{L^{\infty}_{T}H^{\frac{1}{2}+\delta}}\leq C_{2}T^{\frac{\delta}{\gamma}}\|f\|_{C_{T}W^{\alpha}}\|g\|_{C_{T}W^{\alpha}}.$
  • (3) For $f\in L_{T}^{\infty}{H^{\frac{1}{2}+\delta}}$ and $g\in L_{T}^{\infty}W^{\alpha}$, then we have
    $\|B(f,g)\|_{L^{\infty}_{T}H^{\frac{1}{2}+\delta}}\leq C_{3}T^{\frac{\delta}{\gamma}}\|f\|_{L^{\infty}_{T}H^{\frac{1}{2}+\delta}}\|g\|_{L^{\infty}_{T}H^{\alpha}}.$
3.2 Well-posedness
In this section, we study the well-posedness of equation (1.1) for $\gamma>\frac{4}{3}$. We begin by considering the equation
$\partial_{t}v-\Lambda^{\gamma}v=\partial_{x}(v^{2})+2\partial_{x}(v\tilde{X})+\partial_{x}(\tilde{X}^{2}),\quad u(0)=u_{0}.$ (3.3)
and solve it via a contraction mapping argument.
Proposition 3.1.
Let $\alpha>\frac{1}{6}$, $\gamma\in(\frac{4}{3},2]$ and $u_{0}\in H^{\frac{1}{2}+\delta}$ for $2\delta\leq\alpha+\gamma-\frac{3}{2}$. If $\tilde{X}(t)\in{C_{\bar{T}}W^{\alpha}}$ for some $\bar{T}>0$ , then there exists a unique solution $v(t,x)$ of equation (3.3) with a postive $T^{*}$ satisfying $v\in C_{T^{*}}H^{\frac{1}{2}+\delta}$.
Proof.
Let $r=2\max(\|u_{0}\|_{H^{\frac{1}{2}+\delta}},\|\tilde{X}\|_{C_{\bar{T}}W^{\alpha}})$. We rewrite equation (3.3) as
$v(t,x)=P(t)u_{0}+B(v,\tilde{X})+B(\tilde{X},v)+B(\tilde{X},\tilde{X})+B(v,v),$
where $P(t)f$ satisfying linear equation
$\partial_{t}P(t)f-\Lambda^{\gamma}P(t)f=0,\quad P(0)f=f.$ (3.4)
Firstly, we claim that there exist $0<T^{*}\leq 1\wedge\bar{T}$ such that for any $t\in[0,T^{*}]$
$a:=P(t)u_{0}+B(\tilde{X},\tilde{X})\in B(0,r),\quad\|B\|_{C_{t}H^{\frac{1}{2}+\delta}}\leq\frac{1}{8r},\quad\sup_{v}\frac{2\|B(v,\tilde{X})\|_{C_{t}H^{\frac{1}{2}+\delta}}}{\|v\|_{C_{t}H^{\frac{1}{2}+\delta}}}\leq\frac{1}{4}.$ (3.5)
For the first claim of (3.5), fixing $0<T\leq 1\wedge\bar{T}$ small enough to make sure that $C_{2}T^{\frac{\delta}{\gamma}}\|\tilde{X}\|^{2}_{C_{\bar{T}}W^{\alpha}}\leq\frac{r}{4}$, then by Lemma 3.2, we have
$\|a\|_{C_{T}H^{\frac{1}{2}+\delta}}\leq$ $\|P(\cdot)u_{0}\|_{C_{T}H^{\frac{1}{2}+\delta}}+\|B(\tilde{X},\tilde{X})\|_{C_{T}H^{\frac{1}{2}+\delta}}$
$\leq$ $\|u_{0}\|_{C_{T}H^{\frac{1}{2}+\delta}}+C_{2}T^{\frac{\delta}{\gamma}}\|\tilde{X}\|^{2}_{C_{\bar{T}}W^{\alpha}}\leq r.$
For the second claim of (3.5), fixing $0<T\leq 1$ small enough to make sure that $C_{1}T^{\frac{\gamma-1}{\gamma}}\leq\frac{1}{8r}$, then by Lemma 3.2, we have
$\|B(v,v)\|_{L^{\infty}_{T}H^{\frac{1}{2}+\delta}}\leq$ $C_{1}T^{\frac{\gamma-1}{\gamma}}\|v\|_{C_{T}H^{\frac{1}{2}+\delta}}\|v\|_{C_{T}H^{\frac{1}{2}+\delta}}$
$\leq$ $\frac{1}{8r}\|v\|_{C_{T}H^{\frac{1}{2}+\delta}}^{2}.$
For the third claim of (3.5), fixing $0<T\leq 1\wedge\bar{T}$ small enough to make sure that $C_{3}T^{\frac{\delta}{\gamma}}\|\tilde{X}\|_{C_{\bar{T}}W^{\alpha}}\leq\frac{1}{8}$, then by Lemma 3.2, we have
$\frac{\|B(v,\tilde{X})\|_{C_{T}H^{s}}}{\|v\|_{C_{T}H^{s}}}\leq$ $C_{3}T^{\frac{\delta}{\gamma}}\|\tilde{X}\|_{C_{\bar{T}}W^{\alpha}}\leq\frac{1}{8}.$
Let $T^{*}\leq\bar{T}$ satisfy all the condition above, we prove our claim is hold on $0,T^{*}]$. By Lemma 2.8, we finish our proof. ∎
Proof of Theorem 1.1.
Let $\gamma\in(\frac{4}{3},2]$, by Lemma 2.7, we have $X(t)\in C_{T^{*}}W^{\alpha}$ for some $\alpha\in(\frac{1}{6},\frac{\gamma}{2}-\frac{1}{2})$ where $X(t)$ satisfies (1.5). By Proposition 3.1, then we finish our proof by writing equation (1.1) as the (3.3) with fixing v=uXv=u-X and $\tilde{X}(t)=X(t)=\int_{0}^{t}P(t-s)\xi(s)\mathrm{d}s$. ∎
Remark 3.2.
While we prove the existence of the solution in the space $L_{t}^{\infty}H^{s}$, the equation’s structure also allow the continuity of $v$ in time to be readily established. Observing that
$v(t)-v(s)=(S(t)-S(s))u_{0}+B_{s,t}(v+X,v+X),$
where $B_{s,t}(u,u):=\int_{s}^{t}P(t-s)D(u^{2})\mathrm{d}s$, Using the similar argument for proving well-posedness, we can easily get the continuous.
To obtain a global solution, further energy estimates are required. A classical energy estimate for the heat equation (see [BCD11]) is applied to derive the corresponding estimate for the rough Burgers equation.
Lemma 3.3.
Let $\gamma\in(1,2]$. Let $v$ be the solution in $\mathcal{C}([0,T];\mathcal{S}^{\prime}(\mathbb{T}^{d}))$ of the Cauchy problem
[…]
with $f\in L^{2}([0,T];\dot{H}^{s-1})$ and $v_{0}\in\dot{H}^{s}(\mathbb{T}^{d})$. Then,
$v\in(\cap_{p=2}^{\infty}L^{p}([0,T];\dot{H}^{s+\frac{\gamma}{p}}))\cap\mathcal{C}([0,T];\dot{H}^{s}).$
Moreover we have the following estimates:
$\|v(t)\|_{\dot{H}^{s}}^{2}+2\int_{0}^{t}\|\Lambda^{\frac{\gamma}{2}}v(t^{\prime})\|_{\dot{H}^{s}}^{2}\mathrm{d}t^{\prime}=$ $\|v_{0}\|_{\dot{H}^{s}}^{2}+2\int_{0}^{t}\langle f(t^{\prime}),v(t^{\prime})\rangle_{s}\mathrm{d}t^{\prime},$
$(\sum_{k\neq 0}\langle k\rangle^{2s}(\sup_{0\leq t^{\prime}\leq t}|\hat{v}(t^{\prime},\xi)|)^{2})^{\frac{1}{2}}\leq$ $\|v_{0}\|_{\dot{H}^{s}}+\|f\|_{L_{T}^{2}(\dot{H}^{s-\frac{\gamma}{2}})},$
$\|v(t)\|_{L_{T}^{p}(\dot{H}^{s+\frac{\gamma}{p}})}\leq$ $\|v_{0}\|_{\dot{H}^{s}}+\|f\|_{L_{T}^{2}(\dot{H}^{s-\frac{\gamma}{2}})},$
with $\langle a,b\rangle_{s}=\int\langle\xi\rangle^{2s}\hat{a}(\xi)\overline{\hat{b}(\xi)}\mathrm{d}\xi$
Proof.
The case of $\gamma=2$ is proved in [BCD11], the extension to the case $\gamma\in(1,2]$ is straightforward. ∎
Next, we build the well-posedness in $L^{4}_{T}\dot{H}^{s+\frac{\gamma}{4}}$, for $0\leq s<\gamma-\frac{3}{2}$. The key observation is that for $s_{1}\leq s_{2}$, we have
$\|\partial_{x}f\|_{\dot{H}^{s_{1}}}\leq\|\partial_{x}f\|_{\dot{H}^{s_{2}}}.$ (3.7)
Proposition 3.2.
Let $u_{0}\in\dot{H}^{s}$ for $s\in[0,\gamma-\frac{3}{2})$ with $\gamma\in(\frac{8}{5},2]$, then there exists a constant $C$ such that if the following conditions
$CT_{0}^{\frac{1}{4}}\|\tilde{X}\|_{C_{\bar{T}}W^{\alpha}}<\frac{1}{4}\quad for\;0<T_{0}\leq\bar{T}\quad and\quad\alpha\geq\max(\frac{3}{2}-\frac{3\gamma}{4},s+1-\frac{\gamma}{2}),$ (3.8)
hold, then there exists a positive time $T$ such that (3.3) has a unique solution in $L_{T}^{4}([0,T];\dot{H}^{s+\frac{\gamma}{4}})$. Moreover let $T_{u_{0}}$ denote the maximal time of existence of such a solution, then
  • For the same constant $C$,
    $\|u_{0}\|_{\dot{H}^{s}}\leq\frac{1}{16C}\Longrightarrow T_{u_{0}}=T_{0}.$
  • If $T_{u_{0}}\leq T_{0}$, then
    $\int_{0}^{T_{u_{0}}}\|u(t)\|_{\dot{H}^{s+\frac{\gamma}{4}}}^{4}dt=\infty.$ (3.9)
Proof.
We firstly give the local result for small initial data $u_{0}$. Noting that $B(f,g)$ is the solution of
$\partial_{t}B(f,g)-\Lambda^{\gamma}B(f,g)=\partial_{x}(fg).$
Let $\gamma>\frac{8}{5}$ and $0\leq s<\gamma-\frac{3}{2}$, by Lemma 3.3, (3.7) and Lemma 2.1, we have
$\|P(t)u_{0}+B(\tilde{X},\tilde{X})\|_{L_{T}^{4}\dot{H}^{s+\frac{\gamma}{4}}}\leq$ $\|u_{0}\|_{\dot{H}^{s}}+CT^{\frac{1}{2}}\|\tilde{X}\|_{C_{T}W^{\alpha}}^{2},$ (3.10)
$2\|B(v,\tilde{X})\|_{L_{T}^{4}\dot{H}^{s+\frac{\gamma}{4}}}\leq$ $CT^{\frac{1}{4}}\|\tilde{X}\|_{C_{T}W^{\alpha}}\|v\|_{L_{T}^{4}\dot{H}^{s+\frac{\gamma}{4}}},$ (3.11)
$\|B(v,v)\|_{L_{T}^{4}\dot{H}^{s+\frac{\gamma}{4}}}\leq$ $C\|v\|_{L_{T}^{4}\dot{H}^{s+\frac{\gamma}{4}}}^{2}.$ (3.12)
Let $r=\frac{1}{8C}$, we know that if $\|u_{0}\|_{\dot{H}^{s}}\leq\frac{1}{16C}$ and
$CT_{0}^{\frac{1}{4}}\|\tilde{X}\|_{C_{\bar{T}}W^{\alpha}}<\frac{1}{4},\quad$
by Lemma 2.8, there exists a unique solution of (3.3) in the ball with center 0 and radius $\frac{1}{4C}$ in the space $L^{4}([0,T_{0}];\dot{H}^{s+\frac{\gamma}{4}})$. Now we consider the case of large initial data $u_{0}$ in $\dot{H}^{s}$. We split $u_{0}$ into small part in $\dot{H}^{s}$ and a large part with compactly supported Fourier transform. For that we fix some positive real number $\rho_{u_{0}}$ such that
$(\sum_{|k|\geq\rho_{u_{0}}}|k|^{2s}|\widehat{u}_{0}(k)|^{2})^{\frac{1}{2}}\leq\frac{1}{32C}.$
Then by $\|P(t)u_{0}\|_{L_{T}^{4}\dot{H}^{s+\frac{\gamma}{4}}}\leq\|u_{0}\|_{\dot{H}^{s}}$ and $u_{0}^{\flat}=\mathcal{F}^{-1}(1_{B(0,\rho_{u_{0}})}\widehat{u}_{0})$, we get
$\|P(t)u_{0}\|_{L_{T}^{4}\dot{H}^{s+\frac{\gamma}{4}}}\leq$ $\frac{1}{32C}+\|P(t)u^{\flat}_{0}\|_{L_{T}^{4}\dot{H}^{s+\frac{\gamma}{4}}}$
$\leq$ $\frac{1}{32C}+\rho_{u_{0}}^{\frac{\gamma}{4}}\|P(t)u^{\flat}_{0}\|_{L_{T}^{4}H^{s}}$
$\leq$ $\frac{1}{32C}+(\rho_{u_{0}}^{\gamma}T)^{\frac{1}{4}}\|u_{0}\|_{\dot{H}^{s}}.$
Thus if
$T\leq\left(\frac{1}{\rho_{u_{0}}^{\frac{\gamma}{4}}32C\|u_{0}\|_{\dot{H}^{s}}}\right)^{4}\wedge T_{0},$ (3.13)
then we have the existence of a unique solution in the ball with center 0 and radius $\frac{1}{4C}$ in the space $L_{T}^{4}\dot{H}^{s+\frac{\gamma}{4}}$. Noting that if $v$ is a solution in $L^{4}([0,T];\dot{H}^{s+\frac{\gamma}{4}})$, by Lemma 3.3, we have $v\in\mathcal{C}([0,T];\dot{H}^{s})\cap L^{2}([0,T];\dot{H}^{s+\frac{\gamma}{2}})$
Finally, we prove the blow-up criterion. Assume that we have a solution $v$ of (3.3) on the time interval $[0,T[$ such that
$\int_{0}^{T}\|v(t)\|_{\dot{H}^{s+\frac{\gamma}{4}}}^{4}dt<\infty,\quad T<T_{0}<\bar{T}.$
We claim that the lifespan $T_{u_{0}}$ of $v$ is greater than $T$. Indeed, thanks to Lemma 3.3, we have
$\sum_{k\neq 0}|k|^{2s}(\sup_{t\in[0,T[}|\widehat{v}(t,k)|)^{2}\leq$ $\|v_{0}\|_{H^{s}}+\|\partial_{x}(v^{2}+2v\tilde{X}+\tilde{X}^{2})\|_{\dot{H}^{s-\frac{\gamma}{2}}}$
$\leq$ $\|v_{0}\|_{H^{s}}+C\|v\|_{L_{T}^{4}\dot{H}^{s+\frac{\gamma}{4}}}^{2}+CT^{\frac{1}{4}}\|\tilde{X}\|_{C_{\bar{T}}W^{\alpha}}\|v\|_{L_{T}^{4}\dot{H}^{s+\frac{\gamma}{4}}}+CT^{\frac{1}{2}}\|\tilde{X}\|_{C_{\bar{T}}W^{\alpha}}^{2}$
$\leq$ $\|v_{0}\|_{H^{s}}+C\|v\|_{L_{T}^{4}\dot{H}^{s+\frac{\gamma}{4}}}^{2}+\frac{1}{4}\|v\|_{L_{T}^{4}\dot{H}^{s+\frac{\gamma}{4}}}+\frac{1}{16C}<\infty.$
Thus, there exists a positive number $\rho$ exists such that
$\forall t\in[0,T[,\quad\sum_{|k|\geq\rho}|k|^{2s}(\sup_{t\in[0,T[}|\widehat{v}(t,k)|)^{2}\leq\frac{1}{32C}.$
Noting that the choise of $\rho$ is independent of $t$, by condition of (3.13), we can get $T_{0}\geq T_{u_{0}}>T$ which finish our proof. ∎
Proposition 3.3.
Let $\gamma\in(\frac{3}{2},2]$, $\alpha\in(1-\frac{\gamma}{2},\frac{\gamma}{2}-\frac{1}{2})$ and $0<T\leq\bar{T}$, let $v$ be the smooth solution of (3.3) on $[0,T]$ for $u_{0}\in L^{2}$ with $\tilde{X}\in C_{\bar{T}}W^{\alpha}$. Then for any $t\in[0,T]$, we have the following energy estimate
$\|v\|_{L^{2}}^{2}+2(1-\nu)\int_{0}^{t}\|v(t^{\prime})\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}\leq\|u_{0}\|_{L^{2}}^{2}e^{C_{\nu}(1+\|\tilde{X}\|_{C_{\bar{T}}W^{\alpha}}^{\frac{2\gamma}{\gamma-1}})t}+C_{\nu}\|\tilde{X}\|_{C_{\bar{T}}W^{\alpha}}^{4}\int_{0}^{t}e^{C_{\nu}(1+\|\tilde{X}\|_{C_{\bar{T}}W^{\alpha}}^{\frac{2\gamma}{\gamma-1}})(t-t^{\prime})}\mathrm{d}t^{\prime}.$ (3.14)
for some small $\nu<1$.
Proof.
Observing that $v$ satisfies (3.6) with $f=\partial_{x}(v+\tilde{X})^{2}$. By Lemma 3.3 and integrating by parts, we have
$\|v(t)\|_{L^{2}}^{2}+2\int_{0}^{t}\|v(t^{\prime})\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}\leq\|u_{0}\|_{L^{2}}^{2}+2\int_{0}^{t}\|v(t^{\prime})\|_{L^{2}}^{2}\mathrm{d}t^{\prime}+2\int_{0}^{t}\langle v(t^{\prime}),2\partial_{x}(v(t^{\prime})\tilde{X}(t^{\prime}))+\partial_{x}(\tilde{X}^{2}(t^{\prime}))\rangle_{0}\mathrm{d}t^{\prime}.$
By Hölder inequality, Young’s inequality, Lemma 2.1 and Lemma 2.7, there exists $\alpha\in(1-\frac{\gamma}{2},\frac{\gamma}{2}-\frac{1}{2})$ such that
$\langle v,2\partial_{x}(v\tilde{X})+\partial_{x}(\tilde{X}^{2})\rangle_{0}\leq$ $2\|v\|_{H^{\frac{\gamma}{2}}}\|v\tilde{X}\|_{H^{1-\frac{\gamma}{2}}}+\|v\|_{H^{\frac{\gamma}{2}}}\|\tilde{X}^{2}\|_{H^{1-\frac{\gamma}{2}}}$
$\leq$ $2\|v\|_{H^{\frac{\gamma}{2}}}^{1+\frac{1}{\gamma}}\|v\|_{L^{2}}^{1-\frac{1}{\gamma}}\|\tilde{X}\|_{W^{\alpha}}+\|v\|_{H^{\frac{\gamma}{2}}}\|\tilde{X}\|_{W^{\alpha}}^{2}$
$\leq$ $\frac{\nu}{2}\|v\|_{H^{\frac{\gamma}{2}}}^{2}+C_{\nu}(\|v\|_{L^{2}}^{2}\|\tilde{X}\|^{\frac{2\gamma}{\gamma-1}}_{W^{\alpha}}+\|\tilde{X}\|_{W^{\alpha}}^{4}).$
For $0\leq t\leq T$, we have
$\|v(t)\|_{L^{2}}^{2}+2\int_{0}^{t}\|v(t^{\prime})\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}\leq$ $\|u_{0}\|_{L^{2}}^{2}+\nu\int_{0}^{t}\|v(t^{\prime})\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}+C_{\nu}\int_{0}^{t}(1+\|\tilde{X}(t^{\prime})\|_{W^{\alpha}}^{\frac{2\gamma}{\gamma-1}})\|v(t^{\prime})\|_{L^{2}}^{2}\mathrm{d}t^{\prime}$ (3.15)
$+C_{\nu}\int_{0}^{t}\|\tilde{X}(t^{\prime})\|_{W^{\alpha}}^{4}\mathrm{d}t^{\prime}.$
Observing that $\|\tilde{X}\|_{C_{t}W^{\alpha}}\leq\|\tilde{X}\|_{C_{\bar{T}}W^{\alpha}}$, for $\nu<1$, by Gronwall’s inequality, the proposition is proved. ∎
Furthermore, the following lemma shows the estimate of $v$ can be improved to higher regularity.
Lemma 3.4.
Let $\gamma\in(\frac{3}{2},2]$, $\alpha\in(s+1-\frac{\gamma}{2},\frac{\gamma}{2}-\frac{1}{2})$ , $0<s<\gamma-\frac{3}{2}$ and $0<T\leq\bar{T}$. Let $v$ be the smooth solution of (3.3) on $[0,T]$ for $u_{0}\in H^{s}$ with $\tilde{X}\in C_{\bar{T}}W^{\alpha}$. For $t\in[0,T]$, we have the following energy estimate
$\|\Lambda^{s}v\|_{L^{2}}+2(1-\nu)\int_{0}^{t}\|\Lambda^{s}v\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}\leq$ $\|u_{0}\|^{2}_{H^{s}}e^{C_{\nu}\int_{0}^{t}\|v\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}}+C_{\nu}\|\tilde{X}\|_{C_{\bar{T}}W^{\alpha}}^{2}\int_{0}^{t}e^{C_{\nu}\int_{t^{\prime}}^{t}\|v\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime\prime}}\|v\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}$ (3.16)
$+C_{\nu}\|\tilde{X}\|_{C_{\bar{T}}W^{\alpha}}^{4}\int_{0}^{t}e^{C_{\nu}\int_{t^{\prime}}^{t}\|v\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime\prime}}\mathrm{d}t^{\prime}$
for some small $\nu<1$. Moreover, $v\in C_{T}H^{s}\cap L_{T}^{2}H^{s+\frac{\gamma}{2}}$.
Proof.
Recalling $\Lambda^{s}v$ satisfies the following equation
$\partial_{t}(\Lambda^{s}v)-\Lambda^{\gamma}(\Lambda^{s}v)=\Lambda^{s}\partial_{x}(v^{2})+2\Lambda^{s}\partial_{x}(v\tilde{X})+\Lambda^{s}\partial_{x}(\tilde{X}^{2}),$ (3.17)
multiplying $\Lambda^{s}v$ both sides and integrating on $\mathbb{T}$, we have
$\frac{1}{2}\partial_{t}\|\Lambda^{s}v\|_{L^{2}}^{2}+\|\Lambda^{s}v\|_{\dot{H}^{\frac{\gamma}{2}}}^{2}=\langle\Lambda^{s}\partial_{x}(v^{2}),\Lambda^{s}v\rangle+2\langle\Lambda^{s}\partial_{x}(v\tilde{X}),\Lambda^{s}v\rangle+\langle\Lambda^{s}\partial_{x}(\tilde{X}^{2}),\Lambda^{s}v\rangle.$ (3.18)
For the first term in the right side of (3.18), by Parsavel formula and Hölder’s inequality, we have
$\langle\Lambda^{s}\partial_{x}(v^{2}),\Lambda^{s}v\rangle\leq$ $\|v\|_{\dot{H}^{s+\frac{\gamma}{2}}}\|D^{{s+1-\frac{\gamma}{2}}}(v^{2})\|_{L^{2}}$
$\leq$ $\|v\|_{\dot{H}^{s+\frac{\gamma}{2}}}\|D^{{s+1-\frac{\gamma}{2}}}v^{2}-2vD^{s+1-\frac{\gamma}{2}}v\|_{L^{2}}+\|v\|_{\dot{H}^{s+\frac{\gamma}{2}}}\|2vD^{s+1-\frac{\gamma}{2}}v\|_{L^{2}}.$
By the estimate in [KPV93, GO14] and the interpolation, for $0<s+1-\frac{\gamma}{2}<1$ and $\beta=\frac{s}{2}+\frac{1}{2}-\frac{\gamma}{4}$, we have
$\|D^{{s+1-\frac{\gamma}{2}}}v-2vD^{s+1-\frac{\gamma}{2}}v\|_{L^{2}}\lesssim$ $\|D^{\beta}v\|_{L^{4}}^{2}$
$\lesssim$ $\|v\|_{\dot{H}^{s}}\|v\|_{\dot{H}^{\frac{\gamma}{2}}},$
for $\gamma\geq\frac{3}{2}$. By Hölder’s inequality and Sobolev embedding in [BCD11], we have
$\|2vD^{s+1-\frac{\gamma}{2}}v\|_{L^{2}}\lesssim$ $\|v\|_{L^{\frac{2}{1-2s}}}\|D^{s+1-\frac{\gamma}{2}}v\|_{L^{\frac{2}{2s}}}$
$\lesssim$ $\|v\|_{\dot{H}^{s}}\|D^{s+1-\frac{\gamma}{2}}v\|_{\dot{H}^{\frac{1}{2}-s}}$
$\lesssim$ $\|v\|_{\dot{H}^{s}}\|v\|_{\dot{H}^{\frac{\gamma}{2}}},$
for $s<\frac{1}{2}$ and $\gamma>\frac{3}{2}$. Then we have
$\langle\Lambda^{s}\partial_{x}(v^{2}),\Lambda^{s}v\rangle\leq\frac{\nu}{3}\|v\|_{\dot{H}^{s+\frac{\gamma}{2}}}^{2}+C_{\nu}\|v\|_{\dot{H}^{s}}^{2}\|v\|_{\dot{H}^{\frac{\gamma}{2}}}^{2}.$
For the second term and third term of right hand side of (3.18), by Parsavel formula Hölder’s inequality and Lemma 2.1, for $s+1-\frac{\gamma}{2}\leq\alpha<\frac{\gamma}{2}-\frac{1}{2}$, we have
$2\langle\Lambda^{s}\partial_{x}(vX),\Lambda^{s}v\rangle\leq$ $\|vX\|_{\dot{H}^{s+1-\frac{\gamma}{2}}}\|v\|_{\dot{H}^{s+\frac{\gamma}{2}}}$
$\leq$ $\|X\|_{W^{\alpha}}\|v\|_{H^{s+\frac{3}{2}-\frac{\gamma}{2}}}\|v\|_{\dot{H}^{s+\frac{\gamma}{2}}}$
$\leq$ $\|X\|_{W^{\alpha}}\|v\|_{H^{\frac{\gamma}{2}}}\|v\|_{\dot{H}^{s+\frac{\gamma}{2}}}$
$\leq$ $C_{\nu}\|X\|_{W^{\alpha}}^{2}\|v\|^{2}_{H^{\frac{\gamma}{2}}}+\frac{\nu}{3}\|v\|_{\dot{H}^{s+\frac{\gamma}{2}}}^{2},$
and
$\langle\Lambda^{s}\partial_{x}(X^{2}),\Lambda^{s}v\rangle\leq$ $\|X^{2}\|_{\dot{H}^{s+1-\frac{\gamma}{2}}}\|v\|_{\dot{H}^{s+\frac{\gamma}{2}}}$
$\leq$ $C_{\nu}\|X\|_{W^{\alpha}}^{2}\|v\|_{\dot{H}^{s+\frac{\gamma}{2}}}$
$\leq$ $C_{\nu}\|X\|_{W^{\alpha}}^{4}+\frac{\nu}{3}\|v\|_{\dot{H}^{s+\frac{\gamma}{2}}}^{2}.$
Combining the above estimates, and integral (3.18) on $[0,t]$, we have
$\frac{1}{2}\|\Lambda^{s}v(t)\|_{L^{2}}^{2}+(1-\nu)\int_{0}^{t}\|\Lambda^{s}v(t^{\prime})\|_{\dot{H}^{\frac{\gamma}{2}}}\mathrm{d}t^{\prime}\leq$ $\frac{1}{2}\|u_{0}\|_{H^{s}}^{2}+C_{\nu}\int_{0}^{t}\|X(t^{\prime})\|_{W^{\alpha}}^{2}\|v(t^{\prime})\|_{\dot{H}^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}+C_{\nu}\int_{0}^{t}\|X(t^{\prime})\|_{W^{\alpha}}^{4}\mathrm{d}t^{\prime}$
$+C_{\nu}\int_{0}^{t}\|v(t^{\prime})\|_{\dot{H}^{s}}^{2}\|v(t^{\prime})\|_{\dot{H}^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}.$
Since $\int_{0}^{t}\|v\|_{\dot{H}^{\frac{\gamma}{2}}}\mathrm{d}t^{\prime}$ is uniformly bounded by constant depending only on $\|u_{0}\|_{L^{2}}$ and $\|X\|_{C_{\bar{T}}W^{\alpha}}$. By Gronwall’s inequality and $H^{s}\subset\dot{H}^{s}$ for $s\geq 0$, we prove (3.16) for all $v\in C_{T}L^{2}\cap L_{T}^{2}H^{\frac{\gamma}{2}}$. ∎
Now we prove that $v$ given by Proposition 3.2 satisfied the energy estimates.
Lemma 3.5.
Let $u_{0}\in\dot{H}^{\sigma}$ for $\sigma>0$, $\gamma>\frac{5}{3}$, then for $\tilde{X}=X$ satisfying (1.5), there exists a unique weak solution $v$ of equation (3.3) satisfying energy estimate (3.14) and blow up criterion (3.9).
Proof.
Noting that $X$ satisfies (1.5), by Proposition 2.7, we have $X\in C_{T}W^{\alpha}$ for any $\alpha\in(\frac{1}{3},\frac{\gamma}{2}-\frac{1}{2})$. Then for $\gamma>\frac{5}{3}$ and $s\in[0,\gamma-\frac{3}{2})$, we can choose $\alpha$ satisfying
$\alpha\geq\max(\frac{3}{2}-\frac{3\gamma}{4},s+1-\frac{\gamma}{2}).$
Choose a function $\phi:\mathbb{R}\mapsto\mathbb{R}_{+}$ that is even, smooth, compactly supported, decreasing on $\mathbb{R}_{+}$, and such that $\phi(0)=1$, and then set
$\xi^{\epsilon}(x)=\mathcal{F}^{-1}(\phi(\epsilon k)\mathcal{F}(\xi)(k))(x),\quad u_{0}^{\epsilon}(x)=\mathcal{F}^{-1}(\phi(\epsilon k)\mathcal{F}(u_{0})(k))(x).$
Let $X^{\epsilon}$ solve the equation
$\partial_{t}X^{\epsilon}-\Lambda^{\gamma}X^{\epsilon}=\xi^{\epsilon},\quad X^{\epsilon}(0)=0.$
Without loss of generality, we assume that $\|X^{\epsilon}\|_{C_{T}W^{\alpha}}\leq\|X\|_{C_{T}W^{\alpha}}$ and $\|u^{\epsilon}_{0}\|_{\dot{H}^{s}}\leq\|u_{0}\|_{\dot{H}^{s}}$. Let $v^{\epsilon}$ be the solution of equation
$\partial_{t}v^{\epsilon}-\Lambda^{\gamma}v^{\epsilon}=\partial_{x}(v^{\epsilon})^{2}+2\partial_{x}(v^{\epsilon}X^{\epsilon})+\partial_{x}(X^{\epsilon})^{2},\quad v^{\epsilon}(0)=u_{0}^{\epsilon}.$ (3.19)
Then by Proposition 3.2, there exists $T_{u_{0}}$(Noting that $T_{u_{0}^{\epsilon}}\geq T_{u_{0}}$.) such that $v^{\epsilon}\in L^{4}_{T_{u_{0}}}\dot{H}^{s+\frac{\gamma}{4}}$ for any $s\geq 0$. Noting that $X^{\epsilon}\in C_{T}W^{\infty}$, $u_{0}^{\epsilon}\in\dot{H}^{\infty}$, by Lemma 3.3, we have $v^{\epsilon}\in C_{T}\dot{H}^{s}\cap L_{T}^{2}\dot{H}^{s+\frac{\gamma}{2}}$ for any $s\geq 0$. Choosing $s=0$, then by Banach–Alaoglu Theorem, Fatou’s lemma([BCD11]) and Proposition 3.3, there exists $\psi:\mathbb{R}_{+}\mapsto\mathbb{R}_{+}$ and a subsequence $v_{\psi(\epsilon)}$ satisfying
$\lim_{\epsilon\to 0}\psi(\epsilon)\to 0,\quad\lim_{\epsilon\to 0}v_{\psi(\epsilon)}\rightharpoonup v\quad in\>L^{2},\quad\|v\|_{C_{T}L^{2}}^{2}+\int_{0}^{t}\|v\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}s\leq C(\|u_{0}\|_{L^{2}},\|X\|_{C_{T}W^{\alpha}}).$
To prove $v$ is the solution of (3.3), we need the strong convergence of $v_{\psi(\epsilon)}$ in $L^{2}$ to make sure that $\partial_{x}(v_{\psi(\epsilon)})^{2}$ can be weak-converge to $\partial_{x}(v^{2})$(it’s already valid for $\partial_{x}(X^{\epsilon})^{2}$ and $\partial_{x}(X^{\epsilon}v^{\epsilon})$). We use the Aubin-Lions-Simon Theorem in [Sim86]. It suffices to show the following two conditions
$\partial_{t}v^{\epsilon}\in L_{T}^{p}(H^{-\sigma}),\quad and\quad v^{\epsilon}\in C_{T}H^{s},$ (3.20)
uniformly hold for some $\sigma,s>0$, $p>1$. Observing that $\partial_{t}v^{\epsilon}=\Lambda^{\gamma}v^{\epsilon}+\partial_{x}(v^{\epsilon})^{2}+2\partial_{x}(v^{\epsilon}X^{\epsilon})+\partial_{x}(X^{\epsilon})^{2}$, by Lemma 2.1 and Proposition 3.3, we have $\partial_{t}v^{\epsilon}$ is uniformly bounded for any $\sigma\geq 1$. The second condition of (3.20) is a consequence of Lemma (3.4) when we choose $0<s\leq\sigma$.
Now we get a solution $v$ in $C_{T}L^{2}\cap L_{T}^{2}\dot{H}^{\frac{\gamma}{2}}$ for some $T>0$ and satisfying the energy estimate. Noting that
$\|v(t)\|_{L_{T}^{4}\dot{H}^{\frac{\gamma}{4}}}\leq\|v(t)\|_{L_{T}^{4}H^{\frac{\gamma}{4}}}\leq\|v(t)\|_{C_{T}L^{2}}^{\frac{1}{2}}\|v(t)\|_{L_{T}^{2}H^{\frac{\gamma}{2}}}^{\frac{1}{2}},$ (3.21)
we have that $v$ is also the solution in $L_{T}^{4}\dot{H}^{\frac{\gamma}{4}}$ with initial data in $L^{2}$. By the uniqueness of Proposition 3.2, we get $v$ also has the blow up criterion. Finally, we prove that the chooise of $v$ is independent by the choosing of mollifier $\phi$. It suffices to show the uniqueness of $v$ in $C_{T}L^{2}\cap L^{2}\dot{H}^{\frac{\gamma}{2}}$. Let $u_{0}$ belong to $H^{\sigma}$ for $\sigma>0$ and $v^{1}(t;u_{0})$ and $v^{2}(t;u_{0})$ be the solutions on $[0,T]$ given above and $\overline{v}:=v(t;u^{1}_{0})-v(t;u^{2}_{0})$, then $\overline{v}$ satisfies the equation
$\partial_{t}\overline{v}-\Lambda^{\gamma}\overline{v}=\partial_{x}(\overline{v}(v^{1}+v^{2}))+2\partial_{x}(\overline{v}X).$
Noting that $v^{1}$ and $v^{2}$ belong to $C_{T}L^{2}\cap L_{T}^{2}\dot{H}^{\frac{\gamma}{2}}$, multiplying $\bar{v}$ both sides and integrating, we obtain the estimate
$\|\overline{v}(t)\|_{L^{2}}^{2}+\int_{0}^{t}\|\overline{v}(t^{\prime})\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}\leq 2\int_{0}^{t}\|\overline{v}(t^{\prime})\|_{L^{2}}^{2}\mathrm{d}t^{\prime}+\int_{0}^{t}\langle\overline{v}(t^{\prime}),\partial_{x}(\overline{v}(t^{\prime})(v^{1}(t^{\prime})+v^{2}(t^{\prime})))+2\partial_{x}(\overline{v}(t^{\prime})X(t^{\prime}))\rangle_{0}\mathrm{d}t^{\prime}.$
By Hölder’s inequality, Parsavel formula, Lemma 2.1 and (3.7), for $\gamma>\frac{5}{3}$, we have
$\langle\overline{v},\partial_{x}(\overline{v}(v^{1}+v^{2}))\rangle_{0}\leq$ $\|\overline{v}\|_{\dot{H}^{\frac{\gamma}{2}}}\|\overline{v}(v^{1}+v^{2})\|_{\dot{H}^{\frac{\gamma}{2}-\frac{1}{2}}}$
$\leq$ $\|\overline{v}\|_{\dot{H}^{\frac{\gamma}{2}}}\|\overline{v}\|_{\dot{H}^{\frac{\gamma}{4}}}\|(v^{1}+v^{2})\|_{\dot{H}^{\frac{\gamma}{4}}}$
$\leq$ $\|\bar{v}\|^{\frac{3}{2}}_{\dot{H}^{\frac{\gamma}{2}}}\|\bar{v}\|^{\frac{1}{2}}_{L^{2}}\|v^{1}+v^{2}\|_{\dot{H}^{\frac{\gamma}{4}}}$
$\leq$ $\frac{\nu}{2}\|\bar{v}\|^{2}_{\dot{H}^{\frac{\gamma}{2}}}+\|\bar{v}\|^{2}_{L^{2}}\|v^{1}+v^{2}\|^{4}_{\dot{H}^{\frac{\gamma}{4}}}.$
Similarly, by the same estimate as in (3.11), we have
$\langle\overline{v},2\partial_{x}(\overline{v}X))\rangle_{0}\leq$ $2\|\overline{v}\|_{\dot{H}^{\frac{\gamma}{2}}}\|\overline{v}X\|_{\dot{H}^{1-\frac{\gamma}{2}}}$
$\leq$ $\|\overline{v}\|_{\dot{H}^{\frac{\gamma}{2}}}\|\overline{v}\|_{\dot{H}^{\frac{\gamma}{4}}}\|X\|_{W^{\alpha}}$
$\leq$ $\|\bar{v}\|^{\frac{3}{2}}_{\dot{H}^{\frac{\gamma}{2}}}\|\bar{v}\|^{\frac{1}{2}}_{L^{2}}\|X\|_{W^{\alpha}}$
$\leq$ $\frac{\nu}{2}\|\bar{v}\|^{2}_{\dot{H}^{\frac{\gamma}{2}}}+\|\bar{v}\|^{2}_{L^{2}}\|X\|^{4}_{W^{\alpha}}.$
Noting that $\int_{0}^{t}\|v^{1}\|_{\dot{H}^{\frac{\gamma}{4}}}^{4}dt$ and $\int_{0}^{t}\|v^{2}\|_{\dot{H}^{\frac{\gamma}{4}}}^{4}dt$ for $t\in[0,T]$ are finite, by the Gronwall lemma, we get $\|\overline{v}(t)\|_{L^{2}}^{2}+\int_{0}^{t}\|\overline{v}(t^{\prime})\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}\leq 0$ for any $t\in[0,T]$ until the blow-up time, which finishes our proof.
Proof of Theorem 1.2.
Noting that $u_{0}\in\dot{H}^{\sigma}$ for $\sigma>0$, let $v$ be the solution satisfying Lemma 3.5. To obtain the global solution, we proceed with our induction argument. Firstly, find a $T_{0}$ satisfying
$CT_{0}^{\frac{1}{4}}\|X\|_{C_{\bar{T}}W^{\alpha}}\leq\frac{1}{4}.$
We prove that for any $u_{0}\in H^{\sigma}$ for some $\sigma>0$, there exists a local solution on $0,T_{0}]$. Indeed, by Proposition 3.2, there exists $v$ on $0,T_{u_{0}}\wedge T_{0})$ satisfying energy estimate (3.14) and the following blow-up criterion that if $T_{u_{0}}\leq T_{0}$, then
$\int_{0}^{T_{u_{0}}}\|v(t)\|^{4}_{\dot{H}^{\frac{\gamma}{4}}}dt=\infty.$
By interpolation (3.21), if $T_{u_{0}}\leq T_{0}$, then we have
$\forall T<T_{u_{0}},\quad\int_{0}^{T}\|v(t)\|^{4}_{\dot{H}^{\frac{\gamma}{4}}}dt\leq C(\|u_{0}\|_{H^{\sigma}},\|X\|_{C_{\bar{T}}W^{\alpha}},T_{0})<\infty.$
By contradiction, we have $T_{u_{0}}>T_{0}$, which shows that $v$ is the solution on $0,T_{0}]$. Consider equation (3.3) with initial data $v(T_{0})$. Setting $v^{(1)}(t)=v(T_{0}+t)$, $X^{(1)}(t)=X(T_{0}+t)$, we find that $v^{(1)}(t)$ satisfies the equation
$\partial_{t}v^{(1)}(t)-\Lambda^{\gamma}v^{(1)}(t)=\partial_{x}((v^{(1)}+X^{(1)})^{2}),\quad v^{(1)}(0)=v({T_{0}}).$ (3.22)
By a similar argument, we can also get a solution $v^{(1)}$ of equation on $0,T_{v(T_{0})}\wedge T_{0}[$ , since
$\|X^{(1)}\|_{C_{t}W^{\alpha}}=\|X\|_{C_{[T_{0},T_{0}+t]}W^{\alpha}}\leq\|X\|_{C_{\bar{T}}W^{\alpha}},\quad for\>t\in[0,\bar{T}-T_{0}].$
Then using the similar argument as above implies that $v^{(1)}(t)$ is the solution on $0,T_{0}]$. Repeating such process, we can get $v^{(n)},X^{(n)}$ satisfying
$\|X^{(n)}\|_{C_{t}W^{\alpha}}\leq\|X\|_{C_{\bar{T}}W^{\alpha}},\quad v^{(n)}(t)=v^{(n-1)}(T_{0}+t),\quad for\>t\in[0,\bar{T}-nT_{0}],$ (3.23)
where $v^{(n)}$ on $0,T_{0}]$ is also a solution of (3.3). Setting $v:=v^{(0)}$ on $0,T_{0}]$, $v^{new}:=v^{(i)}$ on $iT_{0},(i+1)T_{0}]$ for $i\leq n$, $v^{new}$ is a solution in $C([0,(n+1)T_{0}];L^{2}(\mathbb{T}))\cap L^{2}([0,(n+1)T_{0}];H^{\frac{\gamma}{2}}(\mathbb{T}))$. Choosing $n=[\frac{\bar{T}}{T_{0}}]$, we get $v\in C([0,\bar{T}];L^{2}(\mathbb{T}))\cap L^{2}([0,\bar{T}];H^{\frac{\gamma}{2}}(\mathbb{T}))$. Since $\bar{T}$ is arbitrary, we obtain $v$ as a weak global solution of (3.3). By the proof of Lemma 3.5, it’s unique. Since $X$ satisfies (1.5), we prove that $v$ is the weak global solution of (1.1). ∎
3.3 Application to DP equation
This section demonstrates the application of the aforementioned theory to the study of a class of shallow water wave equations. Taking the DP equation as an example, similar as the equation (3.3), we consider the difference equation of type of DP.
$\partial_{t}v-\Lambda^{\gamma}v=\frac{1+3(1-\partial_{x}^{2})^{-1}}{2}\partial_{x}(v^{2}+2v\tilde{X}+\tilde{X}^{2}),\quad v(0)=u_{0}.$ (3.24)
The only difference between (3.24) and (3.3) lies on $\frac{3}{2}(1-\partial_{x}^{2})^{-1}\partial_{x}(u^{2})$. Note that the operator $(1-\partial_{x}^{2})^{-1}$ is an ‘stable’ operator in $H^{s}$ since
$\|(1-\partial_{x}^{2})^{-1}f\|_{H^{s}}\leq\|f\|_{H^{s-2}},\quad for\>s\in\mathbb{R}.$
Hence, we can easily get the local result for rough DP equation. For the global result, we need the following energy estimate for DP equation.
Proposition 3.4.
For $\gamma\in(\frac{3}{2},2]$, $\alpha\in(1-\frac{\gamma}{2},\frac{\gamma}{2}-\frac{1}{2})$ and $0\leq T\leq\bar{T}<\infty$. Let $v$ be a smooth solution of (3.24) on $[0,T]$ for $u_{0}\in L^{2}$ and $\tilde{X}(t)\in C_{\bar{T}}W^{\alpha}$, then we have the following energy estimate
[…]
for some small $\nu<1$.
Proof.
Let $n=(1-\partial_{x}^{2})v$ and $w=(4-\partial_{x}^{2})^{-1}v$, similar as the proof of Lemma 4.1 in [CZG24], by (3.24), we have
$\int_{\mathbb{T}}v\partial_{x}nw+3\partial_{x}vnw\mathrm{d}x=0.$
Then we can derive it to the following equation
$\partial_{t}\int_{\mathbb{T}}nw\mathrm{d}x=$ $2\int_{\mathbb{T}}\langle\partial_{x}\rangle^{2}\partial_{t}vw\mathrm{d}x$
$=$ $2\int_{\mathbb{T}}\langle\partial_{x}\rangle^{2}\Lambda^{\gamma}vw\mathrm{d}x+2\int_{\mathbb{T}}\langle\partial_{x}\rangle^{2}\partial_{x}(2v\tilde{X}+\tilde{X}^{2})w\mathrm{d}x+3\int_{\mathbb{T}}\partial_{x}(2v\tilde{X}+\tilde{X}^{2})w\mathrm{d}x.$
By Parseval formula, Hölder inequality and interpolation inequality, we have
$\int_{\mathbb{T}}\langle\partial_{x}\rangle^{2}\partial_{x}(v\tilde{X})w\mathrm{d}x=$ $\int_{\mathbb{T}}\frac{1+|\xi|^{2}}{4+|\xi|^{2}}\mathcal{F}(\partial_{x}(v\tilde{X}))\mathcal{F}(v)\mathrm{d}x$ (3.26)
$\leq$ $\|v\|_{H^{\frac{\gamma}{2}}}\|v\tilde{X}\|_{H^{1-\frac{\gamma}{2}}}$
$\leq$ $\|v\|_{H^{\frac{\gamma}{2}}}^{1+\frac{1}{\gamma}}\|v\|_{L^{2}}^{1-\frac{1}{\gamma}}\|\tilde{X}\|_{\mathcal{C}^{\alpha}}$
$\leq$ $\frac{\nu}{6}\|v\|_{H^{\frac{\gamma}{2}}}^{2}+C_{\nu}\|v\|_{L^{2}}^{2}\|\tilde{X}\|_{W^{\alpha}}^{\frac{2\gamma}{\gamma-1}},$
where we use the fact $\alpha>{1-\frac{\gamma}{2}}$. Similarly, by Lemma 2.1, we have
$\int_{\mathbb{T}}\langle\partial_{x}\rangle^{2}\partial_{x}\tilde{X}^{2}w\mathrm{d}x\leq$ $\frac{\nu}{6}\|v\|_{H^{\frac{\gamma}{2}}}^{2}+C_{\nu}\|\tilde{X}\|_{W^{\alpha}}^{4},$ (3.27)
$\int_{\mathbb{T}}\partial_{x}(2v\tilde{X}+\tilde{X}^{2})w\mathrm{d}x\leq$ $\frac{\nu}{3}\|v\|_{H^{\frac{\gamma}{2}}}^{2}+C_{\nu}\|v\|_{L^{2}}^{2}\|\tilde{X}\|_{W^{\alpha}}^{\frac{2\gamma}{\gamma-1}}+C_{\nu}\|\tilde{X}\|_{W^{\alpha}}^{4}.$ (3.28)
Combining above estimates, we have
$\int_{\mathbb{T}}n(t)w(t)\mathrm{d}x$ $+(2-2\nu)\int_{0}^{t}\|v(t^{\prime})\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}$ (3.29)
$\leq$ $\int_{\mathbb{T}}n(0)w(0)\mathrm{d}x+C_{\nu}\int_{0}^{t}\|v(t^{\prime})\|_{L^{2}}^{2}\|\tilde{X}(t^{\prime})\|_{W^{\alpha}}^{\frac{2\gamma}{\gamma-1}}\mathrm{d}t^{\prime}+C_{\nu}\int_{0}^{t}\|\tilde{X}(t^{\prime})\|_{W^{\alpha}}^{4}\mathrm{d}t^{\prime}.$
Noting that
$\|v(t)\|_{L^{2}}^{2}=\|\widehat{v}(t)\|_{L^{2}}^{2}\leq 4\int_{\mathbb{T}}\frac{1+\xi^{2}}{4+\xi^{2}}|\widehat{v}(t)|^{2}\mathrm{d}x=4\int_{\mathbb{T}}n(t)w(t)\mathrm{d}x\leq 4\|v(t)\|_{L^{2}}^{2}.$ (3.30)
Particularly, $\int_{\mathbb{T}}n(0)w(0)\mathrm{d}x\leq 4\|u_{0}\|_{L^{2}}^{2}$, by (3.29) we have
$\|v(t)\|_{L^{2}}^{2}+(8-8\nu)\int_{0}^{t}\|v(t^{\prime})\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}\leq 16\|u_{0}\|_{L^{2}}^{2}+C_{\nu}\int_{0}^{t}\|v(t^{\prime})\|_{L^{2}}^{2}\|\tilde{X}(t^{\prime})\|_{W^{\alpha}}^{\frac{2\gamma}{\gamma-1}}\mathrm{d}t^{\prime}+C_{\nu}\int_{0}^{t}\|\tilde{X}(t^{\prime})\|_{W^{\alpha}}^{4}\mathrm{d}t^{\prime}$ (3.31)
for $\nu<1$ and $t\in[0,T]$. By Gronwall’s inequality, we finish our proof. ∎
Similar to the proof of Lemma 3.4, we can improve the estimate in higher regularity as follows. The proof is omitted.
Lemma 3.6.
Let $\gamma\in(\frac{3}{2},2]$, $\alpha\in(s+1-\frac{\gamma}{2},\frac{\gamma}{2}-\frac{1}{2})$ , $0<s<\gamma-\frac{3}{2}$, and $0\leq T\leq\bar{T}<\infty$. Let $v$ be the smooth solution of (3.24) on $[0,T]$ for $u_{0}\in H^{s}$ with $\tilde{X}\in C_{\bar{T}}W^{\alpha}$. For $t\in[0,T]$, we have the following energy estimate
$\|\Lambda^{s}v(t)\|_{L^{2}}+2(1-\nu)\int_{0}^{t}\|\Lambda^{s}v(t^{\prime})\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}\leq$ $\|u_{0}\|^{2}_{L^{2}}e^{C_{\nu}\int_{0}^{t}\|v(t^{\prime})\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}}+\|\tilde{X}\|_{C_{\bar{T}}W^{\alpha}}^{4}\int_{0}^{t}e^{C_{\nu}\int_{t^{\prime}}^{t}\|v(t^{\prime\prime})\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime\prime}}\mathrm{d}t^{\prime}$ (3.32)
$+\|\tilde{X}\|_{C_{\bar{T}}W^{\alpha}}^{2}\int_{0}^{t}e^{C_{\nu}\int_{t^{\prime}}^{t}\|v(t^{\prime\prime})\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime\prime}}\|v(t^{\prime})\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}$
for some small $\nu<1$. Moreover, $v\in C_{T}H^{s}\cap L_{T}^{2}H^{s+\frac{\gamma}{2}}$.
Using a similar argument as in the proof of Theorem 1.2, we finish the proof of Remark 1.3.
4 Result in the range of $\gamma\in(\frac{5}{4},\frac{4}{3}]$
4.1 Regularity analysis
This section is devoted to isolating a remainder term of Sobolev regularity $\frac{1}{2}+$ under the weaker dissipation condition $\gamma\in(\frac{5}{4},\frac{4}{3}]$. We will achieve this by means of the paracontrolled method, introduced in [GIP15, GST19], a technique especially well‑suited for analyzing the regularity structure of solutions to singular SPDEs.
Our strategy as follows. Recall the difference equation
$\partial_{t}v-\Lambda^{\gamma}v=\mathcal{L}(v)=$ $\mathcal{L}(u-X)$
$=$ $\partial_{x}(u^{2})+\xi-\xi$
$=$ $\partial_{x}(v^{2})+\partial_{x}(X^{2})+2\partial_{x}(vX).$
Assume that $v$ gains the critical regularity $H^{\frac{1}{2}+}$. By the critical analysis, $\partial_{x}(v^{2})$, $\partial_{x}(X^{2})$ and $\partial_{x}(vX)$ must gain $H^{\frac{1}{2}-\gamma+}$-order regularity at least. But it’s impossible in the weaker dissipation conditions $\gamma\in(\frac{5}{4},\frac{4}{3}]$ since $\partial_{x}(X^{2})$ and $\partial_{x}(vX)$ can gain $\alpha-1$-order regularity at most. To overcome this difficult, our strategy is to decompose such terms in “ higher-regularity-term" and “ lower- regularity-term" by para-product decomposition (1.9). Correspondingly, we write
$v=w+u^{\sharp},$ (4.1)
where $\mathcal{L}u^{\sharp}$ corresponds to the “higher-regularity-term", while $\mathcal{L}w$ is designed to absorb the “ lower- regularity-term".
Next, our work is carried out in the framework of the space $W^{s}=\mathcal{C}^{s}\cap H^{s}$. We introduce a method called "para-controlled solution", which was developed in [GIP15, GP17] to denote by $u^{\sharp}$ the part of $u$ that gains $W^{\frac{1}{2}+}$ regularity. To get a higher order term in $W^{\frac{1}{2}+}$. Let distribution $u^{\sharp}$ satisfy equation
$\mathcal{L}u^{\sharp}=\partial_{x}(u^{\sharp})^{2}+R(X,u^{\sharp},u),\quad u^{\sharp}(0)=u_{0}\in W^{\frac{1}{2}+}.$ (4.2)
where $R(X,u^{\sharp},u)$ is to be determined. Assume that $R(X,u^{\sharp},u)\in W^{\frac{1}{2}-\gamma+}$. Then by regularity theory, $u^{\sharp}$ belongs to $W^{\frac{1}{2}+}$. Fix $u=X+w+u^{\sharp}$, then we have
$\mathcal{L}(u^{\sharp})=$ $\mathcal{L}(u-X-w)$
$=$ $\partial_{x}(u^{2})+\xi-\xi-\mathcal{L}w$
$=$ $\partial_{x}(w^{2})+\partial_{x}(X^{2})+2\partial_{x}(wX)+2\partial_{x}(wu^{\sharp})+2\partial_{x}(Xu^{\sharp})+(\partial_{x}((u^{\sharp})^{2})-\mathcal{L}w).$
Since $u^{\sharp}$ possesses sufficient regularity, the product $u^{\sharp}\cdot w$ preserve regularity of $w$. The problematic terms are $\partial_{x}(w\cdot X)$ and $\partial_{x}(X^{2})$, as they can gain at most $\alpha-1$-order regularity which below the required $\frac{1}{2}-\gamma+$ in weak dissipation conditions $\gamma\in(\frac{5}{4},\frac{4}{3}]$. Using para-product decomposition (1.9) and the fact that $X\in C_{T}\mathcal{C}^{\alpha}$, we observe that $\partial_{x}(X\circ X)$, $\partial_{x}(X\prec w)$ and $\partial_{x}(X\circ w)$ all belong to $H^{\frac{1}{2}-\gamma}$. The main difficulty therefore lies in the terms $\partial_{x}(v\prec X)$ and $\partial_{x}(X\prec X)$.
To handle this, we assume that $v$ possesses a suitable structure that can absorb the problematic terms. Motivated by the paracontrolled ansatz, we write
$w=u^{\prime}\prec\!\!\!\prec\mathcal{Q},\quad u^{\mathcal{Q}}:=w+u^{\sharp}.$ (4.3)
with the assumption $u^{\prime}\in\mathcal{W}^{\frac{1}{8}+}(T)$, $\mathcal{Q}\in\mathcal{W}^{\frac{3}{8}+}$ and $u^{\sharp}\in\mathcal{W}^{\frac{1}{2}+}(T)$ where $\mathcal{W}^{s}(T)$ is defined in Section 2. By Lemma 2.5, this implies $w\in\mathcal{W}^{\frac{3}{8}+}(T)$. For $\gamma>\frac{5}{4}$, the regularity analysis proceeds as follows
$2[\partial_{x}(X\prec X)-X\prec\partial_{x}X]+\partial_{x}(w^{2})+2\partial_{x}(wu^{\sharp})+\partial_{x}(X\circ X)\in C_{T}W^{-\frac{3}{4}+},\quad 2\partial_{x}(X\prec u^{\mathcal{Q}})\in C_{T}W^{-\frac{5}{8}+},$
$2\partial_{x}(X\circ u^{\mathcal{Q}})\in C_{T}W^{-\frac{1}{2}+},\quad 2[u^{\prime}\prec\!\!\!\prec\partial_{x}X-u^{\prime}\prec\partial_{x}X]\in C_{T}W^{-\frac{3}{4}+},\quad u^{\prime}\prec\!\!\!\prec\mathcal{L}\mathcal{Q}-\mathcal{L}(u^{\prime}\prec\!\!\!\prec\mathcal{Q})\in C_{T}W^{\frac{1}{2}-\gamma+}.$
Therefore, we can determine $R(X,u^{\sharp},u)$ as
$R(X,u^{\sharp},u):=$ $\partial_{x}(w^{2})+2\partial_{x}(wu^{\sharp})+\partial_{x}(X\circ X)+2[\partial_{x}(X\prec X)-X\prec\partial_{x}X]$
$+2\partial_{x}(X\prec u^{\mathcal{Q}}+X\circ u^{\mathcal{Q}}+[\partial_{x}(u^{\mathcal{Q}}\prec X)-u^{\mathcal{Q}}\prec\partial_{x}X]$
$+u^{\prime}\prec\!\!\!\prec\mathcal{L}\mathcal{Q}-\mathcal{L}(u^{\prime}\prec\!\!\!\prec\mathcal{Q})+[u^{\prime}\prec\partial_{x}X-u^{\prime}\prec\!\!\!\prec\partial_{x}X].$
Here we see that the purpose of using modified paraproduct operator $f\prec\!\!\!\prec g$ is to ensure that the commutator $\mathcal{L}(u^{\prime}\prec\!\!\!\prec\mathcal{Q})-u^{\prime}\prec\!\!\!\prec\mathcal{L}\mathcal{Q}$ aligns with our regularity framework. Let $u^{\prime}=X+u^{\mathcal{Q}}$, we counter the rough term.
$u^{\prime}\prec\mathcal{L}Q=$ $2X\prec\partial_{x}X+2u^{\mathcal{Q}}\prec\partial_{x}X,$
where the last term has higher regularity. By equation (4.2), we construct a paracontrolled solution if we can verify $u^{\sharp}$ has enough regularity. In fact, by regularity theory, for $\gamma>\frac{5}{4}$, (4.2) derive the regularity of $u^{\sharp}$ such as
$u^{\sharp}\in\mathcal{W}^{\frac{1}{2}+\delta}(T),$
which we finish our ansatz setting.
4.2 Construction of para-controlled solution
We frame this problem as solving a system of equations of $(u^{\prime},u^{\sharp})$. Given an equation of $(u^{\prime},u^{\sharp})$ driven by $(\tilde{X},u_{0})$.
$\left\{\begin{aligned} &u^{\prime}=2\tilde{X}+2u^{\mathcal{Q}}\\ &\mathcal{L}u^{\sharp}=[\partial_{x}(\tilde{X}^{2})-2\tilde{X}\prec\partial_{x}\tilde{X}]+\partial_{x}(u^{\mathcal{Q}})^{2}+2[\partial_{x}(u^{\mathcal{Q}}\tilde{X})-u^{\mathcal{Q}}\prec\partial_{x}\tilde{X}]+[u^{\prime}\prec\partial_{x}\tilde{X}-\mathcal{L}(u^{\prime}\prec\!\!\!\prec\mathcal{Q})],\end{aligned}\right$. (4.4)
with $u^{\sharp}(0)=u_{0}$. We have the following proposition
Proposition 4.1.
Let $\alpha\in(\frac{1}{8},\frac{1}{6})$, $\gamma\in(\frac{5}{4},\frac{4}{3}]$ and $u_{0}\in W^{\frac{1}{2}+\delta}$ for $0<\delta\leq\min(\gamma-\frac{5}{4},\alpha-\frac{1}{8})$. If $\tilde{X}\in\mathcal{W}^{\alpha}(\bar{T})$ for some $\bar{T}>0$, then there exists a unique solution $(u^{\prime},u^{\sharp})$ of equation (4.4) with positive $T^{*}$ satisfied $(u^{\prime},u^{\sharp})\in\mathcal{W}^{\frac{1}{8}+\delta}(T^{*})\times\mathcal{W}^{\frac{1}{2}+\delta}(T^{*})$.
Proof.
Firstly, let $u^{\sharp}(0)=u_{0}$ and $f(u^{\sharp},u^{\prime})=u^{\mathcal{Q}}=u^{\prime}\prec\!\!\!\prec Q+u^{\sharp}$, we can rewrite (4.4) in mild form
$u^{\prime}=2\tilde{X}+2u^{\mathcal{Q}},\quad\quad u^{\sharp}=a+g(u^{\prime})+\mathcal{B}(f(u^{\sharp},u^{\prime}),f(u^{\sharp},u^{\prime}))+\mathcal{G}(f(u^{\sharp},u^{\prime})).$
and
$\mathcal{L}a=P(t)u_{0}+[\partial_{x}(\tilde{X}^{2})-2\tilde{X}\prec\partial_{x}\tilde{X}$], $\quad a(0)=0,$
$\mathcal{L}(\mathcal{G}(f(u^{\sharp},u^{\prime})))=2[\partial_{x}(u^{\mathcal{Q}}\tilde{X})-u^{\mathcal{Q}}\prec\partial_{x}\tilde{X}$], $\quad\mathcal{G}(f(u^{\sharp},u^{\prime}))(0)=0,$
$\mathcal{L}{\mathcal{B}}(f(u^{\sharp},u^{\prime}),f(u^{\sharp},u^{\prime}))=\partial_{x}(u^{\mathcal{Q}})^{2},$ $\quad{\mathcal{B}}(f(u^{\sharp},u^{\prime})(0)=0,$
$\mathcal{L}g(u^{\prime})=[u^{\prime}\prec\!\!\!\prec\partial_{x}\tilde{X}-\mathcal{L}(u^{\prime}\prec\!\!\!\prec\mathcal{Q})]+[u^{\prime}\prec\partial_{x}\tilde{X}-u^{\prime}\prec\!\!\!\prec\partial_{x}\tilde{X}$], $\quad g(u^{\prime})(0)=0.$
Define $\mathcal{E}_{T}$ as follows
$\mathcal{E}_{T}:=\{(u,v)\in\mathcal{W}^{\frac{1}{2}+\delta}(T)\times\mathcal{W}^{\frac{1}{8}+\delta}(T):2\|u\|_{\mathcal{W}^{\frac{1}{2}+\delta}(T)}+\|v\|_{\mathcal{W}^{\frac{1}{8}+\delta}(T)}<\infty\}.$
Our goal is to find a $T^{*}\leq 1\wedge\bar{T}$ such that the conditions of Lemma 2.9 hold in $\mathcal{E}_{T^{*}}$. Assume that $T\leq 1\wedge\bar{T}$, by Lemma 2.5, for $\epsilon=\frac{1}{8}+\delta$, $\rho=\gamma-\delta$ and $\delta\leq\min(\gamma-\frac{5}{4},\alpha-\frac{1}{8})$, we have
$\|f(u^{\sharp},u^{\prime})\|_{\mathcal{W}^{\frac{1}{8}+\delta}(T)}\leq$ $\|u^{\sharp}\|_{\mathcal{W}^{\frac{1}{2}+\delta}(T)}+CT^{\frac{\delta}{\gamma}}\|u^{\prime}\|_{\mathcal{W}^{\frac{1}{8}+\delta}(T)}(\|\mathcal{Q}\|_{C_{T}\mathcal{C}^{\frac{1}{8}+\delta}}+\|\partial_{x}\tilde{X}\|_{C_{T}\mathcal{C}^{\frac{1}{8}-\gamma+2\delta}}),$
$\leq$ $\|u^{\sharp}\|_{\mathcal{W}^{\frac{1}{2}+\delta}(T)}+CT^{\frac{\delta}{\gamma}}\|u^{\prime}\|_{\mathcal{W}^{\frac{1}{8}+\delta}(T)}\|\tilde{X}\|_{\mathcal{W}^{\alpha}(\bar{T})}.$
and
$\|f(u^{\sharp},u^{\prime})\|_{\mathcal{W}^{\frac{3}{2}-\gamma+\delta}(T)}\leq$ $\|u^{\sharp}\|_{\mathcal{W}^{\frac{1}{2}+\delta}(T)}+CT^{\frac{\delta}{\gamma}}\|u^{\prime}\|_{\mathcal{W}^{\frac{1}{8}+\delta}(T)}(\|\mathcal{Q}\|_{C_{T}W^{\frac{3}{2}-\gamma+\delta}}+\|\partial_{x}\tilde{X}\|_{\frac{3}{2}-2\gamma+2\delta})$
$\leq$ $\|u^{\sharp}\|_{\mathcal{W}^{\frac{1}{2}+\delta}(T)}+CT^{\frac{\delta}{\gamma}}\|u^{\prime}\|_{\mathcal{W}^{\frac{1}{8}+\delta}(T)}\|\tilde{X}\|_{\mathcal{W}^{\alpha}(\bar{T})}.$
Fixing $T$ small enough, we can make sure that $CT^{1-\frac{\delta}{\gamma}}\|\tilde{X}\|_{\mathcal{W}^{\frac{1}{8}+\delta}(\bar{T})}<\frac{1}{30}$.
For fixed $r=16(1\vee C)\max(\|u_{0}\|_{W^{\frac{1}{2}+\delta}},\|\tilde{X}\|_{\mathcal{W}^{\alpha}(\bar{T})})$, by Lemma A.2 and Lemma B.1, we can choose $T\leq 1\wedge\bar{T}$ small enough such that
$\|(\tilde{X},a)\|_{\mathcal{E}_{T}}\leq$ $\|\tilde{X}\|_{\mathcal{W}^{\alpha}(T)}+2C\|u_{0}\|_{W^{\frac{1}{2}+\delta}}+2CT^{\frac{\delta}{\gamma}}\|\mathcal{L}a\|_{C_{T}W^{\frac{1}{2}-\gamma+2\delta}}$
$\leq$ $\frac{r}{16}+\frac{r}{8}+\frac{r}{16}\leq\frac{r}{4}.$
By Lemma B.3 and Lemma B.1, we have a similar argument as above. We can choose $T\leq 1\wedge\bar{T}$ small enough such that
$\|g(u^{\prime})\|_{\mathcal{W}^{\frac{1}{2}+\delta}(T)}\leq CT^{\frac{\delta}{\gamma}}\|\mathcal{L}g(u^{\prime})\|_{C_{T}W^{\frac{1}{2}-\gamma+2\delta}}\leq C$ $T^{\frac{\delta}{\gamma}}\|u^{\prime}\|_{\mathcal{W}^{\frac{1}{8}+\delta}(T)}(\|\mathcal{Q}\|_{C_{{T}}W^{\frac{3}{8}+\delta}}+\|\partial_{x}\tilde{X}\|_{C_{T}W^{\frac{3}{8}-\gamma+\delta}})$
$\leq$ $CT^{\frac{\delta}{\gamma}}\|u^{\prime}\|_{\mathcal{W}^{\frac{1}{8}+\delta}(T)}\|\tilde{X}\|_{C_{\bar{T}}W^{\alpha}}\leq\frac{1}{16}\|u^{\prime}\|_{\mathcal{W}^{\frac{1}{8}+\delta}(T)}.$
and
$\|\mathcal{G}(f(u^{\sharp},u^{\prime}))\|_{\mathcal{W}^{\frac{1}{2}+\delta}(T)}\leq$ $CT^{\frac{\delta}{\gamma}}\|\mathcal{L}\mathcal{G}(f(u^{\sharp},u^{\prime}))\|_{C_{T}W^{\frac{1}{2}-\gamma+2\delta}}$
$\leq$ $CT^{\frac{\delta}{\gamma}}\|f(u^{\sharp},u^{\prime})\|_{\mathcal{W}^{\frac{1}{8}+\delta}}\|\tilde{X}\|_{\mathcal{W}^{\alpha}(\bar{T})}$
$\leq$ $\frac{1}{16}\|f(u^{\sharp},u^{\prime})\|_{\mathcal{W}^{\frac{1}{8}+\delta}}.$
where we use the fact $\gamma>\frac{3}{2}$. Finally, let $X=\mathcal{W}^{\frac{3}{2}-\gamma+\delta}(T)$, we have $\mathcal{W}^{\frac{1}{2}+\delta}(T)\subset\mathcal{W}^{\frac{3}{2}-\gamma+\delta}(T)\subset\mathcal{W}^{\frac{1}{8}+\delta}(T)$, Choosing $T$ small enough such that
$\|\mathcal{B}(f(u^{\sharp},u^{\prime}),f(u^{\sharp},u^{\prime}))\|_{\mathcal{W}^{\frac{1}{2}+\delta}(T)}\leq$ $CT^{\frac{\delta}{\gamma}}\|\mathcal{L}\mathcal{B}(f(u^{\sharp},u^{\prime}))\|_{C_{T}W^{\frac{1}{2}-\gamma+2\delta}}$
$\leq$ $CT^{\frac{\delta}{\gamma}}\|\partial_{x}(f(u^{\sharp},u^{\prime})^{2})\|_{C_{T}W^{\frac{1}{2}-\gamma+\delta}}$
$\leq$ $CT^{\frac{\delta}{\gamma}}\|f(u^{\sharp},u^{\prime})\|^{2}_{C_{T}W^{\frac{3}{2}-\gamma+\delta}}$
$\leq$ $\frac{1}{16}\|f(u^{\sharp},u^{\prime})\|^{2}_{\mathcal{W}^{\frac{3}{2}-\gamma+\delta}(T)}.$
Setting $T^{*}$ to satisfy all the conditions above, by Lemma 2.9, there exists $(u^{\sharp},u^{\prime})\in\mathcal{E}_{T^{*}}$ for which we prove our proposition. ∎
Proof of Theorem 1.3.
Let $\tilde{X}(t)=X(t)$, where $X(t)$ satisfies equation (1.5). By Remark 2.1, we have $\tilde{X}\in\mathcal{W}^{\alpha}({\bar{T}})$ for any fixed $\bar{T}>0$ and $\alpha\in(\frac{1}{8},\frac{\gamma}{2}-\frac{1}{2})$. Then by Proposition 4.1, there exists $(u^{\prime},u^{\sharp})\in C_{T^{*}}W^{\frac{1}{8}+\delta}\times C_{T^{*}}W^{\frac{1}{2}+\delta}$ satisfying (4.4) for $\delta\leq\min(\gamma-\frac{5}{4},\alpha-\frac{1}{8})<\frac{\gamma}{2}-\frac{5}{8}$ and $0<T^{*}\leq\bar{T}$. It’s not difficult to check that $u=X+u^{\prime}\prec\!\!\!\prec\mathcal{Q}+u^{\sharp}$ is the solution of (1.1) by the analysis in Section 4.1. Since $u\in C_{T^{*}}W^{\alpha}$, we finish the proof of Theorem 1.3. ∎
5 Discussion on the range of $\gamma$
In the final section, we propose a conjecture regarding the admissible range of the dissipation parameter $\gamma$ for the rough or singular Burgers equation driven by $|D|^{\beta}\xi$ by the method in [DNY22].
$u_{t}-\Lambda^{\gamma}u=\partial_{x}(u^{2})+|D|^{\beta}\xi,$ (5.1)
where $\beta\geq 0$. For $\zeta=|D|^{\beta}\xi$, we fix a dyadic number $N\neq 0$, and set $\zeta^{N}:=\mathcal{F}^{-1}(1_{|k|\sim N}\mathcal{F}\xi(k))$, which means that the frequency of $\xi$ is focus on $\frac{1}{2}N\leq|k|\leq N$. Then the Fourier mode of the linear evolution $X:=\int_{0}^{t}e^{(t-s)\Lambda^{\gamma}}\zeta^{N}\mathrm{d}s$ can be written as
$X=N^{\beta-\frac{\gamma}{2}}\sum_{|k|\sim N}G_{k}(t)e^{ikx},\quad G_{k}(t)=N^{\frac{\gamma}{2}}\int_{0}^{t}e^{-(t-s)|k|^{\gamma}}\xi_{k}(s)\mathrm{d}s,$
It’s not difficult to calculate that $G_{k}(t)$ from a collection of independent Gaussian variable with $\mathbb{E}[|G_{k}(t)|^{2}]\sim 1$, then we can get $X$ belong to $C_{T}H^{\frac{\gamma}{2}-\beta-\frac{1}{2}}$. Now we calculate the "first nonlinear iteration". For $u^{(1)}(t):=\int_{0}^{t}e^{(t-s)\Lambda^{\gamma}}\partial_{x}(X^{2})\mathrm{d}s$, and let $u_{k}^{(1)}(t)$ be its Fourier mode
$u_{k}^{(1)}(t)=N^{2\beta-\gamma+1}\sum_{\begin{subarray}{c}l,m\in\mathbb{Z},|l|\sim|m|\sim|k|\sim|N|\\ l+m=k\end{subarray}}\int_{0}^{t}e^{-(t-s)|k|^{\gamma}}G_{l}(s)G_{m}(s)\mathrm{d}s.$
Since the heat integral always provide $N^{-\gamma}$and the sum only has size $N^{1}$, then the inner sum integral has size $N^{\frac{1}{2}-\gamma}$ with high probability by square root cancellation, then we get the ansatz that
$\beta\leq\frac{3}{2}\gamma-\frac{3}{2}.$ (5.2)
We call a pair $(\beta,\gamma)$ “probabilistic admissible pair" for the Rough or singular Burgers equation on $H^{s}$ space(In the $\mathcal{C}^{s}$ space, it will be different), if condition (5.2) holds. It is immediate to verify that the pair $(0,\gamma)$ for $\gamma>\frac{4}{3}$ and $(\frac{1}{2},\frac{3}{2})$ are probabilistic admissible pairs; the former regime is treated in the present paper, while the latter will be our future directions.
5.1 Future directions
In future work, we intend to extend our analysis to more singular regimes of the stochastic Burgers equation. A primary example is the parameter choice $\gamma\geq\frac{3}{2}$ and $\beta=\frac{1}{2}$ in (5.1). The solution $u$ inherits the spatial regularity of the stochastic convolution $X(t)=\int_{0}^{t}P(t-s)\xi(s)\mathrm{d}s$. According to the Lemma 2.7, we have a priori $u\in W^{-\frac{1}{2}-\delta}$ for any $\delta>0$. At such low regularity, the nonlinear term $\partial_{x}(u^{2})$ may become ill-defined, as the assumptions required for Lemma 2.1 no longer hold. A standard way to circumvent this difficulty is to work with generalized solutions. More precisely, we consider a regularized solution $u^{\epsilon}$ satisfying
$\partial_{t}u^{\epsilon}-\Lambda^{\gamma}u^{\epsilon}=D(u^{\epsilon})^{2}+|D|^{\beta}\xi_{\epsilon},\quad u^{\epsilon}(0)=u^{\epsilon}_{0}+Y_{\epsilon}(0),$ (5.3)
and prove that $u^{\epsilon}$ converges in probability to a limiting process $u$. This approach was originally introduced by Martin Hairer in his analysis of the KPZ equation [Hai13]. In his subsequent foundational work [Hai14], he developed the comprehensive framework of regularity structures, which provides a systematic way to define solutions for a broad class of singular stochastic partial differential equations—contributions that were recognized with the Fields Medal. Within that framework, the solution is constructed in a space of modelled distributions $\mathcal{D}^{\gamma,\eta}_{P}$ over $[0,T]$ and satisfies the fixed-point equation
$u=(\mathcal{K}_{\bar{\Gamma}}+R_{\gamma}\mathcal{R})\textbf{R}^{+}F(u)+Gu_{0}.$ (5.4)
Appendix A Estimate for heat flow
In this chapter, we give some estimates for heat flow. Defining operator $P(t)$ as
$\mathcal{F}(P(t)f)=e^{-t|\xi|^{\gamma}}\mathcal{F}f,$ (A.1)
for $f\in\mathcal{S}^{\prime}(\mathbb{R}^{d})$, it’s not difficult to see that $P(t)f$ is the solution of
$\partial_{t}P(t)f-\Lambda^{\gamma}P(t)f=0,\quad u(0)=f.$ (A.2)
For the operator, we firstly claim that it’s the linear bounded operator in $L^{p}$ for $p\geq 1$. In fact, setting $\varphi(z)=e^{-|z|^{\gamma}}$, then $P(t)$ can be written as $\varphi(t^{\frac{1}{\gamma}}D)$. Noting that $\varphi\in L^{1}$, it suffice to show $\mathcal{F}\varphi\in L^{1}$. Then for any $t\geq 0$, we have
$\|P(t)f\|_{L^{p}}=\|\mathcal{F}^{-1}\varphi(t^{\frac{1}{\gamma}}\cdot)\ast f\|_{L^{p}}\lesssim\|f\|_{L^{p}}.$ (A.3)
We can easily get $\mathcal{F}\varphi$ belongs to $L^{1}$ since its the density of a symmetric $\gamma$-stable random variable for $\gamma\in(0,2]$.
Lemma A.1.
For $\gamma>0$, $\lambda\in\mathbb{Z}/\{0\}$, we have the estimate
$\|P(t)f\|_{L^{p}}\leq Ce^{-ct\lambda^{\gamma}}\|f\|_{L^{p}},$
if the support of $\hat{f}$ belongs to some $\lambda\mathcal{C}$, where $\mathcal{C}$ is a fixed annual.
Proof.
The proof is similar to Lemma 2.4 in [BCD11], which only need to notice that $\partial_{x}^{\beta}e^{-t|\xi|^{\gamma}}\leq C(1+t)^{|\beta|}e^{-ct}$ for $\xi\in\lambda\mathcal{C}$ will not cause difficult when $\gamma>0$, since $|\xi|$ has a uniformly lower bound for $\lambda\in\mathbb{Z}/\{0\}$. ∎
For Besov space $B_{p,r}^{s}$, we have the following result.
Lemma A.2.
Let $1\leq p,r\leq\infty$, $\alpha\in\mathbb{R}$ and $\gamma>0$, then for $t\geq 0$, we have the following estimate
$\|P(t)f\|_{B_{p,r}^{\alpha}}\lesssim\|f\|_{B_{p,r}^{\alpha}}.$ (A.4)
Let $\delta\geq 0$ and $0<t\leq 1$, we have
$\|P(t)f\|_{B_{p,r}^{\alpha+\delta}}\leq t^{-\frac{\delta}{\gamma}}\|f\|_{B_{p,r}^{\alpha}}.$ (A.5)
Moreover, fix $f(0),\mathcal{L}f:=(\partial_{t}-\Lambda^{\gamma})f\in B_{p,r}^{\alpha}$ , then for $\delta\in[0,\gamma)$ and $t\leq 1$, we have estimate
$\|f\|_{B_{p,r}^{\alpha+\delta}}\lesssim\|f(0)\|_{B_{p,r}^{\alpha+\delta}}+t^{1-\frac{\delta}{\gamma}}\|\mathcal{L}f\|_{C_{t}B_{p,r}^{\alpha}}.$ (A.6)
Proof.
For fixed $j\geq 0$, by Lemma A.1, for $t>0$, we have
$\|P(t)\Delta_{j}f(t)\|_{L^{p}}\lesssim$ $Ce^{-ct2^{j\gamma}}\|\Delta_{j}f\|_{L^{p}}$
$\lesssim$ $t^{-\frac{\delta}{\gamma}}2^{-j\delta}\|\Delta_{j}f\|_{L^{p}}$
For j=1j=-1, by (A.3) and $t\leq 1$, we have
$\|P(t)\Delta_{j}f(t)\|_{L^{p}}\lesssim$ $\|\Delta_{j}f(t)\|_{L^{p}}$
$\lesssim$ $(2^{-\delta}t^{\frac{\delta}{\gamma}})t^{-\frac{\delta}{\gamma}}2^{\delta}\|\Delta_{j}f(t)\|_{L^{p}}$
$\lesssim$ $t^{-\frac{\delta}{\gamma}}2^{-j\delta}\|\Delta_{j}f\|_{L^{p}}.$
Multiplying $2^{j(\alpha+\delta)}$ both sides and taking the $l^{r}$-norm with respect to $1\leq j$, we prove the second result.
For (A.6), by (A.4) and (A.5)
$\|f\|_{B_{p,r}^{\alpha+\delta}}\leq$ $\|P(t)f(0)\|_{B_{p,r}^{\alpha+\delta}}+\int_{0}^{t}\|P(t-s)\mathcal{L}f(s)\|_{B_{p,r}^{\alpha+\delta}}\mathrm{d}s$
$\lesssim$ $\|f(0)\|_{B_{p,r}^{\alpha+\delta}}+\int_{0}^{t}\|(t-s)^{-\frac{\delta}{\gamma}}\mathcal{L}f(s)\|_{B_{p,r}^{\alpha}}\mathrm{d}s$
$\lesssim$ $\|f(0)\|_{B_{p,r}^{\alpha+\delta}}+t^{1-\frac{\delta}{\gamma}}\|\mathcal{L}f\|_{C_{t}B_{p,r}^{\alpha}}$
where we use $\delta<\gamma$. ∎
A similar result carries over to the space $C_{T}^{\alpha}L^{2}$ and $C_{T}^{\alpha}L^{\infty}$, as stated in the following lemma.
Lemma A.3.
Let $\gamma>0$ and $\alpha\in\mathbb{R}_{+}$, for any $T\geq 0$ we have the following estimate
$\|P(\cdot)f\|_{\mathcal{W}^{\alpha}_{\infty}(T)}\lesssim\|f\|_{\mathcal{C}^{\alpha}},\quad\|P(\cdot)f\|_{\mathcal{W}^{\alpha}_{2}(T)}\lesssim\|f\|_{H^{\alpha}}.$
Moreover, let $\alpha\in(0,\gamma)$, $\delta\in[0,\gamma)$ satisfy $0<\alpha+\delta<\gamma$ and $0\leq T\leq 1$, it has the estimate
$\|f\|_{\mathcal{W}_{\infty}^{\alpha+\delta}(T)}\lesssim\|f(0)\|_{W_{\infty}^{\alpha+\delta}}+T^{1-\frac{\delta}{\gamma}}\|\mathcal{L}f\|_{C_{T}W_{\infty}^{\alpha}},\quad\|f\|_{\mathcal{W}_{2}^{\alpha+\delta}(T)}\lesssim\|f(0)\|_{W_{2}^{\alpha+\delta}}+T^{1-\frac{\delta}{\gamma}}\|\mathcal{L}f\|_{C_{T}W_{2}^{\alpha}}.$
Proof.
We only proof the case of $\mathcal{W}_{2}^{\alpha}(T)$. The case of $\mathcal{W}_{\infty}^{\alpha+\delta}(T)$ can be refer to Lemma 2.9 in [GP17]. The estimate for $C_{T}H^{\alpha}$ and $C_{T}H^{\alpha+\delta}$ is a direct consequence of Lemma A.2. For $C_{T}^{\frac{\alpha}{\gamma}}L^{2}$, by Parsavel formula, we have
$\|P(\cdot)f\|_{C_{T}^{\frac{\alpha}{\gamma}}L^{2}}=$ $\sup_{0\leq s<t\leq T}\frac{\|(P(t)-P(s))f\|_{L^{2}}}{|t-s|^{\frac{\alpha}{\gamma}}}$
$\leq$ $\sup_{0\leq s<t\leq T}\frac{\|e^{-t|k|^{\gamma}}(1-e^{-(t-s)|k|^{\gamma}})\hat{f}(k)\|_{L^{2}}}{|t-s|^{\frac{\alpha}{\gamma}}}\leq\|f\|_{H^{\alpha}},$
for $\alpha\in\mathbb{R}_{+}$, which we prove the first estimate. Note that
$\|f(t)-f(s)\|_{L^{2}}\leq\int_{s}^{t}\|P(t-\tau)\mathcal{L}f(\tau)\|_{L^{2}}\mathrm{d}\tau+\int_{0}^{s}\|[P(t-\tau)-P(s-\tau)]\mathcal{L}f(\tau)\|_{L^{2}}\mathrm{d}\tau.$
The first term can be handled using a method similar to that in Lemma A.2, in fact
$\int_{s}^{t}\|P(t-\tau)\mathcal{L}f(\tau)\|_{L^{2}}\mathrm{d}\tau\lesssim$ $\|\mathcal{L}f\|_{C_{T}H^{\alpha}}\int_{s}^{t}e^{-(t-\tau)|k|^{\gamma}}|k|^{-\alpha}\mathrm{d}\tau$
$\lesssim$ $\|\mathcal{L}f\|_{C_{T}H^{\alpha}}(t-s)^{1+\frac{\alpha}{\gamma}}.$
For the second term, we have estimate
$\int_{0}^{s}\|[P(t-\tau)-P(s-\tau)]\mathcal{L}f(\tau)\|_{L^{2}}\mathrm{d}\tau\lesssim$ $\|\mathcal{L}f\|_{C_{T}H^{\alpha}}\int_{0}^{s}e^{-(s-\tau)|k|^{\gamma}}(e^{-(t-s)|k|^{\gamma}}-1)|k|^{-\alpha}\mathrm{d}s$
$\lesssim$ $\|\mathcal{L}f\|_{C_{T}H^{\alpha}}(t-s)^{\frac{\alpha+\delta}{\gamma}}s^{1-\frac{\delta}{\gamma}}.$
Combining the above estimate, by the definition of $C_{T}^{\frac{\alpha+\delta}{\gamma}}L^{2}$, we have
$\|f\|_{\mathcal{W}_{2}^{\alpha+\delta}(T)}=$ $\sup_{0\leq s<t\leq T}\frac{\|f(t)-f(s)\|_{L^{2}}}{|t-s|^{\frac{\alpha+\delta}{\gamma}}}$
$\leq$ $T^{1-\frac{\delta}{\gamma}}\|\mathcal{L}f\|_{C_{T}H^{\alpha}}.$
Then we finish the proof. ∎
There are also some similar estimates for $\tilde{P}(t)$, which is defined as
$\mathcal{F}(\tilde{P}(t)(f))=\frac{e^{-t|\xi|^{\gamma}}}{1+|\xi|^{2}}\mathcal{F}f.$
To clear that, fix $\varphi(t,z)=\frac{e^{-|z|^{\gamma}}}{1+t^{-\frac{2}{\gamma}}|z|^{2}}.$ We have $\tilde{P}(t)f=\varphi(t,t^{\frac{1}{\gamma}}D)f=\tilde{G}(t,x)f$. For fixed $t\in[0,T]$, it’s easy to see that $\mathcal{F}^{-1}\varphi(t,z)\in L^{1}(\mathbb{R})$. In fact, by Young’s inequality and (A.1)
$\|\mathcal{F}^{-1}_{z}\varphi(t,z)(\xi)\|_{L^{1}}=$ $\|\mathcal{F}^{-1}{e^{-|z|^{\gamma}}}\ast\mathcal{F}(\frac{1}{1+t^{-\frac{2}{\gamma}}|z|^{2}})\|_{L^{1}}$
$\leq$ $\|\check{g}\|_{L^{1}}\|\mathcal{F}(\frac{1}{1+t^{-\frac{2}{\gamma}}|z|^{2}})\|_{L^{1}}$
the first integral is always finite. For the second term, by $\mathcal{F}^{-1}_{z}{\frac{1}{1+|z|^{2}}}=\pi e^{-|\xi|}$, we have $\mathcal{F}(\frac{1}{1+t^{-\frac{2}{\gamma}}|z|^{2}})=\pi t^{\frac{1}{\gamma}}e^{-t^{\frac{1}{\gamma}}|\xi|}$ which also in $L^{1}$, which obtain the following claim
$\|\tilde{P}(t)f\|_{B_{p,r}^{\alpha}}\lesssim\|f\|_{B_{p,r}^{\alpha}}.$
On the other hand, for $j\geq 1$, we have the estimate
$\|(\mathcal{F}^{-1}\varphi(t,t^{\frac{1}{\gamma}}z))\ast\Delta_{j}u\|_{L^{p}}\lesssim$ $\|\mathcal{F}^{-1}(\varphi(t,t^{\frac{1}{\gamma}}z)\psi(2^{-j}z))\|_{L^{1}}\|\Delta_{j}u\|_{L^{p}}$
$\lesssim$ $\|\mathcal{F}^{-1}(\varphi(t,t^{\frac{1}{\gamma}}2^{j}z)\psi)\|_{L^{1}}\|\Delta_{j}u\|_{L^{p}}$
$\lesssim$ $\|(1+|x|)^{d+1}\mathcal{F}^{-1}(\varphi(t,t^{\frac{1}{\gamma}}2^{j}z)\psi)\|_{L^{\infty}}\|\Delta_{j}u\|_{L^{p}}$
$\lesssim$ $\|(1+|D|)^{d+1}\varphi(t,t^{\frac{1}{\gamma}}2^{j}z)\psi\|_{L^{1}}\|\Delta_{j}u\|_{L^{p}}$
$\lesssim$ $(1+2^{j}t^{\frac{1}{\gamma}})^{d+1}\max_{\mu\in\mathbb{N}^{d}:|\mu|\leq d+1}\|\partial^{\mu}\varphi(t,t^{\frac{1}{\gamma}}2^{j}\cdot)\|_{L^{\infty}(\mathrm{supp}(\psi))}\|\Delta_{j}u\|_{L^{p}}.$
By Lebnitz’rule $\partial^{2}(uv)=\sum_{\mu_{1}+\mu_{2}=2}\frac{\alpha!}{\mu_{1}!\mu_{2}!}\partial^{\mu_{1}}u\partial^{\mu_{2}}v$, we have
$\partial^{2}\varphi(t,z)=\sum_{\mu_{1}+\mu_{2}=2}\frac{\alpha!}{\mu_{1}!\mu_{2}!}\partial^{\mu_{1}}(e^{-|z|^{\gamma}})\partial^{\mu_{2}}(\frac{1}{1+t^{-\frac{2}{\gamma}}|z|^{2}})=c_{1}u^{\prime\prime}v+c_{2}u^{\prime}v^{\prime}+c_{3}uv^{\prime\prime},$
where $u(z)=e^{-|z|^{\gamma}}$ and $v(z),v^{\prime}(z),v^{\prime\prime}(z)$ satisfied
$v:=\frac{1}{1+cz^{2}},\quad v^{\prime}:=\frac{-2cz}{(1+cz^{2})^{2}},\quad v^{\prime\prime}=\frac{6c^{2}z^{2}-2c}{(1+cz^{2})^{3}}.$
Since $cz^{2}>0$, it’s obviously that $|v|,|v^{\prime}|,|v^{\prime\prime}|\lesssim 1$ which shows that
$\sup_{|\mu|\leq 2}\sup_{z\geq 1}(1+|z|)^{\delta+2}|\partial_{z}^{\mu}\varphi(t,z)|\lesssim 1.$
On the other hand, since all partial derivatives of $u=e^{-|z|^{\gamma}}$ is decay faster than rational function for $\gamma>0$, we have
$\sup_{|\mu|\leq 2}\sup_{z\geq 1}(1+|z|)^{\delta+2}|\partial_{z}^{\mu}\varphi(t,z)|\lesssim$ $\sup_{z\geq 1}(1+|z|)^{\delta+2d}\sum_{|\mu|\leq 2}|\partial_{z}^{\mu}\varphi(t,z)|$
$\lesssim$ $c^{-1}\sup_{z\geq 1}(1+|z|)^{\delta}|(u^{\prime\prime}+u^{\prime}+u)\frac{cz^{2}}{1+cz^{2}}|$
$+c^{-1}\sup_{z\geq 1}(1+|z|)^{\delta-1}|(u^{\prime}+u)\frac{-2c^{2}z^{4}}{(1+cz^{2})^{2}}|$
$+c^{-1}\sup_{z\geq 1}(1+|z|)^{\delta-2}|u\frac{6c^{3}z^{6}-2c^{2}z^{4}}{(1+cx^{2})^{3}}|$
$\lesssim$ $c^{-1}.$
Let $c=t^{-\frac{2}{\gamma}}$ and $d=1$, then there exists $j_{0}$ satisfying that $2^{j_{0}}t^{\frac{1}{\gamma}}\geq 1$ such that
$\|\tilde{P}(t)\Delta_{j}f(t)\|_{L^{p}}\lesssim$ $(1+2^{j}t^{\frac{1}{\gamma}})^{2}(1+2^{j}t^{\frac{1}{\gamma}})^{-2-\delta}\min(t^{\frac{2}{\gamma}},1)$
$\lesssim$ $\min(t^{\frac{2}{\gamma}},1)t^{-\frac{\delta}{\gamma}}2^{-j\delta}\|\Delta_{j}f(t)\|_{L^{p}}.$
For $1\leq j<j_{0}$, we have the estimate
$\|\tilde{P}(t)\Delta_{j}f(t)\|_{L^{p}}\lesssim$ $\|\Delta_{j}f(t)\|_{L^{p}}$
$\lesssim$ $(\max(t^{-\frac{2}{\gamma}},1)2^{j_{0}\delta}t^{\frac{\delta}{\gamma}})\min(t^{\frac{2}{\gamma}},1)t^{-\frac{\delta}{\gamma}}2^{-j\delta}\|\Delta_{j}f(t)\|_{L^{p}}$
$\lesssim$ $\min(t^{\frac{2}{\gamma}},1)t^{-\frac{\delta}{\gamma}}2^{-j\delta}\|\Delta_{j}f(t)\|_{L^{p}}.$
Combining the proceeding results, we obtain the following lemma
Corollary A.1.
Let $\gamma>0$, $\tilde{P}(t)$ satisfy $\mathcal{F}(\tilde{P}(t)(f))=\frac{e^{-t|\xi|^{\gamma}}}{1+|\xi|^{2}}\mathcal{F}f$. Fixing some $T>0$, then for any $t\in(0,T]$, $\alpha\in\mathbb{R}$, $\delta\in(0,\gamma)$, and $u\in S^{\prime}(\mathbb{R})$, we have the following estimate
$\|\tilde{P}(t)u\|_{B^{\alpha+\delta}_{p,r}}\lesssim\min(1,t^{\frac{2}{\gamma}})t^{-\frac{\delta}{\gamma}}\|u\|_{B^{\alpha}_{p,r}}.$ (A.7)
Moreover, let $\tilde{L}f=(\partial_{t}-\Lambda^{\gamma})(1-\partial_{x}^{2})^{-1}f\in B_{p,r}^{\alpha}$, then for $\delta\in[0,\gamma+2)$ and $t\leq 1$, we have the following estimate
$\|f\|_{B_{p,r}^{\alpha+\delta}}\leq\|f(0)\|_{B_{p,r}^{\alpha+\delta}}+Ct^{1-\frac{\delta-2}{\gamma}}\|\mathcal{L}f\|_{B_{p,r}^{\alpha}}.$ (A.8)
Appendix B Commutator estimates
First we introduce a useful commutator estimates as Lemma 2.99 in [BCD11].
Lemma B.1.
[[BCD11]] Let $f$ be a smooth function on $\mathbb{R}^{d}$ and be homogeneous of degree $m$ away from 0. Let $\rho\in(0,1)$, $s\in\mathbb{R}$ and $p,r\in[1,\infty]$. If $p_{1},p_{2}\in[1,\infty]$ such that $\frac{1}{p_{1}}+\frac{1}{p_{2}}=\frac{1}{p}$, then the bound
$\|a\prec f(D)u-f(D)(a\prec u)\|_{B_{p,r}^{s-m+\rho}}\lesssim\|\nabla a\|_{B_{p_{1},\infty}^{\rho-1}}\|u\|_{B_{p_{2},r}^{s}}.$
is hold.
Remark B.1.
Fixing $f(x)=|x|^{\gamma}$, then we have $\Lambda^{\gamma}(u\prec v)-u\prec\Lambda^{\gamma}v\in W_{p}^{\alpha+\beta-\gamma}$ for $u\in\mathcal{C}^{\alpha}$, $\alpha\in(0,1)$ and $v\in W_{p}^{\beta}$ for $\alpha\in(0,1)$.
Lemma B.2.
Assume that $\alpha\in(0,1)$, $\beta,\gamma\in\mathbb{R}$ and satisfies $\beta+\gamma<0$, $\alpha+\beta+\gamma>0$, let $C(f,g,h)=(f\prec g)\circ h-f(g\circ h)$, then we have the following estimate
$\|C(f,g,h)\|_{H^{\alpha+\beta+\gamma}}\leq\|f\|_{\mathcal{C}^{\alpha}}\|g\|_{H^{\beta}}\|v\|_{\mathcal{C}^{\gamma}}.$
Proof.
We rewrite
$C(f,g,h)=\sum_{j,k\geq-1}\sum_{|i-j|\leq 1}[1_{i\gtrsim k}R_{i}(\Delta_{k}f,g)\Delta_{j}h-1_{i\leq k-N}\Delta_{k}f\Delta_{j}g\Delta_{j}h].$
It’s suffice to estimate above two terms. For any fixed $k$, we have estimate
$\|\sum_{j\geq-1}\sum_{|i-j|\leq 1}1_{i\leq k-N}\Delta_{k}f\Delta_{j}g\Delta_{j}h\|_{L^{2}}\lesssim$ $2^{-k\alpha}\|f\|_{\mathcal{C}^{\alpha}}\sum_{i=-1}^{k-N}2^{-i(\beta+\gamma)}\|g\|_{H^{\beta}}\|h\|_{\mathcal{C}^{\gamma}}$
$\lesssim$ $2^{-k(\alpha+\beta+\gamma)}\|f\|_{\mathcal{C}^{\alpha}}\|g\|_{H^{\beta}}\|h\|_{\mathcal{C}^{\gamma}},$
where we use the fact $\beta+\gamma<0$. Note that the Fourier transform of $\sum_{j,k\geq-1}\sum_{|i-j|\leq 1}1_{i\gtrsim k}R_{i}(\Delta_{k}f,g)\Delta_{j}h$ is supported in a ball $2^{k}\mathcal{B}$, by Lemma 2.84 in [BCD11], we obtain the estimate since $\alpha+\beta+\gamma>0$. For the first series, we can check that the Fourier transform of $\sum_{k\geq-1}\sum_{|i-j|\leq 1}1_{i\gtrsim k}R_{i}(\Delta_{k}f,g)\Delta_{j}h$ is supported in a ball $2^{j}\mathcal{B}$ for fixed $j$. And we calculate that
$\|\sum_{k\geq-1}\sum_{|i-j|\leq 1}1_{i\gtrsim k}R_{i}(\Delta_{k}f,g)\Delta_{j}h\|_{L^{2}}=$ $\|\sum_{|i-j|\leq 1}R_{i}(\sum_{1\leq k\lesssim i}\Delta_{k}f,g)\Delta_{j}h\|_{L^{2}}$
$\lesssim$ $\sum_{|i-j|\leq 1}2^{-i(\beta+\gamma)}\|\sum_{k\lesssim i}\Delta_{k}f\|_{L^{\infty}}\|g\|_{H^{\beta}}2^{-j\gamma}\|h\|_{\mathcal{C}^{\gamma}}$
$\lesssim$ $2^{-j(\alpha+\beta+\gamma)}\|f\|_{\mathcal{C}^{\alpha}}\|g\|_{H^{\beta}}\|h\|_{\mathcal{C}^{\gamma}}.$
Combining all estimates above, we finish our proof. ∎
Lemma B.3.
Let $\alpha\in(0,\gamma)$ and $\beta\in\mathbb{R}$. Then for $t>0$, we have the following estimate
$\|f\prec\!\!\!\prec g-f\prec g\|_{W^{\alpha+\beta}}\lesssim\|f\|_{\mathcal{W}_{\infty}^{\alpha}(t)}\|g\|_{W^{\beta}}.$ (B.1)
Moreover, define operator $\mathcal{L}=\partial_{t}-\Lambda^{\gamma}$, then for $\gamma>0$ and $\alpha\in(0,1)$, we have the following estimate
$\|\mathcal{L}(f\prec\!\!\!\prec g)(t)-(f\prec\!\!\!\prec\mathcal{L}g)(t)\|_{W^{\alpha+\beta-\gamma}}\lesssim\|f\|_{\mathcal{W}_{\infty}^{\alpha}(t)}\|g\|_{W^{\beta}}.$ (B.2)
Proof.
We only proof the case of $H^{\alpha}$, the case of $\mathcal{C}^{\alpha}$ is similar. Firstly, control
$\|(\int_{\mathbb{R}}2^{\gamma j}\varphi(2^{\gamma j}(t-s))S_{j-1}f(s)\mathrm{d}s-S_{j-1}f(t))\Delta_{j}g(t)\|_{L^{2}}.$
By $\int_{\mathbb{R}}\varphi(s)\mathrm{d}s=1$, we have
$\|(\int_{\mathbb{R}}2^{\gamma j}\varphi(2^{\gamma j}(t-s))S_{j-1}f(s)\mathrm{d}s$- $S_{j-1}f(t))\Delta_{j}g(t)\|_{L^{2}}$
$=$ $\|\int_{\mathbb{R}}2^{\gamma j}\varphi(2^{2j}s)S_{j-1}f(t-s)-S_{j-1}f(t)\mathrm{d}s\Delta_{j}g(t)\|_{L^{2}}$
$\lesssim$ $\|\Delta_{j}g(t)\|_{L^{2}}\int_{\mathbb{R}}2^{\gamma j}\varphi(2^{\gamma j}s)\|S_{j-1}f(t-s)-S_{j-1}f(t)\|_{L^{\infty}}\mathrm{d}s$
$\lesssim$ $\|\Delta_{j}g(t)\|_{L^{2}}\|f(t)\|_{C_{t}^{\frac{\alpha}{\gamma}}L^{\infty}}\int_{\mathbb{R}}2^{\gamma j}\varphi(2^{\gamma j}s)|s|^{\frac{\alpha}{\gamma}}\mathrm{d}s$
$\lesssim$ $2^{-j(\alpha+\beta)}(2^{j\beta}\|\Delta_{j}g(t)\|_{L^{2}})\|f(t)\|_{\mathcal{W}_{\infty}^{\alpha}(t)}.$
Multiplying $2^{-j(\alpha+\beta)}$ both sides to the above inequality and summing over $j$ in the $l^{2}$, we prove (B.1).
For the second result, we firstly control the term of fractional Laplacian commutator
$\|\Lambda^{\gamma}(f\prec\!\!\!\prec g)-f\prec\!\!\!\prec\Lambda^{\gamma}g\|_{C_{T}W^{\alpha+\beta-\gamma}}$
For $f\in C_{t}\mathcal{C}^{\alpha}$, rewriting $\Lambda^{\gamma}(f\prec\!\!\!\prec g)(t)-f\prec\!\!\!\prec\Lambda^{\gamma}g(t)$ as
$\sum_{j\geq 1}\int_{0}^{t}2^{\gamma j}$ $\varphi(2^{\gamma j}(t-s))[\Lambda^{\gamma}(S_{j-1}f(s)\Delta_{j}g(t))-S_{j-1}f(s)\Lambda^{\gamma}(\Delta_{j}g(t))]\mathrm{d}s$
$=$ $\sum_{j\geq 1}\int_{0}^{t}2^{\gamma j}\varphi(2^{\gamma j}(t-s))[S_{j-1}f,\Lambda^{\gamma}\widetilde{\Delta}_{j}]\Delta_{j}g(t)\mathrm{d}s$
where $\tilde{\rho}$ is a smooth function supported in an annuals and with value 1 on a neighborhood of $\mathrm{Supp}\rho+\mathrm{Supp}\chi(\cdot/4)$. Observing that there exists $N_{0}$ such that
$\forall j\geq N_{0},\quad\Lambda^{\gamma}\widetilde{\Delta}_{j}=2^{j\gamma}(|\cdot|^{\gamma}\tilde{\rho})(2^{-j}D).$
Then by Lemma 2.97 in [BCD11], we have for any $j\geq N_{0}$
$\|[S_{j-1}f,\Lambda^{\gamma}\widetilde{\Delta}_{j}]\Delta_{j}g(t)\|_{L^{p}}\leq C2^{j(\gamma-1)}\|S_{j-1}f(s)\|_{L^{\infty}}\|\Delta_{j}g(t)\|_{L^{p}},$
and for $1\leq j<N_{0}$, we have
$\|[S_{j-1}f,\Lambda^{\gamma}\widetilde{\Delta}_{j}]\Delta_{j}g(t)\|_{L^{p}}\leq C2^{j(\gamma-1)}\|\nabla S_{j-1}f(s)\|_{L^{\infty}}\|\Delta_{j}g(t)\|_{L^{p}}$
Since $\|\nabla S_{j-1}f(s)\|_{L^{\infty}}\leq C2^{j(1-\alpha)}\|\nabla f(s)\|_{\mathcal{C}^{\alpha-1}}\leq C2^{j(1-\alpha)}\|f\|_{C_{t}\mathcal{C}^{\alpha}}$ for $\alpha<1$ and $0<s<t$. then we have
$2^{j(\beta+\alpha-\gamma)}\|\int_{0}^{t}2^{\gamma j}\varphi(2^{\gamma j}(t-s))[S_{j-1}f,\Lambda^{\gamma}\widetilde{\Delta}_{j}]\Delta_{j}g(t)\mathrm{d}s\|_{L^{p}}\lesssim\|f\|_{C_{t}\mathcal{C}^{\alpha}}2^{j\beta}\|\Delta_{j}g(t)\|_{L^{p}}$
since $S_{j-1}f,\Lambda^{\gamma}\widetilde{\Delta}_{j}]$ is spectrally supported in dyadic annuli, by Lemma 2.23 in [BCD11], we obtain the estimate.
For $\partial_{t}(f\prec\!\!\!\prec g)-f\prec\!\!\!\prec\partial_{t}g$, it’s suffice to estimate
$\sum_{j}\partial_{t}(\int_{0}^{t}2^{\gamma j}\varphi(2^{\gamma j}(t-s))S_{j-1}f(s)\mathrm{d}s)\Delta_{j}g(t).$
For fixd $j$ we recall that $\mathrm{supp}(\varphi)\subset\mathbb{R}_{+}$, and therefore
$\int_{0}^{t}2^{2\gamma j}\varphi^{\prime}(2^{\gamma j}(t-s))S_{j-1}f(s)\mathrm{d}s=2^{\gamma j}\int_{\mathbb{R}}2^{\gamma j}\varphi^{\prime}(2^{\gamma j}(t-s))S_{j-1}f(s)1_{s\geq 0}\mathrm{d}s.$
Since $\varphi^{\prime}(0)=\varphi(0)=0$, (B.2) then follows as the first part of the proof. ∎
Acknowledgments This work was partially supported by the National Natural Science Foundation of China (No.12571261).